### v2

# Hölder regularity of harmonic functions on metric measure spaces

Abstract
We introduce the Hölder regularity condition for harmonic functions on metric measure spaces and prove that under mild volume regular condition and upper heat kernel estimate, the Hölder regularity condition, the weak Bakry-Émery non-negative curvature condition, the heat kernel Hölder continuity with or without exponential terms and the heat kernel near-diagonal lower bound are equivalent. As applications, firstly, we prove the validity of the so-called generalized reverse Hölder inequality on the Sierpiński carpet cable system, which was left open by Devyver, Russ, Yang (Int. Math. Res. Not. IMRN (2023), no. 18, 15537–15583). Secondly, we prove that two-sided heat kernel estimates alone imply gradient estimate for the heat kernel on strongly recurrent fractal-like cable systems, which improves the main results of the aforementioned paper. Thirdly, we obtain Hölder (Lipschitz) estimate for heat kernel on general metric measure spaces, which extends the classical Li-Yau gradient estimate for heat kernel on Riemannian manifolds.
Date: August 18, 2024MSC2010: 28A80, 35K08Keywords: Hölder regularity, gradient estimate, harmonic function, heat kernel, Sierpiński carpet.Jin Gao was supported by National Natural Science Foundation of China (12271282). Jin Gao acknowledged Jiaxin Hu and Eryan Hu for reading the preliminary draft. Part of the work was carried out while Meng Yang was participating a Mini-Symposia in the 9th European Congress of Mathematics (9ECM) in Seville (Spain) between 15 and 19 July 2024, Meng Yang was very grateful to Fabrice Baudoin and Li Chen for the invitation to the symposia.
1 Introduction
Let us recall the following classical de Giorgi-Nash-Moser theorem from PDE theory. Let $D$ be a bounded domain in $\mathbb{R}^{d}$. Consider the following divergence form elliptic operator
$Lu=\sum_{i,j=1}^{d}\frac{\partial}{\partial x_{i}}\left(a_{ij}\frac{\partial u}{\partial x_{j}}\right),$
where $a_{ij}$, $i,j=1,\ldots,d$, are measurable functions in $D$ satisfying that $a_{ij}=a_{ji}$ for $i,j=1,\ldots,d$ and there exist some positive constants $\lambda$, $\Lambda$ such that
$\lambda|\xi|^{2}\leq\sum_{i,j=1}^{d}a_{ij}(x)\xi_{i}\xi_{j}\leq\Lambda|\xi|^{2}\text{ for any }x\in D,\xi=(\xi_{1},\ldots,\xi_{d})\in\mathbb{R}^{d}.$
Then any weak solution $u\in W^{1,2}_{loc}(D)$ of $Lu=0$ in $D$, or ($L$-)harmonic function $u\in W^{1,2}_{loc}(D)$, is locally Hölder continuous in $D$ or $u\in C^{\alpha}(D)$, where $\alpha\in(0,1)$ is the Hölder exponent. One only needs the measurability of the coefficients $a_{ij}$ to obtain the Hölder regularity of the weak solution $u$. To prove this theorem, very powerful iteration techniques were developed, including de Giorgi iteration [24] and Moser iteration [49], and were applied in many areas later, see [18, 36, 21, 32, 13, 29] and references therein for analysis and heat kernel on metric measure spaces. However, iteration techniques give the existence of some Hölder exponent $\alpha$ which depends only on $d$, $\lambda$, $\Lambda$, but may not give an explicit value of $\alpha$.
The main purpose of this paper is to consider Hölder regularity of harmonic functions on metric measure spaces, where certain explicit Hölder exponent can be obtained. These metric measure spaces will originate from certain strongly recurrent fractals.
Fractals are new examples with very different phenomena from Riemannian manifolds. On a complete non-compact Riemannian manifold, it was independently discovered by Grigor’yan [27] and Saloff-Coste [50, 51] that the following two-sided Gaussian estimate of the heat kernel
$\frac{C_{1}}{V(x,\sqrt{t})}\exp\left(-C_{2}\frac{d(x,y)^{2}}{t}\right)\leq p_{t}(x,y)\leq\frac{C_{3}}{V(x,\sqrt{t})}\exp\left(-C_{4}\frac{d(x,y)^{2}}{t}\right),$ HK(2)
is equivalent to the conjunction of the volume doubling condition and the scale-invariant $L^{2}$-Poincaré inequality on balls. However, one important estimate on fractals is the following two-sided sub-Gaussian estimate of the heat kernel
$\frac{C_{1}}{V(x,t^{1/\beta})}\exp\left(-C_{2}\left(\frac{d(x,y)}{t^{1/\beta}}\right)^{\frac{\beta}{\beta-1}}\right)\leq p_{t}(x,y)\leq\frac{C_{3}}{V(x,t^{1/\beta})}\exp\left(-C_{4}\left(\frac{d(x,y)}{t^{1/\beta}}\right)^{\frac{\beta}{\beta-1}}\right),$
where $\beta$ is a new parameter called the walk dimension, which is always strictly greater than 2 on fractals. Moreover, we say that a fractal is strongly recurrent if its Hausdorff dimension $\alpha$ is strictly less than its walk dimension $\beta$. For example, on the Sierpiński gasket (see Figure 2), we have $\alpha=\log 3/\log 2<\beta=\log 5/\log 2$, see [14, 41], on the Sierpiński carpet (see Figure 2), we have $\alpha=\log 8/\log 3<\beta\approx 2.09697$, see [6, 7, 11, 8, 35, 46].
Figure 1: The Sierpiński gasket
Figure 2: The Sierpiński carpet
Another purpose of this paper is to improve some recent results by Devyver, Russ and the second named author [25], where gradient estimate for heat kernel was considered. Gradient estimate for heat kernel plays an important role in the $L^{p}$-boundedness of the Riesz transform for $p>2$, see [3, 22]. To obtain pointwise upper sub-Gaussian estimate for the gradient of the heat kernel on fractal-like cable systems, they introduced the so-called generalized reverse Hölder inequality and verified this condition on the Vicsek and the Sierpiński gasket cable systems, where the proof can be generalized to the cable systems corresponding to a class of p.c.f. self-similar sets. Since the harmonic extension algorithm was intrinsically needed, their proof can not apply on the Sierpiński carpet cable system. In this paper, we introduce the Hölder regularity condition, which is stronger than the generalized reverse Hölder inequality, and, a little bit surprisingly, can be ensured by pointwise two-sided heat kernel estimates. Therefore, on the one hand, we prove the validity of the generalized reverse Hölder inequality on the Sierpiński carpet cable system, which was left open in [25, Page 15545, Line 2]. On the other hand, to obtain gradient estimate for the heat kernel, we do not need to assume gradient type conditions, such as the generalized reverse Hölder inequality, but only pointwise two-sided heat kernel estimates and some mild volume condition are enough, thus we improve main results of [25, Theorem 1.1, Corollary 1.3].
The next purpose of this paper is to generalize the recent result by Baudoin, Chen [15, Theorem 3.3] about Lipschitz estimate for heat kernel on the unbounded Vicsek set. Here we obtain Hölder (Lipschitz) estimate for heat kernel on general unbounded metric measure spaces, which may behave differently at small scale and large scale. We follow a concise analytical approach, thus avoid to going into the probabilistic framework of fractional diffusions by Barlow [5].
To end the Introduction, let us state briefly our main result, more detailed statement will be given in Section 2. We introduce firstly two scaling functions. Let $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\in(0,+\infty)$ satisfy
$22\leq\beta_{i}\leq\alpha_{i}+1,\alpha_{i}<\beta_{i}\text{ for }i=1,$ 2 .
Let $\Phi,\Psi:(0,+\infty)\to(0,+\infty)$ be given by
$\Phi(r)$ $=r^{\alpha_{1}}1_{(0,1)}(r)+r^{\alpha_{2}}1_{[1,+\infty)}(r),$
$\Psi(r)$ $=r^{\beta_{1}}1_{(0,1)}(r)+r^{\beta_{2}}1_{[1,+\infty)}(r).$
In an unbounded metric measure Dirichlet space, we introduce the Hölder regularity condition HR($\Phi,\Psi$) and prove that under mild volume regular condition V($\Phi$) and upper heat kernel estimate UHK($\Psi$), the following conditions are indeed equivalent.
Since we do not need to assume the existence of any gradient operator associated with the metric measure Dirichlet space, in particular, a “carré du champ”, the above equivalence generalizes the corresponding result by Coulhon, Jiang, Koskela, Sikora [22, Theorem 1.2].
In addition, we also extend the classical result on Riemannian manifolds to metric measure Dirichlet spaces, that is, the conjunction of UHK($\Psi$) and $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ implies NLE($\Psi$) on metric measure Dirichlet spaces, which extends that the conjunction of upper and gradient estimates of heat kernel implies lower estimate of heat kernel on Riemannian manifolds, see Cheeger, Yau [19] and Li, Yau [47].
Throughout the paper, the letters $C,C_{1},C_{2},C_{A},C_{B}$ will always refer to some positive constants and may change at each occurrence. The sign $\asymp$ means that the ratio of the two sides is bounded from above and below by positive constants. The sign $\lesssim$ ($\gtrsim$) means that the LHS is bounded by positive constant times the RHS from above (below).
2 Statement of the main results
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded metric measure Dirichlet (MMD) space, that is, $(X,d)$ is a locally compact separable unbounded metric space, $m$ is a positive Radon measure on $X$ with full support, $(\mathcal{E},\mathcal{F})$ is a strongly local regular Dirichlet form on $L^{2}(X;m)$. Throughout this paper, we always assume that all metric balls are relatively compact and $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is conservative.
For any $x\in X$, for any $r\in(0,+\infty)$, denote the (metric) ball $B(x,r)=\{y\in X:d(x,y)<r\}$, denote $V(x,r)=m(B(x,r))$. If $B=B(x,r)$, then we denote $\delta B=B(x,\delta r)$ for any $\delta\in(0,+\infty)$. Let $C(X)$ denote the space of all real-valued continuous functions on $X$ and let $C_{c}(X)$ denote the space of all real-valued continuous functions on $X$ with compact support.
Consider the strongly local regular Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$. Let $\Delta$ be the corresponding generator which is a non-negative definite self-adjoint operator. Let $\Gamma$ be the corresponding energy measure. Denote $\mathcal{E}_{1}(\cdot,\cdot)=\mathcal{E}(\cdot,\cdot)+(\cdot,\cdot)$, where (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) is the inner product in $L^{2}(X;m)$. We refer to [26] for related results about Dirichlet forms.
We need to introduce some conditions to state our main results.
We say that the volume doubling condition VD holds if there exists $C_{VD}\in(0,+\infty)$ such that
$V(x,2r)\leq C_{VD}V(x,r)\text{ for any }x\in X,r\in(0,+\infty).$ VD
We say that the volume regular condition V($\Phi$) holds if there exists $C_{VR}\in(0,+\infty)$ such that
$\frac{1}{C_{VR}}\Phi(r)\leq V(x,r)\leq C_{VR}\Phi(r)\text{ for any }x\in X,r\in(0,+\infty).$ V($\Phi$)
Consider the regular Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$. Let $\left\{P_{t}\right\}$ be the corresponding heat semi-group. Let $\left\{X_{t},t\geq 0,\mathbb{P}_{x},x\in X\backslash\mathcal{N}_{0}\right\}$ be the corresponding Hunt process, where $\mathcal{N}_{0}$ is a properly exceptional set, that is, $m(\mathcal{N}_{0})=0$ and $\mathbb{P}_{x}(X_{t}\in\mathcal{N}_{0}\text{ for some }t>0)=0$ for any $x\in X\backslash\mathcal{N}_{0}$. For any bounded Borel function $f$, we have $P_{t}f(x)=\mathbb{E}_{x}f(X_{t})$ for any $t>0$, for any $x\in X\backslash\mathcal{N}_{0}$.
The heat kernel $p_{t}(x,y)$ associated with the heat semi-group $\left\{P_{t}\right\}$ is a measurable function defined on $(0,+\infty)\times(X\backslash\mathcal{N}_{0})\times(X\backslash\mathcal{N}_{0})$ satisfying that:
  • For any bounded Borel function $f$, for any $t>0$, for any $x\in X\backslash\mathcal{N}_{0}$, we have
    $P_{t}f(x)=\int_{X\backslash\mathcal{N}_{0}}p_{t}(x,y)f(y)m(\mathrm{d}y).$
  • For any $t,s>0$, for any $x,y\in X\backslash\mathcal{N}_{0}$, we have
    $p_{t+s}(x,y)=\int_{X\backslash\mathcal{N}_{0}}p_{t}(x,z)p_{s}(z,y)m(\mathrm{d}z).$
  • For any $t>0$, for any $x,y\in X\backslash\mathcal{N}_{0}$, we have $p_{t}(x,y)=p_{t}(y,x)$.
See [32] for more details.
We say that the heat kernel upper bound UHK($\Psi$) holds if there exist a properly exceptional set $\mathcal{N}$ and $C_{1}$, $C_{2}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y\in X\backslash\mathcal{N}$, we have
$p_{t}(x,y)\leq\frac{C_{1}}{V\left(x,\Psi^{-1}(t)\right)}\exp\left(-\Upsilon\left(C_{2}d(x,y),t\right)\right),$ UHK($\Psi$)
where
$\Upsilon(R,t)=\sup_{s\in(0,+\infty)}\left(\frac{R}{s}-\frac{t}{\Psi(s)}\right)\asymp\begin{cases}\left(\frac{R}{t^{1/{\beta_{1}}}}\right)^{\frac{\beta_{1}}{\beta_{1}-1}},&\text{if }t<R,\\ \left(\frac{R}{t^{1/{\beta_{2}}}}\right)^{\frac{\beta_{2}}{\beta_{2}-1}},&\text{if }t\geq R.\end{cases}$
Then UHK($\Psi$) can also be re-written as follows:
$p_{t}(x,y)\leq\begin{cases}\frac{C_{1}}{V\left(x,\Psi^{-1}(t)\right)}\exp\left(-C_{2}\left(\frac{d(x,y)}{t^{1/{\beta_{1}}}}\right)^{\frac{\beta_{1}}{\beta_{1}-1}}\right),&\text{if }t<d(x,y),\\ \frac{C_{1}}{V\left(x,\Psi^{-1}(t)\right)}\exp\left(-C_{2}\left(\frac{d(x,y)}{t^{1/{\beta_{2}}}}\right)^{\frac{\beta_{2}}{\beta_{2}-1}}\right),&\text{if }t\geq d(x,y).\end{cases}$
If a lower bound, similar to UHK($\Psi$), with different constants $C_{i}$ also holds, then we say that HK($\Psi$) holds.
We say that the heat kernel near-diagonal lower bound NLE($\Psi$) holds if there exist a properly exceptional set $\mathcal{N}$ and $C,\varepsilon\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y\in X\backslash\mathcal{N}$ with $d(x,y)<\varepsilon\Psi^{-1}(t)$, we have
$p_{t}(x,y)\geq\frac{C}{V(x,\Psi^{-1}(t))}.$ NLE($\Psi$)
If the metric $d$ is furthermore assumed to satisfy the chain condition, then the conjunction of UHK($\Psi$) and NLE($\Psi$) is equivalent to HK($\Psi$), see [32, Section 6].
Let $D$ be an open subset of $X$. Let
$\mathcal{F}_{D}=\text{ the }\mathcal{E}_{1}\text{-closure of }\mathcal{F}\cap C_{c}(D).$
Let $f\in L^{1}_{loc}(D)$. We say that $u\in\mathcal{F}$ is a solution of the Poisson equation $\Delta u=f$ in $D$ if
$\mathcal{E}(u,\varphi)=\int_{D}f\varphi\mathrm{d}m\text{ for any }\varphi\in\mathcal{F}\cap C_{c}(D).$
If $\Delta u=f$ in $D$ with $f\in L^{2}(D)$, then the above equation also holds for any $\varphi\in\mathcal{F}_{D}$. We say that $u\in\mathcal{F}$ is harmonic in $D$ if $\Delta u=0$ in $D$.
Denote $\gamma_{i}=\beta_{i}-\alpha_{i}\in(0,+\infty)$ for $i=1,2$, then
$\left(\frac{\Psi}{\Phi}\right)(r)=r^{\gamma_{1}}1_{(0,1)}(r)+r^{\gamma_{2}}1_{[1,+\infty)}(r).$
We say that the Hölder regularity condition HR($\Phi,\Psi$) holds if there exists $C_{H}\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $u\in\mathcal{F}$ which is harmonic in $2B$, for $m$-a.e. $x,y\in B$ with $x\neq y$, we have
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u(x)-u(y)|\leq C_{H}\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$ HR($\Phi,\Psi$)
The above condition is some kind of generalization of the so-called generalized reversed Hölder inequality GRH($\Phi,\Psi$) introduced in [25].
We say that the Hölder estimate HHK($\Phi,\Psi$) for the heat kernel holds if there exists $C_{HHK}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x_{1},x_{2},y_{1},y_{2}\in X$, we have
$|p_{t}(x_{1},y_{1})-p_{t}(x_{2},y_{2})|\leq C_{HHK}\frac{\left(\frac{\Psi}{\Phi}\right)(d(x_{1},x_{2}))+\left(\frac{\Psi}{\Phi}\right)(d(y_{1},y_{2}))}{t}.$ HHK($\Phi,\Psi$)
The above condition implies that for any $t\in(0,+\infty)$, for any $x\in X$, $p_{t}(x,\cdot)$ is $\gamma_{1}$-Hölder continuous in the small scale and $\gamma_{2}$-Hölder continuous in the large scale.
We say that the Hölder estimate $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ with exponential terms for the heat kernel holds if there exist $C_{1},C_{2}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y_{1},y_{2}\in X$, we have
$|p_{t}(x,y_{1})-p_{t}(x,y_{2})|$
$\leq C_{1}\frac{\left(\frac{\Psi}{\Phi}\right)(d(y_{1},y_{2}))}{t}\left(\exp\left(-\Upsilon\left(C_{2}d(x,y_{1}),t\right)\right)+\exp\left(-\Upsilon\left(C_{2}d(x,y_{2}),t\right)\right)\right).$ $\text{HHK}_{\text{exp}}(\Phi,\Psi)$
We say that the weak Bakry-Émery non-negative curvature condition wBE($\Phi,\Psi$) holds if there exists $C_{wBE}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $f\in L^{\infty}(X;m)$, for any $x,y\in X$, we have
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|P_{t}f(x)-P_{t}f(y)|\leq C_{wBE}\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\lVert f\rVert_{L^{\infty}(X;m)}.$ wBE($\Phi,\Psi$)
The above condition is some kind of generalization of the weak Bakry-Émery curvature condition introduced in [1, Definition 3.1].
The main result of this paper is as follows.
Theorem 2.1.
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$) and
UHK($\Psi$). Then the followings are equivalent.
  1. (1)
  2. (2)
  3. (3)
  4. (4)
  5. (5)
The key novality of the above result is that the pointwise two-sided heat kernel estimates are enough to give some proper explicit Hölder regularity of harmonic functions. In previous literatures (for example [36, Section 5.3], [21, Proposition 4.5], [32, Section 5.3], [13, Section 4.2.1]), using Moser iteration technique or de Giorgi iteration technique and oscillation inequality, one can obtain only some Hölder exponent depending implicitly on other constants.
Our proof will follow the following diagrams.
[…]
and
[…]
Remark 2.2.
In the setting of fractional metric spaces with fractional diffusions introduced in [5, Section 3], when the dimension $\alpha$ of the state space is smaller than the walk dimension $\beta$, or $\alpha_{1}=\alpha_{2}=\alpha<\beta_{1}=\beta_{2}=\beta$ in our setting, it was proved in [1, Lemma 3.4, Remark 3.6, Theorem 3.7] that
[…]
Remark 2.3.
In the fractal setting, HHK($\Phi,\Psi$) was also proved on the unbounded Sierpiński gasket [14, Theorem 1.5 (c)] and on the unbounded Sierpiński carpet [8, Theorem 1.1 (e)].
Remark 2.4.
The above equivalence can be regarded as some “Hölder” version generalization of the “gradient” version result [22, Theorem 1.2]. It was proved that in a non-compact doubling Dirichlet metric measure space $(X,d,\mu,\mathscr{E})$ endowed with a “carré du champ”. Assume the upper Gaussian bound $(UE)$ for the heat kernel and a local $L^{\infty}$-Poincaré inequality $(P_{\infty,\text{loc}})$. Then the following conditions are equivalent.
  • The quantitative reverse $L^{\infty}$-Hölder inequality ($\text{RH}_{\infty}$) for gradients of harmonic functions: There exists $C\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $u$ which is harmonic in $2B$, we have
    $\lVert|\nabla u|\rVert_{L^{\infty}(B)}\leq\frac{C}{r}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}\mu.$ ($\text{RH}_{\infty}$)
  • The pointwise Li-Yau gradient estimate ($\text{GLY}_{\infty}$) for heat kernel: There exist $C,c\in(0,+\infty)$ such that
    $|\nabla_{x}h_{t}(x,y)|\leq\frac{C}{\sqrt{t}V(y,\sqrt{t})}\exp\left\{-c\frac{d(x,y)^{2}}{t}\right\}$ ($\text{GLY}_{\infty}$)
    for any $t\in(0,+\infty)$, for $\mu$-a.e. $x,y\in X$.
  • The $L^{\infty}$-boundedness ($\text{G}_{\infty}$) of the gradient of the heat semi-group $|\nabla H_{t}|$: There exists $C\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, we have
    $\lVert|\nabla H_{t}|\rVert_{\infty\to\infty}\leq\frac{C}{\sqrt{t}},$
    that is, for any $f\in L^{\infty}(X)$, we have
    $\lVert|\nabla H_{t}f|\rVert_{L^{\infty}(X)}\leq\frac{C}{\sqrt{t}}\lVert f\rVert_{L^{\infty}(X)}.$ ($\text{G}_{\infty}$)
Our conditions are “Hölder” version of the above “gradient” conditions as in Figure 3.
“Hölder” conditions “gradient” conditions
HR($\Phi,\Psi$) ($\text{RH}_{\infty}$)
$\text{HHK}_{\text{exp}}(\Phi,\Psi)$ ($\text{GLY}_{\infty}$)
wBE($\Phi,\Psi$) ($\text{G}_{\infty}$)
Figure 3: The correspondence between “Hölder” and “gradient” conditions
In their proof to obtain ($\text{RH}_{\infty}$), they established a reproducing formula for harmonic functions using the finite propagation speed property, which may not be available on general MMD spaces. We leave it as an open question to generalize the reproducing formula technique to prove HR($\Phi,\Psi$) on MMD spaces or the following GRH($\Phi,\Psi$) on cable systems, in particular, to develop the corresponding finite propagation speed theory.
Remark 2.5.
Let us compare the stability of “gradient” and “Hölder” conditions as follows.
By a standard technique, it is easy to see that under $(UE)$, any one of the equivalent “gradient” conditions implies the heat kernel near-diagonal lower Gaussian bound $(NLE)$. In particular, under $(UE)$, we have ($\text{GLY}_{\infty}$) implies $(NLE)$. However, the converse is usually not true. Indeed, the conjunction of $(UE)$ and $(NLE)$, or equivalently, HK(2), is equivalent to the conjunction of the volume doubling condition and the scale-invariant $L^{2}$-Poincaré inequality on balls, hence is quasi-isometry stable. However, to obtain ($\text{GLY}_{\infty}$), some curvature assumptions are usually needed, see [47, 16], hence ($\text{GLY}_{\infty}$) is quasi-isometry unstable.
For comparison, by our main result, Theorem 2.1, under UHK($\Psi$), all the “Hölder” conditions are equivalent to NLE($\Psi$). By the quasi-isometry stability of UHK($\Psi$) (see [2]) and the conjunction of UHK($\Psi$) and NLE($\Psi$) (see [31]), our “Hölder” conditions are quasi-isometry stable. Roughly speaking, the stability originates from the strongly recurrent condition, which requires our MMD spaces to behave like $\mathbb{R}^{1}$ in the classical Riemannian setting.
Let $(V,E)$ be an infinite, locally bounded, connected (undirected) graph. Let
$(X,d,m,\mathcal{E},\mathcal{F})$ be the corresponding unbounded cable system with an elementary gradient operator $\nabla$, see [25, Section 3] for more details. Let $\alpha_{1}=1$, $\beta_{1}=2$, $\alpha_{2}=\alpha$, $\beta_{2}=\beta$.
We say that the generalized reverse Hölder inequality GRH($\Phi,\Psi$) holds if there exists $C_{H}\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $u\in\mathcal{F}$ which is harmonic in $2B$, we have
$\lVert\lvert\nabla u\rvert\rVert_{L^{\infty}(B)}\leq C_{H}\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$ GRH($\Phi,\Psi$)
It is obvious that on an unbounded cable system, HR($\Phi,\Psi$) implies GRH($\Phi,\Psi$). Indeed, assume HR($\Phi,\Psi$), then the LHS of HR($\Phi,\Psi$) becomes $|u(x)-u(y)|/d(x,y)$ for $d(x,y)\in(0,1)$. Letting $y\to x$ and using the Lebesgue density theorem, we obtain the limit $|\nabla u(x)|$. Taking essential supremum with respect to $x\in B$, we have GRH($\Phi,\Psi$). Roughly speaking, HR($\Phi,\Psi$) encodes more Hölder continuity information than GRH($\Phi,\Psi$) about the behavior of harmonic functions at large scale.
We give some results as important applications of our main result, Theorem 2.1, on unbounded cable systems. Firstly, due to the fact that two-sided heat kernel estimates have already obtained on a large class of fractals (see [5, 34, 42, 43]), including the Sierpiński carpet (see [8, 9, 35]), we have GRH($\Phi,\Psi$) holds on a large class of fractal-like cable systems, including the Sierpiński carpet cable system, whose validity was left open in [25].
Corollary 2.6.
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded cable system satisfying V($\Phi$) and
HK($\Psi$) with $\alpha<\beta$. Then GRH($\Phi,\Psi$) holds. In particular, GRH($\Phi,\Psi$) holds on the Sierpiński carpet cable system with $\alpha=\log 8/\log 3<2<\beta\approx 2.09697$.
Remark 2.7.
In [25], the authors verified directly GRH($\Phi,\Psi$) on the Vicsek cable system and the Sierpiński gasket cable system using certain Hölder regularity  [^1] The Hölder regularity follows indeed from certain extension algorithm for harmonic functions. of harmonic functions on the corresponding fractals [52], which is also available on some class of p.c.f. self-similar sets [54] without any technical difficulty, but not available on general fractals, especially on the Sierpiński carpet.
Secondly, we improve the results about gradient estimate for heat kernel as follows.
Corollary 2.8.
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded cable system satisfying V($\Phi$) and
HK($\Psi$) with $\alpha<\beta$.
  1. (1)
    Then $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ holds. In particular, we have the gradient estimate GHK($\Phi,\Psi$) for the heat kernel as follows: there exist $C_{1},C_{2}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for $m$-a.e. $x,y\in X$, we have
    $|\nabla_{y}p_{t}(x,y)|$ $\leq\frac{C_{1}\Phi(\Psi^{-1}(t))}{tV\left(x,\Psi^{-1}(t)\right)}\exp\left(-\Upsilon\left(C_{2}d(x,y),t\right)\right).$ GHK($\Phi,\Psi$)
  2. (2)
    There exist $C_{3},C_{4}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for $m$-a.e. $x,y\in X$, we have
    $\left|\nabla_{y}p_{t}(x,y)\right|\leq C_{3}\frac{\Phi(\Psi^{-1}(t))}{t}p_{C_{4}t}(x,y).$
Remark 2.9.
(1) is certain improvement of [25, Theorem 1.1] and (2) is certain improvement of [25, Corollary 1.3], where GRH($\Phi,\Psi$) was needed. Now we can obtain gradient estimate result GHK($\Phi,\Psi$) without assuming gradient type conditions like GRH($\Phi,\Psi$).
As a direct consequence of $\text{HHK}_{\text{exp}}(\Phi,\Psi)$, we also obtain the following Hölder estimate for the heat kernel on unbounded MMD spaces.
Corollary 2.10.
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$)
and HK($\Psi$). Then there exist $C,c\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y_{1},y_{2}\in X$, we have
$|p_{t}(x,y_{1})-p_{t}(x,y_{2})|\leq C\frac{\left(\frac{\Psi}{\Phi}\right)(d(y_{1},y_{2}))}{\left(\frac{\Psi}{\Phi}\right)(\Psi^{-1}(t))}\left(p_{ct}(x,y_{1})+p_{ct}(x,y_{2})\right).$
Remark 2.11.
If $\beta_{1}=\alpha_{1}+1$, then the above Hölder estimate gives a local Lipschitz estimate, which will imply the existence and some estimate of the gradient for the heat kernel. We give some examples as follows.
  • Unbounded cable systems satisfying V($\Phi$) and HK($\Psi$). We have $\alpha_{1}=1$ and $\beta_{1}=2$ and this gives indeed also Corollary 2.8.
  • The unbounded Vicsek set. We have $\alpha_{1}=\alpha_{2}=\log 5/\log 3$, $\beta_{1}=\beta_{2}=\log 15/\log 3$ and the above Hölder estimate is indeed a Lipschitz estimate, which was also given by Baudoin, Chen [15, Theorem 3.3]. However, a resolvent kernel estimate [5, Theorem 3.40] was intrinsically needed in their proof.
  • The blowups of the Vicsek set by another fractal. We have $\alpha_{1}=\log 5/\log 3$ and $\beta_{1}=\log 15/\log 3$. $\alpha_{2}$ and $\beta_{2}$ will be the Hausdorff dimension and walk dimension of another fractal, say the Sierpiński carpet or the unit interval, see Figure 7 or Figure 8, respectively. The estimates of the gradient for heat kernel can behave differently at small scale and large scale. We will give a detailed discussion in Section 9 about these two examples.
[…]
3 Heat kernel estimates
In this section, we list some characterizations of heat kernel estimates for later use.
Let $D$ be an open subset of $X$. Denote by $\lambda_{1}(D)$ the smallest Dirichlet eigenvalue for $D$, that is,
$\lambda_{1}(D)=\inf\left\{\frac{\mathcal{E}(u,u)}{\lVert u\rVert_{2}^{2}}:u\in\mathcal{F}_{D}\backslash\left\{0\right\}\right\}.$
We say that the relative Faber-Krahn inequality FK($\Psi$) holds if there exist $C_{F}\in(0,+\infty)$, $\nu\in(0,1)$ such that for any ball $B=B(x,r)$, for any open subset $D$ of $B$, we have
$\lambda_{1}(D)\geq\frac{C_{F}}{\Psi(r)}\left(\frac{m(B)}{m(D)}\right)^{\nu}.$ FK($\Psi$)
We will sometimes use the notation FK($\Psi,\nu$) to emphasize the role of the value of $\nu$.
We say that the Poincaré inequality PI($\Psi$) holds if there exists $C_{P}\in(0,+\infty)$ such that for any ball $B=B(x,r)$, for any $u\in\mathcal{F}$, we have
$\int_{B}|u-u_{B}|^{2}\mathrm{d}m\leq C_{P}\Psi(r)\int_{2B}\mathrm{d}\Gamma(u,u),$ PI($\Psi$)
where $u_{A}$ is the mean value of $u$ on a measurable set $A$ with $m(A)\in(0,+\infty)$, that is,
$u_{A}=\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{A}u\mathrm{d}m=\frac{1}{m(A)}\int_{A}u\mathrm{d}m.$
Let $U,V$ be two open subsets of $X$ satisfying $U\subseteq\overline{U}\subseteq V$. We say that $\varphi\in\mathcal{F}$ is a cutoff function for $U\subseteq V$ if $0\leq\varphi\leq 1$ $m$-a.e., $\varphi=1$ $m$-a.e. in an open neighborhood of $\overline{U}$ and $\mathrm{supp}(\varphi)\subseteq V$, where $\mathrm{supp}(f)$ refers to the support of the measure of $|f|\mathrm{d}m$ for any given function $f$.
We say that the cutoff Sobolev inequality CS($\Psi$) holds if there exists $C_{S}\in(0,+\infty)$ such that for any $x\in X$, for any $R,r\in(0,+\infty)$, there exists a cutoff function $\varphi\in\mathcal{F}$ for $B(x,R)\subseteq B(x,R+r)$ such that for any $f\in\mathcal{F}$, we have
$\int_{B(x,R+r)\backslash\overline{B(x,R)}}f^{2}\mathrm{d}\Gamma(\varphi,\varphi)$
$\leq\frac{1}{8}\int_{B(x,R+r)\backslash\overline{B(x,R)}}\varphi^{2}\mathrm{d}\Gamma(f,f)+\frac{C_{S}}{\Psi(r)}\int_{B(x,R+r)\backslash\overline{B(x,R)}}f^{2}\mathrm{d}m.$ CS($\Psi$)
We have the characterisations of heat kernel estimates as follows.
Proposition 3.1.
([2, Theorem 1.12]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD. Then the followings are equivalent.
  1. (1)
  2. (2)
Proposition 3.2.
([31, THEOREM 1.2]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$). Then the followings are equivalent.
  1. (1)
  2. (2)
Remark 3.3.
On any complete non-compact Riemannian manifold, CS($\Psi$) with $\beta_{1}=\beta_{2}=2$, or equivalently, $\Psi(r)=r^{2}$ for any $r\in(0,+\infty)$, holds automatically, for example, one can simply take
$\varphi=1-\left(\frac{d(x,\cdot)-(R+\frac{r}{3})}{\frac{r}{3}}\vee 0\right)\wedge 1,$
so that the above equivalences hold without CS($\Psi$) and are classical, see [27, 50, 28]. However, on a general MMD space, CS($\Psi$) does not always hold and is involved in the formulation of the previous equivalences.
4 Poisson equation on unbounded MMD spaces
[…]
First, we have the existence, the uniqueness and the regularity for the solutions of Poisson equation as follows.
Lemma 4.1.
([25, Lemma 2.2, Lemma 2.6]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying FK($\Psi,\nu$). Let $q=\frac{2}{1-\nu}$. Then for any $p\in\left[\frac{q}{q-1},+\infty\right)$, for any ball $B=B(x_{0},r)$, for any $f\in L^{p}(B)$, there exists a unique 222in the sense that if $u_{1},u_{2}\in\mathcal{F}_{B}$ satisfy $\Delta u_{1}=\Delta u_{2}=f$ in $B$, then $u_{1}=u_{2}$ $m$-a.e.. $u\in\mathcal{F}_{B}$ such that $\Delta u=f$ in $B$. There exists $C\in(0,+\infty)$ depending only on $\nu$, $C_{F}$ such that
[…]
Second, we have the pointwise estimate for the solutions of Poisson equation as follows.
Lemma 4.2.
([25, Lemma 2.7]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD, FK($\Psi,\nu$) and CS($\Psi$). Let $q=\frac{2}{1-\nu}$. Then for any $p\in\left[\frac{q}{q-1},+\infty\right)$, there exists $C\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $f\in L^{\infty}(2B)$, if $u\in\mathcal{F}$ satisfies $\Delta u=f$ in $2B$, then for $m$-a.e. $x\in B$, we have
$|u(x)|\leq C\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+F_{1}(x)\right),$
where
$F_{1}(x)=\sum_{j\leq[\log_{2}r]}\Psi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
The proof of the above two results in [25] was given for the case $\beta_{1}=2$. It is easy to see that the same proof holds for more general $\Psi$, including the case in our paper.
Third, we have Hölder estimate for the solutions of Poisson equation as follows. The proof is similar to [25, Proposition 5.1] and [22, Theorem 3.2] to do gradient estimate. Some technical modifications will be needed.
Proposition 4.3.
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$),
FK($\Psi,\nu$), CS($\Psi$) and HR($\Phi,\Psi$). Let $q=\frac{2}{1-\nu}$. Then for any $p\in\left[\frac{q}{q-1},+\infty\right)$, there exists $C\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $f\in L^{\infty}(2B)$, if $u\in\mathcal{F}$ satisfies $\Delta u=f$ in $2B$, then for $m$-a.e. $x,y\in\frac{1}{16}B$ with $x\neq y$, we have
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u(x)-u(y)|\leq C\left(\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+F_{2}(x)+F_{2}(y)\right),$
where
$F_{2}(x)=\sum_{j\leq[\log_{2}r]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
Proof.
Let $k_{0}=[\log_{2}d(x,y)]$ and $k_{1}=[\log_{2}r]$, then $2^{k_{0}}\leq d(x,y)<2^{k_{0}+1}$ and $2^{k_{1}}\leq r<2^{k_{1}+1}$. Since $x,y\in\frac{1}{16}B$, we have $d(x,y)<\frac{1}{8}r$. Hence $2^{k_{0}}\leq d(x,y)<\frac{1}{8}r<\frac{1}{8}2^{k_{1}+1}=2^{k_{1}-2}$, which implies $k_{0}+3\leq k_{1}$.
For any $k=k_{0}+3,\ldots,k_{1}$, by Lemma 4.1, there exists a unique $u_{k}\in\mathcal{F}_{B(x,2^{k})}$ such that $\Delta u_{k}=f$ in $B(x,2^{k})$ and
[…]
Then
$|u(x)-u(y)|$
$\leq|(u-u_{k_{1}})(x)-(u-u_{k_{1}})(y)|$
$+\sum_{k=k_{0}+4}^{k_{1}}|(u_{k}-u_{k-1})(x)-(u_{k}-u_{k-1})(y)|$
$+|u_{k_{0}+3}(x)|+|u_{k_{0}+3}(y)|.$
For any $k=k_{0}+3,\ldots,k_{1}$, we have $d(x,y)<2^{k_{0}+1}\leq 2^{k-2}$, that is, $y\in B(x,2^{k-2})$.
Since $\Delta(u-u_{k_{1}})=0$ in $B(x,2^{k_{1}})$, by HR($\Phi,\Psi$), we have
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|(u-u_{k_{1}})(x)-(u-u_{k_{1}})(y)|$
$\lesssim\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|u-u_{k_{1}}|\mathrm{d}m$
$\leq\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|u|\mathrm{d}m+\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|u_{k_{1}}|\mathrm{d}m\right)$
$\lesssim\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+\Psi\left(2^{k_{1}}\right)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+\Phi(2^{k_{1}})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
Similarly, for any $k=k_{0}+4,\ldots,k_{1}$, since $\Delta(u_{k}-u_{k-1})=0$ in $B(x,2^{k-1})$, by HR($\Phi,\Psi$), we have
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|(u_{k}-u_{k-1})(x)-(u_{k}-u_{k-1})(y)|$
$\lesssim\frac{\Phi(2^{k-2})}{\Psi(2^{k-2})}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k-1})}|u_{k}-u_{k-1}|\mathrm{d}m$
$\leq\frac{\Phi(2^{k-2})}{\Psi(2^{k-2})}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k-1})}|u_{k}|\mathrm{d}m+\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k-1})}|u_{k-1}|\mathrm{d}m\right)$
$\lesssim\frac{\Phi(2^{k-2})}{\Psi(2^{k-2})}\left(\Psi\left(2^{k}\right)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\Psi\left(2^{k-1}\right)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k-1})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
$\lesssim\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\Phi(2^{k-1})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k-1})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
Since $\Delta u_{k_{0}+3}=f$ in $B(x,2^{k_{0}+3})$, by Lemma 4.2, we have
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u_{k_{0}+3}(x)|$
$\lesssim\frac{\Phi(d(x,y))}{\Psi(d(x,y))}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|u_{k_{0}+3}|\mathrm{d}m+\sum_{j\leq k_{0}+2}\Psi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
$\lesssim\frac{\Phi(d(x,y))}{\Psi(d(x,y))}\left(\Psi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Psi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
$\lesssim\frac{\Phi(2^{k_{0}+3})}{\Psi(2^{k_{0}+3})}\Psi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}$
$+\sum_{j\leq k_{0}+2}\frac{\Phi(2^{k_{0}+2})}{\Psi(2^{k_{0}+2})}\frac{\Psi(2^{j})}{\Phi(2^{j})}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}$
$\leq\Phi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p},$
where in the last inequality, we use the fact that $\gamma_{1},\gamma_{2}>0$, which implies that
$\frac{\Phi(2^{k_{0}+2})}{\Psi(2^{k_{0}+2})}\frac{\Psi(2^{j})}{\Phi(2^{j})}\leq 1\text{ for any }j\leq k_{0}+2.$
Similarly, since $y\in B(x,2^{k_{0}+1})\subseteq B(x,2^{k_{0}+2})$, by Lemma 4.2, we also have
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u_{k_{0}+3}(y)|$
$\lesssim\frac{\Phi(d(x,y))}{\Psi(d(x,y))}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|u_{k_{0}+3}|\mathrm{d}m+\sum_{j\leq k_{0}+2}\Psi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
$\lesssim\Phi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
In summary, we have
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u(x)-u(y)|$
$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+\Phi(2^{k_{1}})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|f|^{p}\mathrm{d}m\right)^{1/p}$
$+\sum_{k=k_{0}+4}^{k_{1}}\left(\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\Phi(2^{k-1})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k-1})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
$+\Phi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}$
$+\Phi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}$
$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+\sum_{k=k_{0}+3}^{k_{1}}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}$
$+\sum_{k\leq k_{0}+3}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{k\leq k_{0}+2}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}$
$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+\sum_{k\leq k_{1}}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{k\leq k_{1}}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
[…]
The idea is to consider the heat equation “$\Delta(P_{t}f)=\frac{\partial}{\partial t}(P_{t}f)$” for fixed $t$ as a Poisson equation.
[…]
For any $t\in(0,+\infty)$, for any $f\in L^{\infty}(X;m)$, we have $P_{t}f$ is a solution of the Poisson equation $\Delta(P_{t}f)=\frac{\partial}{\partial t}(P_{t}f)$. For any $x,y\in X$ with $x\neq y$. If $d(x,y)<\Psi^{-1}(t)$, then taking $r=16\Psi^{-1}(t)>16d(x,y)$, by Proposition 4.3, we have
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|P_{t}f(x)-P_{t}f(y)|$
$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,32\Psi^{-1}(t))}|P_{t}f|\mathrm{d}m$
$+\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}$
$+\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}.$
For the first term, we have
$\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,32\Psi^{-1}(t))}|P_{t}f|\mathrm{d}m$
$\lesssim\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,32\Psi^{-1}(t))}\lVert P_{t}f\rVert_{L^{\infty}(X;m)}\mathrm{d}m$
$\leq\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\lVert f\rVert_{L^{\infty}(X;m)}.$
For the second term, by UHK($\Psi$) and [23, THEOREM 4], we have the following estimate of the time derivative of the heat kernel
$\left|\frac{\partial}{\partial t}p_{t}(z,w)\right|\leq\frac{C_{1}}{tV\left(z,\Psi^{-1}(t)\right)}\exp\left(-\Upsilon\left(C_{2}d(z,w),t\right)\right),$
hence
$|\frac{\partial}{\partial t}(P_{t}f)(z)|$
$\leq\int_{X}\lvert\frac{\partial}{\partial t}p_{t}(z,w)\rvert\cdot|f(w)|m(\mathrm{d}w)$
$\leq\lVert f\rVert_{L^{\infty}(X;m)}\int_{X}\frac{C_{1}}{tV\left(z,\Psi^{-1}(t)\right)}\exp\left(-\Upsilon\left(C_{2}d(z,w),t\right)\right)m(\mathrm{d}w).$
By writing
$\int_{X}=\int_{B(z,\Psi^{-1}(t))}+\sum_{n=0}^{\infty}\int_{B(z,2^{n+1}\Psi^{-1}(t))\backslash B(z,2^{n}\Psi^{-1}(t))},$
and using V($\Phi$), we have the integral
$\int_{X}\frac{1}{V\left(z,\Psi^{-1}(t)\right)}\exp\left(-\Upsilon\left(C_{2}d(z,w),t\right)\right)m(\mathrm{d}w)$
is bounded by some constant depending only on $C_{2}$ and $\Phi$, see also [3, Page 944, Line -5], [20, Equation (2.5)], hence
$|\frac{\partial}{\partial t}(P_{t}f)(z)|\lesssim\frac{\lVert f\rVert_{L^{\infty}(X;m)}}{t}.$
Hence
$\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}$
$\lesssim\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}\left(\frac{\lVert f\rVert_{L^{\infty}(X;m)}}{t}\right)^{p}\mathrm{d}m\right)^{1/p}$
$=\frac{\lVert f\rVert_{L^{\infty}(X;m)}}{t}\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\frac{\Phi(2^{k})}{\Phi(16\Psi^{-1}(t))}\frac{\Phi(16\Psi^{-1}(t))}{\Phi(\Psi^{-1}(t))}\Phi(\Psi^{-1}(t))$
$\lesssim\frac{\lVert f\rVert_{L^{\infty}(X;m)}}{t}\Phi(\Psi^{-1}(t))$
$=\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\lVert f\rVert_{L^{\infty}(X;m)}.$
For the third term, similarly, we have
$\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}\lesssim\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\lVert f\rVert_{L^{\infty}(X;m)}.$
In summary, we have
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|P_{t}f(x)-P_{t}f(y)|\lesssim\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\lVert f\rVert_{L^{\infty}(X;m)}.$
If $d(x,y)\geq\Psi^{-1}(t)$, then obviously
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|P_{t}f(x)-P_{t}f(y)|$
$\leq 2\frac{\Phi(d(x,y))}{\Psi(d(x,y))}\lVert P_{t}f\rVert_{L^{\infty}(X;m)}\leq 2\frac{\Phi(d(x,y))}{\Psi(d(x,y))}\lVert f\rVert_{L^{\infty}(X;m)}$
$=2\frac{\frac{\Psi(\Psi^{-1}(t))}{\Phi(\Psi^{-1}(t))}}{\frac{\Psi(d(x,y))}{\Phi(d(x,y))}}\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\lVert f\rVert_{L^{\infty}(X;m)}\leq 2\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\lVert f\rVert_{L^{\infty}(X;m)}.$
[…]
By UHK($\Psi$) and [23, THEOREM 4], we have the following estimates of the heat kernel
$p_{t}(x,y)\leq\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,y),t)\right),$
and its time derivative
$\left|\frac{\partial}{\partial t}p_{t}(z,w)\right|\leq\frac{C_{1}}{tV\left(z,\Psi^{-1}(t)\right)}\exp\left(-\Upsilon\left(C_{2}d(z,w),t\right)\right).$
Take any $t\in(0,+\infty)$, $x,y_{1},y_{2}\in X$. We now distinguish the following four cases.
  1. (i)
    $d(y_{1},y_{2})\geq\Psi^{-1}(t)$.
  2. (ii)
    $d(y_{1},y_{2})<\Psi^{-1}(t)$, $d(x,y_{1})<128\Psi^{-1}(t)$ and $d(x,y_{2})<128\Psi^{-1}(t)$.
  3. (iii)
    $d(y_{1},y_{2})<\Psi^{-1}(t)$ and $d(x,y_{1})\geq 128\Psi^{-1}(t)$.
  4. (iv)
    $d(y_{1},y_{2})<\Psi^{-1}(t)$ and $d(x,y_{2})\geq 128\Psi^{-1}(t)$.
For the case (i), we obtain the result directly from UHK($\Psi$) as follows.
$|p_{t}(x,y_{1})-p_{t}(x,y_{2})|\leq p_{t}(x,y_{1})+p_{t}(x,y_{2})$
$\leq\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,y_{1}),t)\right)+\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,y_{2}),t)\right)$
$\asymp\frac{1}{\Phi(\Psi^{-1}(t))}\left(\exp\left(-\Upsilon(C_{2}d(x,y_{1}),t)\right)+\exp\left(-\Upsilon(C_{2}d(x,y_{2}),t)\right)\right)$
$=\frac{\Psi(\Psi^{-1}(t))}{\Phi(\Psi^{-1}(t))}\frac{1}{t}\left(\exp\left(-\Upsilon(C_{2}d(x,y_{1}),t)\right)+\exp\left(-\Upsilon(C_{2}d(x,y_{2}),t)\right)\right)$
$\leq\frac{\left(\frac{\Psi}{\Phi}\right)(d(y_{1},y_{2}))}{t}\left(\exp\left(-\Upsilon(C_{2}d(x,y_{1}),t)\right)+\exp\left(-\Upsilon(C_{2}d(x,y_{2}),t)\right)\right).$
For the remaining three cases, by Proposition 4.3, letting $r=16\Psi^{-1}(t)>16d(y_{1},y_{2})$, we have
$\frac{\Phi(d(y_{1},y_{2}))}{\Psi(d(y_{1},y_{2}))}|p_{t}(x,y_{1})-p_{t}(x,y_{2})|$
$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}p_{t}(x,z)m(\mathrm{d}z)$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},2^{j})}\left|\frac{\partial}{\partial t}p_{t}(x,z)\right|^{p}m(\mathrm{d}z)\right)^{1/p}$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{2},2^{j})}\left|\frac{\partial}{\partial t}p_{t}(x,z)\right|^{p}m(\mathrm{d}z)\right)^{1/p}.$ (6.1)
For the case (ii), we have $d(x,y_{1})<128\Psi^{-1}(t)$ and $d(x,y_{2})<128\Psi^{-1}(t)$. By Equation (6.1), we have
$\frac{\Phi(d(y_{1},y_{2}))}{\Psi(d(y_{1},y_{2}))}|p_{t}(x,y_{1})-p_{t}(x,y_{2})|$
$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}m(\mathrm{d}z)$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
$=\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}+2\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\frac{C_{1}}{tV(x,\Psi^{-1}(t))}$
$\asymp\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\frac{1}{\Phi(\Psi^{-1}(t))}+\frac{1}{t\Phi(\Psi^{-1}(t))}\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})$
$=\frac{1}{t}+\frac{1}{t\Phi(\Psi^{-1}(t))}\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j}).$
For any $r\in(0,+\infty)$. If $r<1$, then
$\sum_{j\leq[\log_{2}r]}\Phi(2^{j})=\sum_{j\leq[\log_{2}r]}(2^{j})^{\alpha_{1}}\asymp 2^{[\log_{2}r]\alpha_{1}}\asymp r^{\alpha_{1}}=\Phi(r).$
If $r\geq 1$, then
$\sum_{j\leq[\log_{2}r]}\Phi(2^{j})=\sum_{j\leq-1}(2^{j})^{\alpha_{1}}+\sum_{j=0}^{[\log_{2}r]}(2^{j})^{\alpha_{2}}\asymp 1+2^{[\log_{2}r]\alpha_{2}}\asymp 1+r^{\alpha_{2}}\asymp r^{\alpha_{2}}=\Phi(r).$
Hence
$\sum_{j\leq[\log_{2}r]}\Phi(2^{j})\asymp\Phi(r)\text{ for any }r\in(0,+\infty).$
Then
$\frac{\Phi(d(y_{1},y_{2}))}{\Psi(d(y_{1},y_{2}))}|p_{t}(x,y_{1})-p_{t}(x,y_{2})|\lesssim\frac{1}{t}+\frac{1}{t\Phi(\Psi^{-1}(t))}\Phi(16\Psi^{-1}(t))$
$\asymp\frac{1}{t}+\frac{1}{t\Phi(\Psi^{-1}(t))}\Phi(\Psi^{-1}(t))=\frac{2}{t}.$
By assumption, for $i=1,2$, we have
$\Upsilon(C_{2}d(x,y_{i}),t)=\sup_{s\in(0,+\infty)}\left(\frac{C_{2}d(x,y_{i})}{\Psi^{-1}(s)}-\frac{t}{s}\right)\leq\sup_{s\in(0,+\infty)}\left(\frac{128C_{2}\Psi^{-1}(t)}{\Psi^{-1}(s)}-\frac{t}{s}\right).$
By the following elementary result, Lemma 6.1, there exists $C_{3}\in(0,+\infty)$ depending only on $C_{2}$, $\beta_{1}$, $\beta_{2}$ such that
$\sup_{t,s\in(0,+\infty)}\left(\frac{128C_{2}\Psi^{-1}(t)}{\Psi^{-1}(s)}-\frac{t}{s}\right)\leq C_{3}.$
Hence
$\frac{\Phi(d(y_{1},y_{2}))}{\Psi(d(y_{1},y_{2}))}|p_{t}(x,y_{1})-p_{t}(x,y_{2})|\lesssim\frac{e^{C_{3}}}{t}(2e^{-C_{3}})$
$\leq\frac{e^{C_{3}}}{t}\left(\exp(-\Upsilon(C_{2}d(x,y_{1}),t))+\exp(-\Upsilon(C_{2}d(x,y_{2}),t))\right).$
For the case (iii), we have $d(x,y_{1})\geq 128\Psi^{-1}(t)$. Then in Equation (6.1), for any $z\in B(y_{1},32\Psi^{-1}(t))$, we have
$d(x,z)\geq d(x,y_{1})-d(y_{1},z)\geq d(x,y_{1})-32\Psi^{-1}(t)$
$\geq d(x,y_{1})-\frac{1}{4}d(x,y_{1})=\frac{3}{4}d(x,y_{1})\geq\frac{1}{2}d(x,y_{1}).$
For any $j\leq[\log_{2}16\Psi^{-1}(t)]$, for any $z\in B(y_{1},2^{j})$, we have
$d(x,z)\geq d(x,y_{1})-d(y_{1},z)\geq d(x,y_{1})-16\Psi^{-1}(t)$
$\geq d(x,y_{1})-\frac{1}{8}d(x,y_{1})=\frac{7}{8}d(x,y_{1})\geq\frac{1}{2}d(x,y_{1}),$
for any $z\in B(y_{2},2^{j})$, we have
$d(x,z)\geq d(x,y_{1})-d(y_{1},y_{2})-d(y_{2},z)\geq d(x,y_{1})-\Psi^{-1}(t)-16\Psi^{-1}(t)$
$\geq d(x,y_{1})-\frac{1}{4}d(x,y_{1})=\frac{3}{4}d(x,y_{1})\geq\frac{1}{2}d(x,y_{1}).$
Hence Equation (6.1) gives
$\frac{\Phi(d(y_{1},y_{2}))}{\Psi(d(y_{1},y_{2}))}|p_{t}(x,y_{1})-p_{t}(x,y_{2})|$
$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)m(\mathrm{d}z)$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
$\leq\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)m(\mathrm{d}z)$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
$=\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)$
$+2\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)$
$\asymp\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\frac{1}{\Phi(\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)$
$+\frac{1}{t\Phi(\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})$
$=\frac{1}{t}\left(1+\frac{1}{\Phi(\Psi^{-1}(t))}\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\right)\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)$
$\asymp\frac{1}{t}\left(1+\frac{1}{\Phi(\Psi^{-1}(t))}\Phi(16\Psi^{-1}(t))\right)\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)$
$\asymp\frac{1}{t}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right).$
For the case (iv), similar to the case (iii), we also have
$\frac{\Phi(d(y_{1},y_{2}))}{\Psi(d(y_{1},y_{2}))}|p_{t}(x,y_{1})-p_{t}(x,y_{2})|\lesssim\frac{1}{t}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{2}),t)\right).$
In summary, for any $t\in(0,+\infty)$, $x,y_{1},y_{2}\in X$, we have
$|p_{t}(x,y_{1})-p_{t}(x,y_{2})|$
$\lesssim\frac{1}{t}\frac{\Psi(d(y_{1},y_{2}))}{\Phi(d(y_{1},y_{2}))}\left(\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)+\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{2}),t)\right)\right).$
Lemma 6.1.
([25, Lemma 5.2]) Let $A\in(0,+\infty)$. Then there exists $C\in(0,+\infty)$ depending only on $A$, $\beta_{1}$, $\beta_{2}$ such that
$\sup_{t,s\in(0,+\infty)}\left(A\frac{\Psi^{-1}(t)}{\Psi^{-1}(s)}-\frac{t}{s}\right)\leq C.$
Proof.
Let $f:(0,+\infty)\times(0,+\infty)\to\mathbb{R}$ be given by
$f(t,s)=A\frac{\Psi^{-1}(t)}{\Psi^{-1}(s)}-\frac{t}{s}.$
By considering separately the cases $t,s\in(0,1)$, $t,s\in[1,+\infty)$, $0<t<1\leq s$ and $0<s<1\leq t$, we easily obtain that
$f(t,s)\leq A\max\left\{\left(\frac{t}{s}\right)^{1/\beta_{1}},\left(\frac{t}{s}\right)^{1/\beta_{2}}\right\}-\frac{t}{s}.$
Since the function $(0,+\infty)\to\mathbb{R}$, $x\mapsto A\max\{x^{1/\beta_{1}},x^{1/\beta_{2}}\}-x$ is bounded from above by some positive constant $C$ depending only on $A$, $\beta_{1}$, $\beta_{2}$, we have
$\sup_{t,s\in(0,+\infty)}f(t,s)\leq C.$
[…]
For any $t\in(0,+\infty)$, for any $x,y_{1},y_{2}\in X$, we have
$p_{t}(x,y_{1})=p_{t}(y_{1},x)=\int_{X}p_{t/2}(y_{1},z)p_{t/2}(z,x)m(\mathrm{d}z)$
$=\int_{X}p_{t/2}(y_{1},z)p_{t/2}(x,z)m(\mathrm{d}z)=P_{t/2}(p_{t/2}(x,\cdot))(y_{1}),$
and $p_{t}(x,y_{2})=P_{t/2}(p_{t/2}(x,\cdot))(y_{2})$. By wBE($\Phi,\Psi$) and UHK($\Psi$), we have
$\frac{\Phi(d(y_{1},y_{2}))}{\Psi(d(y_{1},y_{2}))}|p_{t}(x,y_{1})-p_{t}(x,y_{2})|$
$=\frac{\Phi(d(y_{1},y_{2}))}{\Psi(d(y_{1},y_{2}))}|P_{t/2}(p_{t/2}(x,\cdot))(y_{1})-P_{t/2}(p_{t/2}(x,\cdot))(y_{2})|$
$\leq C_{wBE}\frac{\Phi(\Psi^{-1}(\frac{t}{2}))}{\Psi(\Psi^{-1}(\frac{t}{2}))}\lVert p_{t/2}(x,\cdot)\rVert_{L^{\infty}(X;m)}$
$\leq C_{wBE}\frac{\Phi(\Psi^{-1}(\frac{t}{2}))}{\frac{t}{2}}\frac{C_{1}}{V(x,\Psi^{-1}(t/2))}$
$\leq 2C_{wBE}\frac{\Phi(\Psi^{-1}(\frac{t}{2}))}{t}\frac{C_{1}C_{VR}}{\Phi(\Psi^{-1}(t/2))}$
$=\left(2C_{1}C_{VR}C_{wBE}\right)\frac{1}{t},$
that is,
$|p_{t}(x,y_{1})-p_{t}(x,y_{2})|\lesssim\frac{1}{t}\frac{\Psi(d(y_{1},y_{2}))}{\Phi(d(y_{1},y_{2}))}.$
Hence for any $x_{1},x_{2}\in X$, we have
$|p_{t}(x_{1},y_{1})-p_{t}(x_{2},y_{2})|$
$\leq|p_{t}(x_{1},y_{1})-p_{t}(x_{1},y_{2})|+|p_{t}(x_{1},y_{2})-p_{t}(x_{2},y_{2})|$
$\lesssim\frac{\left(\frac{\Psi}{\Phi}\right)(d(y_{1},y_{2}))+\left(\frac{\Psi}{\Phi}\right)(d(x_{1},x_{2}))}{t}.$
[…]
First, we give a lower bound of $p_{t}(x,x)$. For any $N\geq 0$, by UHK($\Psi$), we have
$\int_{X\backslash B(x,2^{N}\Psi^{-1}(t))}p_{t}(x,y)m(\mathrm{d}y)$
$=\sum_{n=N}^{\infty}\int_{B(x,2^{n+1}\Psi^{-1}(t))\backslash B(x,2^{n}\Psi^{-1}(t))}p_{t}(x,y)m(\mathrm{d}y)$
$\leq\sum_{n=N}^{\infty}\int_{B(x,2^{n+1}\Psi^{-1}(t))\backslash B(x,2^{n}\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,y),t)\right)m(\mathrm{d}y)$
$\leq\sum_{n=N}^{\infty}C_{1}\frac{V(x,2^{n+1}\Psi^{-1}(t))}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}2^{n}\Psi^{-1}(t),t)\right).$
By V($\Phi$), we have
$\frac{V(x,2^{n+1}\Psi^{-1}(t))}{V(x,\Psi^{-1}(t))}\leq C_{VR}^{2}(2^{n+1})^{\alpha_{1}\vee\alpha_{2}}.$
Note that
$\Upsilon(C_{2}2^{n}\Psi^{-1}(t),t)=\sup_{s\in(0,+\infty)}\left(\frac{C_{2}2^{n}\Psi^{-1}(t)}{s}-\frac{t}{\Psi(s)}\right)$
$=\sup_{s\in(0,+\infty)}\left(\frac{C_{2}2^{n}\Psi^{-1}(t)}{\Psi^{-1}(s)}-\frac{t}{s}\right)\geq C_{2}2^{n}-1.$
Hence
$\int_{X\backslash B(x,2^{N}\Psi^{-1}(t))}p_{t}(x,y)m(\mathrm{d}y)\leq C_{1}C_{VR}^{2}\sum_{n=N}^{\infty}(2^{n+1})^{\alpha_{1}\vee\alpha_{2}}\exp(1-C_{2}2^{n}).$
Since
$\sum_{n=0}^{\infty}(2^{n+1})^{\alpha_{1}\vee\alpha_{2}}\exp(1-C_{2}2^{n})<+\infty,$
there exists $N\geq 0$ depending only on $C_{1}$, $C_{2}$, $C_{VR}$, $\alpha_{1}$, $\alpha_{2}$ sufficiently large such that
$\int_{X\backslash B(x,2^{N}\Psi^{-1}(t))}p_{t}(x,y)m(\mathrm{d}y)\leq C_{1}C_{VR}^{2}\sum_{n=N}^{\infty}(2^{n+1})^{\alpha_{1}\vee\alpha_{2}}\exp(1-C_{2}2^{n})\leq\frac{1}{2}.$
Since $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is conservative, that is, $\int_{X}p_{t}(x,y)m(\mathrm{d}y)=1$, we have
$\int_{B(x,2^{N}\Psi^{-1}(t))}p_{t}(x,y)m(\mathrm{d}y)=1-\int_{X\backslash B(x,2^{N}\Psi^{-1}(t))}p_{t}(x,y)m(\mathrm{d}y)\geq\frac{1}{2}.$
On the other hand
$\int_{B(x,2^{N}\Psi^{-1}(t))}p_{t}(x,y)m(\mathrm{d}y)$
$\leq\left(\int_{B(x,2^{N}\Psi^{-1}(t))}p_{t}(x,y)^{2}m(\mathrm{d}y)\right)^{1/2}V(x,2^{N}\Psi^{-1}(t))^{1/2}$
$\leq\left(\int_{X}p_{t}(x,y)^{2}m(\mathrm{d}y)\right)^{1/2}V(x,2^{N}\Psi^{-1}(t))^{1/2}$
$=p_{2t}(x,x)^{1/2}V(x,2^{N}\Psi^{-1}(t))^{1/2},$
where we use the Cauchy-Schwarz inequality in the first inequality and the semi-group property in the last equality, hence
$p_{2t}(x,x)\geq\frac{1}{4V(x,2^{N}\Psi^{-1}(t))},$
hence
$p_{t}(x,x)\geq\frac{1}{4V(x,2^{N}\Psi^{-1}(\frac{t}{2}))}\geq\frac{\widetilde{C}_{1}}{\Phi(\Psi^{-1}(t))},$
where $\widetilde{C}_{1}$ is a positive constant depending only on $C_{1}$, $C_{2}$, $C_{VR}$, $N$, $\Phi$, $\Psi$.
Second, we prove NLE($\Psi$). Indeed, for any $\varepsilon\in(0,\frac{1}{2})$, for any $y\in B(x,\varepsilon\Psi^{-1}(t))$, by HHK($\Phi,\Psi$), we have
$|p_{t}(x,x)-p_{t}(x,y)|$
$\leq C_{HHK}\frac{\left(\frac{\Psi}{\Phi}\right)(d(x,y))}{t}$
$\leq C_{HHK}\frac{\left(\frac{\Psi}{\Phi}\right)(\varepsilon\Psi^{-1}(t))}{\left(\frac{\Psi}{\Phi}\right)(\Psi^{-1}(t))}\frac{\left(\frac{\Psi}{\Phi}\right)(\Psi^{-1}(t))}{t}$
$\leq C_{HHK}\varepsilon^{\gamma_{1}\wedge\gamma_{2}}\frac{1}{\Phi(\Psi^{-1}(t))}.$
There exists $\varepsilon\in(0,\frac{1}{2})$ depending only on $\widetilde{C}_{1}$, $C_{HHK}$, $\gamma_{1}$, $\gamma_{2}$ sufficiently small such that
$C_{HHK}\varepsilon^{\gamma_{1}\wedge\gamma_{2}}\leq\frac{\widetilde{C}_{1}}{2},$
hence
$|p_{t}(x,x)-p_{t}(x,y)|\leq\frac{\widetilde{C}_{1}}{2\Phi(\Psi^{-1}(t))}\leq\frac{1}{2}p_{t}(x,x).$
Thus we have
$p_{t}(x,y)\geq p_{t}(x,x)-|p_{t}(x,x)-p_{t}(x,y)|\geq p_{t}(x,x)-\frac{1}{2}p_{t}(x,x)=\frac{1}{2}p_{t}(x,x)\geq\frac{\widetilde{C}_{1}}{2\Phi(\Psi^{-1}(t))}$
for any $y\in B(x,\varepsilon\Psi^{-1}(t))$, that is, we have NLE($\Psi$). ∎
[…]
Recall that a Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is recurrent if for any non-negative function $f\in L^{1}(X;m)$, we have either $Gf=+\infty$ or $Gf=0$ $m$-a.e., where $G$ is the Green operator, see [26, Equation (1.6.2)]. In our setting, the Dirichlet form is recurrent as follows.
Lemma 8.1.
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$) and NLE($\Psi$). Then $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is recurrent.
Proof.
Taking any non-zero non-negative function $f\in L^{1}(X;m)$, we show that
$Gf(x):=\int_{0}^{+\infty}\left(\int_{X}p_{t}(x,y)f(y)m(\mathrm{d}y)\right)\mathrm{d}t=+\infty\text{ for }m\text{-a.e. }x\in X.$
Indeed, by V($\Phi$) and NLE($\Psi$), we have
$\int_{0}^{+\infty}p_{t}(x,y)\mathrm{d}t\geq\int_{\Psi(\frac{d(x,y)}{\varepsilon})}^{+\infty}p_{t}(x,y)\mathrm{d}t\geq\int_{\Psi(\frac{d(x,y)}{\varepsilon})}^{+\infty}\frac{C}{V(x,\Psi^{-1}(t))}\mathrm{d}t$
$\asymp\int_{\Psi(\frac{d(x,y)}{\varepsilon})}^{+\infty}\frac{1}{\Phi(\Psi^{-1}(t))}\mathrm{d}t\geq\int_{\Psi(\frac{d(x,y)}{\varepsilon})\vee 1}^{+\infty}\frac{\mathrm{d}t}{t^{\alpha_{2}/\beta_{2}}}=+\infty.$
Since $f$ is non-zero non-negative, there exists $a>0$ such that $m(\{f>a\})>0$, then
$Gf(x)=\int_{X}\left(\int_{0}^{+\infty}p_{t}(x,y)\mathrm{d}t\right)f(y)m(\mathrm{d}y)\geq a\int_{\{f>a\}}\left(\int_{0}^{+\infty}p_{t}(x,y)\mathrm{d}t\right)m(\mathrm{d}y)=+\infty$
for $m$-a.e. $x\in X$. Hence $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is recurrent. ∎
We use some techniques from resistance forms. The theory of resistance forms was developed by Kigami [44]. To begin with, let us introduce effective resistances. Let $A,B$ be two non-empty subsets of $X$. The effective resistance between $A$ and $B$ is defined as follows.
$R(A,B)=\left(\inf\left\{\mathcal{E}(u,u):u\in\mathcal{F},u=0\text{ on }A,u=1\text{ on }B\right\}\right)^{-1},$
here we use the convention that $\inf\emptyset=+\infty$, $0^{-1}=+\infty$ and $(+\infty)^{-1}=0$. For $x,y\in X$, write
$R(x,A)=R(\{x\},A),R(x,y)=R(\{x\},\{y\}).$
By definition, it is obvious that
$R(x,y)=\sup\left\{\frac{|u(x)-u(y)|^{2}}{\mathcal{E}(u,u)}:u\in\mathcal{F},\mathcal{E}(u,u)>0\right\},$
which implies that
$|u(x)-u(y)|^{2}\leq R(x,y)\mathcal{E}(u,u)\text{ for any }u\in\mathcal{F},x,y\in X,$
and if $A_{1}\subseteq A_{2}$, $B_{1}\subseteq B_{2}$, then
$R(A_{1},B_{1})\geq R(A_{2},B_{2}).$
Proposition 8.2.
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$),
UHK($\Psi$) and NLE($\Psi$). Then we have two-sided resistance estimates R($\Phi,\Psi$) as follows. There exists $C_{R}\in(0,+\infty)$ such that for any $x,y\in X$ with $x\neq y$, we have
$\frac{1}{C_{R}}\frac{\Psi(d(x,y))}{\Phi(d(x,y))}\leq R(x,y)\leq C_{R}\frac{\Psi(d(x,y))}{\Phi(d(x,y))}.$ R($\Phi,\Psi$)
Remark 8.3.
In the strongly recurrent setting, two-sided heat kernel estimates and two-sided resistance estimates are indeed equivalent. This equivalence was proved in [12] on graphs, in [45] on graphs and resistance forms and in [37] on metric spaces using an analytical approach.
Under the assumptions of the above result, by Proposition 3.2, we have PI($\Psi$) and CS($\Psi$).
The following result will play an important role in this section.
Lemma 8.4.
(Morrey-Sobolev inequality) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$) and PI($\Psi$). Then we have the following Morrey-Sobolev inequality MS($\Phi,\Psi$). There exists $C_{MS}\in(0,+\infty)$ such that for any $u\in\mathcal{F}$, for $m$-a.e. $x,y\in X$ with $x\neq y$, we have
$|u(x)-u(y)|^{2}\leq C_{MS}\frac{\Psi(d(x,y))}{\Phi(d(x,y))}\mathcal{E}(u,u).$ MS($\Phi,\Psi$)
Hence any function in $\mathcal{F}$ has a continuous version, or equivalently, $\mathcal{F}\subseteq C(X)$.
The proof is standard using telescopic technique, see also [12, Page 1654].
Proof.
Let $x,y$ be two different Lebesgue points of $u\in\mathcal{F}\subseteq L^{2}(X;m)$. Denote $r=d(x,y)$. Then
$|u(x)-u(y)|\leq|u(x)-u_{B(x,r)}|+|u_{B(x,r)}-u_{B(y,r)}|+|u(y)-u_{B(y,r)}|.$
For any $n\in\mathbb{Z}$, denote $B_{n}=B(x,2^{-n}r)$. By V($\Phi$) and PI($\Psi$), we have
$|u_{B_{n+1}}-u_{B_{n}}|\leq\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B_{n+1}}|u-u_{B_{n}}|\mathrm{d}m\lesssim\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B_{n}}|u-u_{B_{n}}|\mathrm{d}m\leq\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B_{n}}|u-u_{B_{n}}|^{2}\mathrm{d}m\right)^{1/2}$
$\lesssim\left(\frac{1}{\Phi(2^{-n}r)}\Psi(2^{-n}r)\int_{B_{n-1}}\mathrm{d}\Gamma(u,u)\right)^{1/2}\leq\left(\frac{\Psi(2^{-n}r)}{\Phi(2^{-n}r)}\mathcal{E}(u,u)\right)^{1/2},$
hence
[…]
$=\sum_{n=0}^{\infty}|u_{B_{n+1}}-u_{B_{n}}|\lesssim\sum_{n=0}^{\infty}\left(\frac{\Psi(2^{-n}r)}{\Phi(2^{-n}r)}\mathcal{E}(u,u)\right)^{1/2}$
$=\left(\frac{\Psi(r)}{\Phi(r)}\mathcal{E}(u,u)\right)^{1/2}\sum_{n=0}^{\infty}\left(\frac{\left(\frac{\Psi}{\Phi}\right)(2^{-n}r)}{\left(\frac{\Psi}{\Phi}\right)(r)}\right)^{1/2}\leq\left(\frac{\Psi(r)}{\Phi(r)}\mathcal{E}(u,u)\right)^{1/2}\sum_{n=0}^{\infty}2^{-\frac{(\gamma_{1}\wedge\gamma_{2})n}{2}}$
$\asymp\left(\frac{\Psi(r)}{\Phi(r)}\mathcal{E}(u,u)\right)^{1/2}.$
Similarly, we have
$|u(y)-u_{B(y,r)}|\lesssim\left(\frac{\Psi(r)}{\Phi(r)}\mathcal{E}(u,u)\right)^{1/2}.$
Note that
$|u_{B(x,r)}-u_{B(y,r)}|\leq\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,r)}|u(z_{1})-u(z_{2})|m(\mathrm{d}z_{1})m(\mathrm{d}z_{2})$
[…]
$\asymp\frac{1}{\Phi(r)}\left(\int_{B(x,r)}\int_{B(y,r)}|u(z_{1})-u(z_{2})|^{2}m(\mathrm{d}z_{1})m(\mathrm{d}z_{2})\right)^{1/2}$
$\leq\frac{1}{\Phi(r)}\left(\int_{B(x,2r)}\int_{B(x,2r)}|u(z_{1})-u(z_{2})|^{2}m(\mathrm{d}z_{1})m(\mathrm{d}z_{2})\right)^{1/2}$
$=\frac{1}{\Phi(r)}\left(2V(x,2r)\int_{B(x,2r)}\left(u-u_{B(x,2r)}\right)^{2}\mathrm{d}m\right)^{1/2}$
$\lesssim\frac{1}{\Phi(r)^{1/2}}\left(\Psi(2r)\int_{B(x,4r)}\mathrm{d}\Gamma(u,u)\right)\lesssim\left(\frac{\Psi(r)}{\Phi(r)}\mathcal{E}(u,u)\right)^{1/2}.$
In summary, we have
$|u(x)-u(y)|\lesssim\left(\frac{\Psi(r)}{\Phi(r)}\mathcal{E}(u,u)\right)^{1/2}.$
Proof of Proposition 8.2.
The upper bound is obvious by MS($\Phi,\Psi$). We use the idea of an analytical proof of [37, Theorem 3.4] to give the lower bound. By the spectral calculus, it is easy to see that
$\mathcal{E}(P_{t}u,P_{t}u)\leq\frac{1}{2et}(u,u)\text{ for any }t\in(0,+\infty),u\in L^{2}(X;m),$
where $\{P_{t}\}$ is the heat semi-group corresponding to the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$.
For any $t\in(0,+\infty)$, $x\in X$, take $u=p_{t}(x,\cdot)\in L^{2}(X;m)$. By the semi-group property, we have $P_{t}u=p_{2t}(x,\cdot)$ and $(u,u)=p_{2t}(x,x)$, hence
$\mathcal{E}(p_{2t}(x,\cdot),p_{2t}(x,\cdot))\leq\frac{1}{2et}p_{2t}(x,x),$
clearly
$\mathcal{E}(p_{t}(x,\cdot),p_{t}(x,\cdot))\leq\frac{1}{et}p_{t}(x,x).$
For any $y\in X$, we have
$p_{t}(x,x)-p_{t}(x,y)\leq R(x,y)^{1/2}\mathcal{E}(p_{t}(x,\cdot),p_{t}(x,\cdot))^{1/2}\leq R(x,y)^{1/2}\left(\frac{1}{et}p_{t}(x,x)\right)^{1/2}.$ (8.1)
By NLE($\Psi$), we have
$p_{t}(x,x)\geq\frac{C_{1}}{V(x,\Psi^{-1}(t))}.$
By UHK($\Psi$), we have
$p_{t}(x,y)\leq\frac{C_{2}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{3}d(x,y),t)\right),$
and
$p_{t}(x,x)\leq\frac{C_{2}}{V(x,\Psi^{-1}(t))}.$
Hence
$p_{t}(x,x)-p_{t}(x,y)$
$\geq\frac{C_{1}}{V(x,\Psi^{-1}(t))}-\frac{C_{2}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{3}d(x,y),t)\right)$
$=\frac{1}{V(x,\Psi^{-1}(t))}\left(C_{1}-C_{2}\exp\left(-\Upsilon(C_{3}d(x,y),t)\right)\right),$
where
$\Upsilon(C_{3}d(x,y),t)=\sup_{s\in(0,+\infty)}\left(\frac{C_{3}d(x,y)}{s}-\frac{t}{\Psi(s)}\right)$
$=\sup_{s\in(0,+\infty)}\left(\frac{C_{3}d(x,y)}{\Psi^{-1}(s)}-\frac{t}{s}\right)\geq\frac{C_{3}d(x,y)}{\Psi^{-1}(t)}-1.$
Take $\varepsilon\in(0,1)$, let $t=\Psi(\varepsilon d(x,y))$, then
$p_{t}(x,x)-p_{t}(x,y)\geq\frac{1}{V(x,\Psi^{-1}(t))}\left(C_{1}-C_{2}\exp\left(1-\frac{C_{3}}{\varepsilon}\right)\right)\geq\frac{C_{1}}{2V(x,\Psi^{-1}(t))},$ (8.2)
where $\varepsilon\in(0,1)$ depending only on $C_{1}$, $C_{2}$, $C_{3}$ is sufficiently small such that
$C_{2}\exp\left(1-\frac{C_{3}}{\varepsilon}\right)\leq\frac{C_{1}}{2}.$
Plugging Equation (8.2) into Equation (8.1), we have
$\frac{C_{1}}{2V(x,\Psi^{-1}(t))}\leq R(x,y)^{1/2}\left(\frac{1}{et}p_{t}(x,x)\right)^{1/2}\leq R(x,y)^{1/2}\left(\frac{1}{et}\frac{C_{2}}{V(x,\Psi^{-1}(t))}\right)^{1/2},$
therefore
$R(x,y)^{1/2}\geq\frac{C_{1}}{2V(x,\Psi^{-1}(t))}\left(\frac{etV(x,\Psi^{-1}(t))}{C_{2}}\right)^{1/2}$
$\asymp\left(\frac{t}{\Phi(\Psi^{-1}(t))}\right)^{1/2}=\left(\frac{\Psi(\varepsilon d(x,y))}{\Phi(\varepsilon d(x,y))}\right)^{1/2}\asymp\left(\frac{\Psi(d(x,y))}{\Phi(d(x,y))}\right)^{1/2},$
that is,
$R(x,y)\gtrsim\frac{\Psi(d(x,y))}{\Phi(d(x,y))}.$
Let us recall the definition of resistance forms as follows.
Definition 8.5.
([44, DEFINITION 3.1]) Let $X$ be a set. A pair $(\mathcal{E},\mathcal{G})$ is called a resistance form on $X$ if the following five conditions are satisfied.
  1. (RF1)
    $\mathcal{G}$ is a linear subspace of $l(X)$ containing constant functions, where $l(X)$ is the space of all real-valued functions on $X$. $\mathcal{E}$ is a non-negative symmetric quadratic form on $\mathcal{G}$. For $u\in\mathcal{G}$, $\mathcal{E}(u,u)=0$ if and only if $u$ is a constant function.
  2. (RF2)
    Let $\sim$ be the equivalence relation on $\mathcal{G}$ given by $u,v\in\mathcal{G}$, $u\sim v$ if and only if $u-v$ is a constant function. Then $(\mathcal{G}/\sim,\mathcal{E})$ is a Hilbert space.
  3. (RF3)
    For any $x,y\in X$ satisfying $x\neq y$, there exists $u\in\mathcal{G}$ such that $u(x)\neq u(y)$.
  4. (RF4)
    For any $x,y\in X$, we have
    $\sup\left\{\frac{|u(x)-u(y)|^{2}}{\mathcal{E}(u,u)}:u\in\mathcal{G},\mathcal{E}(u,u)>0\right\}$
    is finite, where we use the convention that $\sup\emptyset=0$.
  5. (RF5)
    For any $u\in\mathcal{G}$, let $\overline{u}=(u\vee 0)\wedge 1$, then $\overline{u}\in\mathcal{G}$ and $\mathcal{E}(\overline{u},\overline{u})\leq\mathcal{E}(u,u)$.
Recall that for a Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$, its extended Dirichlet space is defined as the family of all $m$-measurable functions $u$ which is finite $m$-a.e. and there exists an $\mathcal{E}$-Cauchy sequence $\left\{u_{n}\right\}\subseteq\mathcal{F}$ such that $u_{n}\to u$ $m$-a.e.. The sequence $\left\{u_{n}\right\}\subseteq\mathcal{F}$ is called an approximation sequence of $u\in\mathcal{F}_{e}$. We can extend $\mathcal{E}$ to $\mathcal{F}_{e}$ and $\mathcal{F}=\mathcal{F}_{e}\cap L^{2}(X;m)$. See [26, Section 1.5] for more details. We can also define the harmonicity of a function in the extended Dirichlet space as follows. Let $D$ be a non-empty open subset of $X$. We say that $u\in\mathcal{F}_{e}$ is harmonic in $D$ if $\mathcal{E}(u,\varphi)=0$ for any $\varphi\in\mathcal{F}\cap C_{c}(D)$.
We collect some results about extended Dirichlet spaces as follows.
Lemma 8.6.
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$), UHK($\Psi$) and NLE($\Psi$).
  1. (1)
    The Morrey-Sobolev inequality MS($\Phi,\Psi$) holds for any $u\in\mathcal{F}_{e}$, hence $\mathcal{F}_{e}\subseteq C(X)$.
  2. (2)
    $(\mathcal{E},\mathcal{F}_{e})$ is a resistance form on $X$.
  3. (3)
    For any $x,y\in X$, we have
    $R(x,y)=\sup\left\{\frac{|u(x)-u(y)|^{2}}{\mathcal{E}(u,u)}:u\in\mathcal{F}_{e},\mathcal{E}(u,u)>0\right\}.$
  4. (4)
    Let $u\in\mathcal{F}_{e}$ be harmonic in a non-empty open subset $D$ of $X$. Then $\mathcal{E}(u,\varphi)=0$ for any $\varphi\in\mathcal{F}_{e}\subseteq C(X)$ with $\varphi=0$ on $X\backslash D$.
Proof.
(1) It is obvious by the definition of $\mathcal{F}_{e}$ and Lemma 8.4.
(2) Since the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is recurrent, by [26, Theorem 1.6.3], we have $1\in\mathcal{F}_{e}$ and $\mathcal{E}(1,1)=0$. For $u\in\mathcal{F}_{e}$, if $\mathcal{E}(u,u)=0$, then by MS($\Phi,\Psi$), we have $u$ is a constant function. Hence we have (RF1).
By (RF1), we have $(\mathcal{F}_{e}/\sim,\mathcal{E})$ is an inner product space. We only need to prove the completeness. Indeed, let $\{u_{n}\}\subseteq\mathcal{F}_{e}/\sim$ be an $\mathcal{E}$-Cauchy sequence, then $\{u_{n}\}$ is $\mathcal{E}$-bounded. Fix arbitrary $x_{0}\in X$. By replacing $u_{n}$ by $u_{n}-u_{n}(x_{0})$ in $\mathcal{F}_{e}/\sim$, we may assume that $u_{n}(x_{0})=0$ for any $n\geq 1$.
For any $N\geq 1$, for any $x\in B(x_{0},N)$, for any $n\geq 1$, by MS($\Phi,\Psi$), we have
$|u_{n}(x)|^{2}=|u_{n}(x)-u_{n}(x_{0})|^{2}\leq C_{MS}\frac{\Psi(d(x,x_{0}))}{\Phi(d(x,x_{0}))}\mathcal{E}(u_{n},u_{n})\leq C_{MS}\frac{\Psi(N)}{\Phi(N)}\sup_{n\geq 1}\mathcal{E}(u_{n},u_{n}).$
By MS($\Phi,\Psi$) again, using the Arzelà-Ascoli lemma and the diagonal argument, there exists a subsequence, still denoted by $\{u_{n}\}$, and $u\in C(X)$ such that $u_{n}$ converges uniformly to $u$ on any relatively compact open subset of $X$. By the definition of $\mathcal{F}_{e}$, we have $u\in\mathcal{F}_{e}$ and $\{u_{n}\}$ is $\mathcal{E}$-convergent to $u$. Hence $(\mathcal{F}_{e}/\sim,\mathcal{E})$ is complete, hence we have (RF2).
(RF3) follows from the regularity of the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$.
(RF5) follows from [26, Corollary 1.6.3].
(3) We need to prove that
$\sup\left\{\frac{|u(x)-u(y)|^{2}}{\mathcal{E}(u,u)}:u\in\mathcal{F},\mathcal{E}(u,u)>0\right\}=\sup\left\{\frac{|u(x)-u(y)|^{2}}{\mathcal{E}(u,u)}:u\in\mathcal{F}_{e},\mathcal{E}(u,u)>0\right\}.$
Indeed, $\leq$” is trivial by $\mathcal{F}\subseteq\mathcal{F}_{e}$. $\geq$”: For any $u\in\mathcal{F}_{e}$ with $\mathcal{E}(u,u)>0$, there exists $\{u_{n}\}\subseteq\mathcal{F}$ which is $\mathcal{E}$-Cauchy and converges to $u$ $m$-a.e.. Since $\{u_{n}\}$ is $\mathcal{E}$-convergent to $u$, we may assume that $\mathcal{E}(u_{n},u_{n})>0$ for any $n\geq 1$. Take $x_{0}\in X$ such that $\lim_{n\to+\infty}u_{n}(x_{0})=u(x_{0})\in\mathbb{R}$, then $\{u_{n}(x_{0})\}$ is bounded. For any $N\geq 1$, for any $x\in B(x_{0},N)$, for any $n\geq 1$, by MS($\Phi,\Psi$), we have
$|u_{n}(x)|^{2}\leq 2\left(|u_{n}(x)-u_{n}(x_{0})|^{2}+|u_{n}(x_{0})|^{2}\right)$
$\leq 2\left(C_{MS}\frac{\Psi(d(x,x_{0}))}{\Phi(d(x,x_{0}))}\mathcal{E}(u_{n},u_{n})+\sup_{n\geq 1}|u_{n}(x_{0})|^{2}\right)$
$\leq 2\left(C_{MS}\frac{\Psi(N)}{\Phi(N)}\sup_{n\geq 1}\mathcal{E}(u_{n},u_{n})+\sup_{n\geq 1}|u_{n}(x_{0})|^{2}\right).$
By MS($\Phi,\Psi$) again, using the Arzelà-Ascoli lemma and the diagonal argument, there exists a subsequence, still denoted by $\{u_{n}\}$, and $v\in C(X)$ such that $u_{n}$ converges uniformly to $v$ on any relatively compact open subset of $X$. Since $u_{n}$ also converges to $u$ $m$-a.e. and $u,v\in C(X)$, we have $u=v$. Hence
$\frac{|u(x)-u(y)|^{2}}{\mathcal{E}(u,u)}=\lim_{n\to+\infty}\frac{|u_{n}(x)-u_{n}(y)|^{2}}{\mathcal{E}(u_{n},u_{n})}\leq\text{LHS}.$
Taking supremum with respect to $u\in\mathcal{F}_{e}$ with $\mathcal{E}(u,u)>0$, we have $\geq$”.
(4) Since $(\mathcal{E},\mathcal{F}_{e})$ is a resistance form on $X$, by [44, PROPOSITION 9.13], we have the continuity and the quasi-continuity coincide. Then $\mathcal{F}_{e,D}$ given by [26, Equation (2.3.14)] can be re-written as
$\mathcal{F}_{e,D}:=\left\{u\in\mathcal{F}_{e}:\widetilde{u}=0\text{ q.e. on }X\backslash D\right\}=\left\{u\in\mathcal{F}_{e}\subseteq C(X):u=0\text{ on }X\backslash D\right\},$
where $\widetilde{u}$ is a quasi-continuous modification of $u\in\mathcal{F}_{e}$.
Since the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is regular, by [26, Exercise 1.4.1], we have $\mathcal{F}\cap C_{c}(X)$ is a special standard core of $\mathcal{E}$. Then (4) follows from the equivalence in [26, Theorem 2.3.3 (ii)]. ∎
By the resistance form theory, we can solve “boundary value problems” or take a trace of resistance forms as follows.
Lemma 8.7.
([44, Section 8]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$), UHK($\Psi$) and NLE($\Psi$). Let $Y\subsetneqq X$ be a non-empty closed subset and $\mathcal{F}_{e}|_{Y}=\{u|_{Y}:u\in\mathcal{F}_{e}\}$. Then for any $u\in\mathcal{F}_{e}$, there exists a unique $v\in\mathcal{F}_{e}$ which is harmonic in $X\backslash Y$ such that $v=u$ on $Y$. Moreover, $v$ is the unique function in $\mathcal{F}_{e}$ that minimizes the following variational problem
$\inf\left\{\mathcal{E}(v,v):v\in\mathcal{F}_{e},v=u\text{ on }Y\right\}.$
Let $h_{Y}:\mathcal{F}_{e}|_{Y}\to\mathcal{F}_{e}$ be given by $u|_{Y}\mapsto v$ for $u\in\mathcal{F}_{e}$ as above. Let
$\mathcal{E}|_{Y}(u,v)=\mathcal{E}(h_{Y}(u),h_{Y}(v))$
for $u,v\in\mathcal{F}_{e}|_{Y}$. Then $(\mathcal{E}|_{Y},\mathcal{F}_{e}|_{Y})$ is a resistance form on $Y$.
We have the following maximum principle.
Corollary 8.8.
(Maximum principle) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$), UHK($\Psi$) and NLE($\Psi$). Let $D\subsetneqq X$ be a non-empty open subset of $X$. Let $u,v\in\mathcal{F}_{e}$ be harmonic in $D$ and satisfy $u\leq v$ on $X\backslash D$, then $u\leq v$ in $D$. In particular, $\inf_{X\backslash D}u\leq u\leq\sup_{X\backslash D}u$.
Proof.
We only need to show that if $u\in\mathcal{F}_{e}$ is harmonic in $D$ and satisfies $u\leq 0$ on $X\backslash D$, then $u\leq 0$ in $D$. Indeed, $u\wedge 0\in\mathcal{F}_{e}$ satisfies $u\wedge 0=u$ on $X\backslash D$, by Lemma 8.7, we have
$\mathcal{E}(u,u)\geq\mathcal{E}(u\wedge 0,u\wedge 0)\geq\inf\left\{\mathcal{E}(v,v):v\in\mathcal{F}_{e},v=u\text{ on }X\backslash D\right\}=\mathcal{E}(u,u).$
Hence the above inequalities are indeed equalities. By the uniqueness of $u\in\mathcal{F}_{e}$ from Lemma 8.7, we have $u\wedge 0=u$, hence $u\leq 0$ in $D$. ∎
The effective resistance between certain two sets can be achieved by the energy of some functions as follows.
Corollary 8.9.
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$), UHK($\Psi$) and NLE($\Psi$). Let $A,B$ be two non-empty closed subsets of $X$ satisfying $A\cap B=\emptyset$ and $A\cup B\subsetneqq X$. If one of the following conditions holds.
  1. (a)
    $X\backslash A$ is relatively compact and $B$ is compact.
  2. (b)
    $A,B$ are compact.
Then there exists a unique $v\in\mathcal{F}_{e}$ with $v=0$ on $A$ and $v=1$ on $B$ such that
$R(A,B)^{-1}=\mathcal{E}(v,v)=\inf\left\{\mathcal{E}(v,v):v\in\mathcal{F}_{e},v=0\text{ on }A,v=1\text{ on }B\right\}.$
Moreover, $1_{A},1_{B}\in\mathcal{F}_{e}|_{A\cup B}$, $v=h_{A\cup B}(1_{B})\in\mathcal{F}_{e}$ and
$R(A,B)^{-1}=\mathcal{E}(h_{A\cup B}(1_{A}),h_{A\cup B}(1_{A}))=\mathcal{E}(h_{A\cup B}(1_{B}),h_{A\cup B}(1_{B})).$
Proof.
Since $B$ is compact, $X\backslash A$ is open and $B\subseteq X\backslash A$, by the regularity of $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$, there exists $u\in\mathcal{F}\cap C_{c}(X)\subseteq\mathcal{F}_{e}$ such that $0\leq u\leq 1$ in $X$, $u=1$ on $B$ and $\mathrm{supp}(u)\subseteq X\backslash A$. Hence $u|_{A\cup B}=1_{B}\in\mathcal{F}_{e}|_{A\cup B}$, $(1-u)|_{A\cup B}=1_{A}\in\mathcal{F}_{e}|_{A\cup B}$.
Since $A\cup B\subsetneqq X$ is a non-empty closed subset, by Lemma 8.7, there exists a unique $v=h_{A\cup B}(u|_{A\cup B})=h_{A\cup B}(1_{B})\in\mathcal{F}_{e}$ which is harmonic in $X\backslash(A\cup B)$ and $v=u$ on $A\cup B$, that is, $v|_{A\cup B}=1_{B}$, or $v=0$ on $A$ and $v=1$ on $B$. Moreover, $v\in\mathcal{F}_{e}$ also minimizes the variational problem
$\inf\left\{\mathcal{E}(v,v):v\in\mathcal{F}_{e},v=0\text{ on }A,v=1\text{ on }B\right\},$
and
$\mathcal{E}(v,v)=\mathcal{E}|_{A\cup B}(1_{B},1_{B})=\mathcal{E}|_{A\cup B}(1_{A},1_{A})$
$=\mathcal{E}(h_{A\cup B}(1_{B}),h_{A\cup B}(1_{B}))=\mathcal{E}(h_{A\cup B}(1_{A}),h_{A\cup B}(1_{A})).$
It remains to prove that if either (a) or (b) holds, then
$\mathcal{E}(v,v)=\inf\left\{\mathcal{E}(v,v):v\in\mathcal{F},v=0\text{ on }A,v=1\text{ on }B\right\},$
where the RHS is equal to $R(A,B)^{-1}$ by definition. Since $\mathcal{F}\subseteq\mathcal{F}_{e}$, we only need to show $\geq$”. By the maximum principle Corollary 8.8, we have $0\leq v\leq 1$ in $X$.
For (a), since $0\leq v\leq 1$ in $X$, $v=0$ on $A$ and $X\backslash A$ is a relatively compact open set, we have $v\in\mathcal{F}_{e}\cap L^{2}(X;m)=\mathcal{F}$, $\geq$” is obvious.
For (b), since $A,B$ are both compact, by a similar proof to Lemma 8.6 (2), there exists an approximation sequence $\{v_{n}\}\subseteq\mathcal{F}\subseteq C(X)$ of $v\in\mathcal{F}_{e}$ that converges uniformly to $v$ on $A\cup B$. For any $\varepsilon\in(0,1/4)$, there exists $N\geq 1$ such that for any $n\geq N$, we have $|v_{n}-v|<\varepsilon$ on $A\cup B$, hence $-\varepsilon<v_{n}<\varepsilon$ on $A$ and $1-\varepsilon<v_{n}<1+\varepsilon$ on $B$. Let $\varphi:\mathbb{R}\to\mathbb{R}$ be given by
$\varphi(t)=\frac{\frac{1}{2}}{\frac{1}{2}-\varepsilon}\left((t\wedge(1-\varepsilon))\vee\varepsilon-\frac{1}{2}\right)+\frac{1}{2},t\in\mathbb{R},$
that is, $\varphi\in C(\mathbb{R})$ satisfies $\varphi=0$ on $(-\infty,\varepsilon]$, $\varphi=1$ on $[1-\varepsilon,+\infty)$ and $\varphi$ is linear on $[\varepsilon,1-\varepsilon]$. Then $\varphi(0)=0$ and $|\varphi(t)-\varphi(s)|\leq\frac{\frac{1}{2}}{\frac{1}{2}-\varepsilon}|t-s|$ for any $t,s\in\mathbb{R}$. For any $n\geq N$, let $\widetilde{v}_{n}=\varphi(v_{n})$, then $\widetilde{v}_{n}\in\mathcal{F}$ satisfies $\widetilde{v}_{n}=0$ on $A$, $\widetilde{v}_{n}=1$ on $B$ and $\mathcal{E}(\widetilde{v}_{n},\widetilde{v}_{n})\leq\left(\frac{\frac{1}{2}}{\frac{1}{2}-\varepsilon}\right)^{2}\mathcal{E}(v_{n},v_{n})$, hence
$\inf\left\{\mathcal{E}(v,v):v\in\mathcal{F},v=0\text{ on }A,v=1\text{ on }B\right\}\leq\mathcal{E}(\widetilde{v}_{n},\widetilde{v}_{n})\leq\left(\frac{\frac{1}{2}}{\frac{1}{2}-\varepsilon}\right)^{2}\mathcal{E}(v_{n},v_{n}).$
Letting $n\to+\infty$, we have
$\inf\left\{\mathcal{E}(v,v):v\in\mathcal{F},v=0\text{ on }A,v=1\text{ on }B\right\}\leq\left(\frac{\frac{1}{2}}{\frac{1}{2}-\varepsilon}\right)^{2}\mathcal{E}(v,v).$
Since $\varepsilon\in(0,\frac{1}{4})$ is arbitrary, we have
$\inf\left\{\mathcal{E}(v,v):v\in\mathcal{F},v=0\text{ on }A,v=1\text{ on }B\right\}\leq\mathcal{E}(v,v),$
that is, we have $\geq$”. ∎
Remark 8.10.
We will be interested in the case $A=X\backslash B(x,r)$, $B=\{y\}$ for (a), and the case $A=\{x\}$, $B=\{y\}$ for (b), where $x,y\in X$, $r\in(0,+\infty)$.
Lemma 8.11.
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$), UHK($\Psi$) and NLE($\Psi$). Then there exists $C\in(0,+\infty)$ such that for any $x_{0}\in X$, for any $r\in(0,+\infty)$, we have
$R(x_{0},X\backslash B(x_{0},r))\geq C\frac{\Psi(r)}{\Phi(r)}.$
The proof is essentially the same as the proof of [12, LEMMA 2.4], [45, Lemma 4.1] and [37, Proposition 5.3].
Proof.
Denote $B=B(x_{0},r)$. For any $x\in B\backslash\left(\frac{1}{2}B\right)$, by Corollary 8.9, there exists $h_{x}\in\mathcal{F}_{e}$ with $h_{x}(x_{0})=1$, $h_{x}(x)=0$ and $0\leq h_{x}\leq 1$ in $X$ such that
$\mathcal{E}(h_{x},h_{x})=R(x_{0},x)^{-1}\leq\left(\frac{1}{C_{R}}\frac{\Psi(d(x_{0},x))}{\Phi(d(x_{0},x))}\right)^{-1}=C_{R}\frac{\Phi(d(x_{0},x))}{\Psi(d(x_{0},x))}\leq C_{R}\frac{\Phi(\frac{r}{2})}{\Psi(\frac{r}{2})},$
where $C_{R}$ is the constant in R($\Phi,\Psi$). Take $\varepsilon\in(0,\frac{1}{4})$ to be determined. For any $y\in B(x,\varepsilon r)$, we have
$|h_{x}(y)|=|h_{x}(x)-h_{x}(y)|\leq R(x,y)^{1/2}\mathcal{E}(h_{x},h_{x})^{1/2}$
$\leq\left(C_{R}\frac{\Psi(d(x,y))}{\Phi(d(x,y))}C_{R}\frac{\Phi(\frac{r}{2})}{\Psi(\frac{r}{2})}\right)^{1/2}\leq C_{R}\left(\frac{\frac{\Psi(\varepsilon r)}{\Phi(\varepsilon r)}}{\frac{\Psi(\frac{r}{2})}{\Phi(\frac{r}{2})}}\right)^{1/2}\leq C_{R}(2\varepsilon)^{\frac{{\gamma_{1}\wedge\gamma_{2}}}{2}}.$
Then there exists $\varepsilon\in(0,\frac{1}{4})$ depending only on $C_{R},\gamma_{1},\gamma_{2}$ sufficiently small such that
$|h_{x}(y)|\leq C_{R}(2\varepsilon)^{\frac{{\gamma_{1}\wedge\gamma_{2}}}{2}}\leq\frac{1}{2}\text{ for any }y\in B(x,\varepsilon r).$
By V($\Phi$), there exists $N\geq 1$ depending only on $C_{VR},\alpha_{1},\alpha_{2},\varepsilon$, hence only on $C_{VR}$, $C_{R}$, $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, $\beta_{2}$, such that $B\backslash\left(\frac{1}{2}B\right)$ can ben covered by at most $N$ balls with center in $B\backslash\left(\frac{1}{2}B\right)$ and radius $\varepsilon r$. Take $x_{1},\ldots,x_{N}\in B\backslash\left(\frac{1}{2}B\right)$ such that $B\backslash\left(\frac{1}{2}B\right)\subseteq\cup_{1\leq i\leq N}B(x_{i},\varepsilon r)$, let $h=\min_{1\leq i\leq N}h_{x_{i}}$. Then $h\in\mathcal{F}_{e}$ satisfies $h(x_{0})=1$, $0\leq h\leq 1$ in $X$, $h\leq\frac{1}{2}$ in $B\backslash\left(\frac{1}{2}B\right)$ and
$\mathcal{E}(h,h)\leq\sum_{i=1}^{N}\mathcal{E}(h_{x_{i}},h_{x_{i}})\leq\sum_{i=1}^{N}C_{R}\frac{\Phi(\frac{r}{2})}{\Psi(\frac{r}{2})}=C_{R}N\frac{\Phi(\frac{r}{2})}{\Psi(\frac{r}{2})}.$
Let $g=2((h-\frac{1}{2})\vee 0)$, then $g\in\mathcal{F}_{e}$ satisfies $g(x_{0})=1$, $0\leq g\leq 1$ in $X$, $g=0$ in $B\backslash\left(\frac{1}{2}B\right)$ and
$\mathcal{E}(g,g)=4\mathcal{E}((h-\frac{1}{2})\vee 0,(h-\frac{1}{2})\vee 0)\leq 4\mathcal{E}(h,h).$
By the regularity of the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$, there exists $\varphi\in\mathcal{F}\cap C_{c}(X)$ satisfying $0\leq\varphi\leq 1$ in $X$, $\varphi=1$ in an open neighborhood of $\overline{\frac{1}{2}B}$ and $\mathrm{supp}(\varphi)\subseteq B$. Then $\varphi g\in\mathcal{F}_{e}$ satisfies $(\varphi g)(x_{0})=1$, $0\leq\varphi g\leq 1$ in $X$ and $\varphi g=g1_{B}=g1_{\frac{1}{2}{B}}$. By the strongly local property and Corollary 8.9, we have
$R(x_{0},X\backslash B)^{-1}\leq\mathcal{E}(\varphi g,\varphi g)=\Gamma(\varphi g,\varphi g)(X)=\Gamma(\varphi g,\varphi g)(B)$
$=\Gamma(g,g)(B)\leq\Gamma(g,g)(X)=\mathcal{E}(g,g)\leq 4\mathcal{E}(h,h)\leq 4C_{R}N\frac{\Phi(\frac{r}{2})}{\Psi(\frac{r}{2})},$
hence
$R(x_{0},X\backslash B)\geq\frac{1}{4C_{R}N}\frac{\Psi(\frac{r}{2})}{\Phi(\frac{r}{2})}=\frac{1}{4C_{R}N}\frac{\frac{\Psi(\frac{r}{2})}{\Phi(\frac{r}{2})}}{\frac{\Psi(r)}{\Phi(r)}}\frac{\Psi(r)}{\Phi(r)}\geq\frac{1}{4\cdot 2^{\gamma_{1}\vee\gamma_{2}}C_{R}N}\frac{\Psi(r)}{\Phi(r)}.$
Proposition 8.12.
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$),
UHK($\Psi$) and NLE($\Psi$). Then for any ball $B=B(x_{0},r)$, for any bounded $u\in\mathcal{F}_{e}$ which is harmonic in $B$, for any $x,y\in B$, we have
$|u(x)-u(y)|\leq\frac{R(x,y)}{R(x,X\backslash B)}\mathrm{osc}_{X\backslash B}u,$
where
$\mathrm{osc}_{A}u=\sup\nolimits_{A}u-\inf\nolimits_{A}u$
is the oscillation of the function $u\in l(A)$ on the set $A$.
The idea of the proof is from a recent result [40, Theorem 6.27], where a general result in the setting of $p$-energy forms was given to prove certain $p$-resistance defines indeed a metric.
Proof.
Let $Y=X\backslash B$, then $Y\subsetneqq X$ is a non-empty closed subset of $X$. For fixed $x,y\in B$, we have $u-u(x)\in\mathcal{F}_{e}$ is harmonic in $B\backslash\{x\}$, hence $u-u(x)=h_{Y\cup\{x\}}(u-u(x))$. By the maximum principle Corollary 8.8, we have $u-u(x)\leq(\mathrm{osc}_{Y}u)1_{Y}$ on $Y\cup\{x\}$, hence
$u-u(x)=h_{Y\cup\{x\}}(u-u(x))\leq h_{Y\cup\{x\}}\left((\mathrm{osc}_{Y}u)1_{Y}\right)=(\mathrm{osc}_{Y}u)h_{Y\cup\{x\}}(1_{Y}).$
It remains to estimate $h_{Y\cup\{x\}}(1_{Y})(y)$. Indeed, by Corollary 8.9, we have
$R(x,X\backslash B)^{-1}=R(x,Y)^{-1}=\mathcal{E}(h_{Y\cup\{x\}}(1_{Y}),h_{Y\cup\{x\}}(1_{Y}))$
$=\mathcal{E}(h_{Y\cup\{x\}}(1_{Y}),h_{Y\cup\{x\}}(1_{Y})-1)=-\mathcal{E}(h_{Y\cup\{x\}}(1_{Y}),h_{Y\cup\{x\}}(1_{\{x\}})).$
Since $\mathcal{F}_{e}\ni h_{Y\cup\{x,y\}}(1_{\{x\}})-h_{Y\cup\{x\}}(1_{\{x\}})=0$ on $Y\cup\{x\}$ and $h_{Y\cup\{x\}}(1_{Y})\in\mathcal{F}_{e}$ is harmonic in $B\backslash\{x\}$, by Lemma 8.6 (4), we have
$\mathcal{E}(h_{Y\cup\{x\}}(1_{Y}),h_{Y\cup\{x,y\}}(1_{\{x\}})-h_{Y\cup\{x\}}(1_{\{x\}}))=0,$
that is,
$\mathcal{E}(h_{Y\cup\{x\}}(1_{Y}),h_{Y\cup\{x\}}(1_{\{x\}}))=\mathcal{E}(h_{Y\cup\{x\}}(1_{Y}),h_{Y\cup\{x,y\}}(1_{\{x\}})).$
Let
$u_{1}$ $=h_{Y\cup\{x\}}(1_{Y})-h_{Y\cup\{x\}}(1_{Y})(y)h_{\{x,y\}}(1_{\{y\}}),$
$v_{1}$ $=h_{Y\cup\{x,y\}}(1_{\{x\}}),$
then $u_{1},v_{1}\in\mathcal{F}_{e}$ are harmonic in $B\backslash\{x,y\}$. Let $u_{2}=u_{1}|_{Y\cup\{x,y\}}$, $v_{2}=v_{1}|_{Y\cup\{x,y\}}$, then $u_{2},v_{2}\in\mathcal{F}_{e}|_{Y\cup\{x,y\}}$ and
$\mathcal{E}(u_{1},v_{1})=\mathcal{E}|_{Y\cup\{x,y\}}(u_{2},v_{2}),$
where $(\mathcal{E}|_{Y\cup\{x,y\}},\mathcal{F}_{e}|_{Y\cup\{x,y\}})$ is a resistance form on $Y\cup\{x,y\}$, which follows from Lemma 8.7. On $Y$, we have $u_{2}\in[0,1]$, $v_{2}=0$, $u_{2}\wedge v_{2}=0$, $u_{2}\vee v_{2}=u_{2}=u_{2}+v_{2}$. On $\{x\}$, we have $u_{2}=0$, $v_{2}=1$, $u_{2}\wedge v_{2}=0$, $u_{2}\vee v_{2}=1=v_{2}=u_{2}+v_{2}$. On $\{y\}$, we have $u_{2}=0$, $v_{2}=0$, $u_{2}\wedge v_{2}=0$, $u_{2}\vee v_{2}=0=u_{2}+v_{2}$. Hence $u_{2}\wedge v_{2}=0$, $u_{2}\vee v_{2}=u_{2}+v_{2}$. By the following easy result Lemma 8.13, we have
$\mathcal{E}(u_{1},v_{1})=\mathcal{E}|_{Y\cup\{x,y\}}(u_{2},v_{2})\leq 0,$
that is,
$\mathcal{E}(h_{Y\cup\{x\}}(1_{Y}),h_{Y\cup\{x,y\}}(1_{\{x\}}))\leq\mathcal{E}(h_{Y\cup\{x\}}(1_{Y})(y)h_{\{x,y\}}(1_{\{y\}}),h_{Y\cup\{x,y\}}(1_{\{x\}})),$
hence
$R(x,X\backslash B)^{-1}=-\mathcal{E}(h_{Y\cup\{x\}}(1_{Y}),h_{Y\cup\{x\}}(1_{\{x\}}))=-\mathcal{E}(h_{Y\cup\{x\}}(1_{Y}),h_{Y\cup\{x,y\}}(1_{\{x\}}))$
$\geq-\mathcal{E}(h_{Y\cup\{x\}}(1_{Y})(y)h_{\{x,y\}}(1_{\{y\}}),h_{Y\cup\{x,y\}}(1_{\{x\}}))$
$=-h_{Y\cup\{x\}}(1_{Y})(y)\mathcal{E}(h_{\{x,y\}}(1_{\{y\}}),h_{Y\cup\{x,y\}}(1_{\{x\}})).$
Since $\mathcal{F}_{e}\ni h_{Y\cup\{x,y\}}(1_{\{x\}})-h_{\{x,y\}}(1_{\{x\}})=0$ on $\{x,y\}$ and $h_{\{x,y\}}(1_{\{y\}})\in\mathcal{F}_{e}$ is harmonic in $X\backslash\{x,y\}$, by Lemma 8.6 (4), we have
$\mathcal{E}(h_{\{x,y\}}(1_{\{y\}}),h_{Y\cup\{x,y\}}(1_{\{x\}})-h_{\{x,y\}}(1_{\{x\}}))=0,$
that is,
$\mathcal{E}(h_{\{x,y\}}(1_{\{y\}}),h_{Y\cup\{x,y\}}(1_{\{x\}}))=\mathcal{E}(h_{\{x,y\}}(1_{\{y\}}),h_{\{x,y\}}(1_{\{x\}})).$
Hence
$R(x,X\backslash B)^{-1}\geq-h_{Y\cup\{x\}}(1_{Y})(y)\mathcal{E}(h_{\{x,y\}}(1_{\{y\}}),h_{Y\cup\{x,y\}}(1_{\{x\}}))$
$=-h_{Y\cup\{x\}}(1_{Y})(y)\mathcal{E}(h_{\{x,y\}}(1_{\{y\}}),h_{\{x,y\}}(1_{\{x\}}))$
$=h_{Y\cup\{x\}}(1_{Y})(y)\mathcal{E}(1-h_{\{x,y\}}(1_{\{y\}}),h_{\{x,y\}}(1_{\{x\}}))$
$=h_{Y\cup\{x\}}(1_{Y})(y)\mathcal{E}(h_{\{x,y\}}(1_{\{x\}}),h_{\{x,y\}}(1_{\{x\}}))$
$=h_{Y\cup\{x\}}(1_{Y})(y)R(x,y)^{-1},$
which implies
$h_{Y\cup\{x\}}(1_{Y})(y)\leq\frac{R(x,y)}{R(x,X\backslash B)}.$
Thus we have
$u(y)-u(x)\leq\left(\mathrm{osc}_{Y}u\right)\frac{R(x,y)}{R(x,X\backslash B)}.$
Similarly, since $u-u(x)\geq-(\mathrm{osc}_{Y}u)1_{Y}$ on $Y\cup\{x\}$, we have
$u-u(x)=h_{Y\cup\{x\}}(u-u(x))\geq-h_{Y\cup\{x\}}\left((\mathrm{osc}_{Y}u)1_{Y}\right)=-\left(\mathrm{osc}_{Y}u\right)h_{Y\cup\{x\}}(1_{Y}),$
which implies
$u(y)-u(x)\geq-\left(\mathrm{osc}_{Y}u\right)h_{Y\cup\{x\}}(1_{Y})(y)\geq-\left(\mathrm{osc}_{Y}u\right)\frac{R(x,y)}{R(x,X\backslash B)}.$
Therefore, we have
$|u(x)-u(y)|\leq\frac{R(x,y)}{R(x,X\backslash B)}\mathrm{osc}_{Y}u.$
Lemma 8.13.
Let $X$ be a set. Let $(\mathcal{E},\mathcal{G})$ be a resistance form on $X$. If $u,v\in\mathcal{G}$ satisfy $u\wedge v=0$ and $u\vee v=u+v$, then $\mathcal{E}(u,v)\leq 0$.
Proof.
By the parallelogram identity and (RF5), we have
$\mathcal{E}(u\wedge v,u\wedge v)+\mathcal{E}(u\vee v,u\vee v)\leq\mathcal{E}(u,u)+\mathcal{E}(v,v).$
By assumption, we have
$\mathcal{E}(u,u)+\mathcal{E}(v,v)\geq\mathcal{E}(u+v,u+v)=\mathcal{E}(u,u)+\mathcal{E}(v,v)+2\mathcal{E}(u,v),$
hence $\mathcal{E}(u,v)\leq 0$. ∎
We need the following $L^{1}$-mean value inequality.
Lemma 8.14.
([31, THEOREM 6.3, LEMMA 9.2]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD, FK($\Psi$) and CS($\Psi$). Then there exists $C\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $u\in\mathcal{F}$ which is harmonic in $2B$, we have
$\lVert u\rVert_{L^{\infty}(B)}\leq C\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$
[…]
Note that we need to prove HR($\Phi,\Psi$) from UHK($\Psi$) and NLE($\Psi$). By Proposition 8.2, we have R($\Phi,\Psi$). Let $B=B(x_{0},r)$ and $u\in\mathcal{F}$ harmonic in $2B$. By VD and Lemma 8.14, we have
$M:=\lVert u\rVert_{L^{\infty}(\frac{3}{2}B)}\lesssim\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$
Let $v=(u\vee(-M))\wedge M$, then $v\in\mathcal{F}\cap L^{\infty}(X;m)$, $\lVert v\rVert_{L^{\infty}(X)}\leq M$ and $v=u$ in $\frac{3}{2}B$, hence $v\in\mathcal{F}\subseteq\mathcal{F}_{e}$ is bounded in $X$ and harmonic in $\frac{3}{2}B$. By Proposition 8.12, for any $x,y\in B\subseteq\frac{3}{2}B$, we have
$|v(x)-v(y)|\leq\frac{R(x,y)}{R(x,X\backslash\left(\frac{3}{2}B\right))}\mathrm{osc}_{X\backslash\left(\frac{3}{2}B\right)}v$
$\leq 2\frac{R(x,y)}{R(x,X\backslash B(x,\frac{r}{2}))}\lVert v\rVert_{L^{\infty}(X)}\leq 2M\frac{R(x,y)}{R(x,X\backslash B(x,\frac{r}{2}))},$
hence
$|u(x)-u(y)|\leq 2\lVert u\rVert_{L^{\infty}(\frac{3}{2}B)}\frac{R(x,y)}{R(x,X\backslash B(x,\frac{r}{2}))}\lesssim\frac{R(x,y)}{R(x,X\backslash B(x,\frac{r}{2}))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$
By R($\Phi,\Psi$) and Lemma 8.11, we have
$|u(x)-u(y)|\lesssim\frac{\left(\frac{\Psi}{\Phi}\right)(d(x,y))}{\left(\frac{\Psi}{\Phi}\right)(\frac{r}{2})}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m\asymp\frac{\left(\frac{\Psi}{\Phi}\right)(d(x,y))}{\left(\frac{\Psi}{\Phi}\right)(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m,$
hence HR($\Phi,\Psi$) holds. ∎
Remark 8.15.
A natural question is to consider the sharpness of the Hölder exponents $\beta_{1}-\alpha_{1}$ and $\beta_{2}-\alpha_{2}$ from HR($\Phi,\Psi$). Indeed, on a class of unbounded p.c.f. self-similar sets (see [54]) and the corresponding cable systems, these exponents are sharp, see the proof of [25, Proposition 4.1, Proposition 4.3] for the Vicsek and the SG cable systems. However, on the unbounded SC and the SC cable system, it was conjectured in [1, Conjecture 5.4, Open Question 5.5] that these exponents are not sharp, and a new fractal dimension, that is, the so-called topological Hausdorff dimension (see [4]), should be involved.
9 Some examples related to the Sierpiński carpet and the Vicsek set
In this section, we give some examples related to the Sierpiński carpet and the Vicsek set including the blowups of a fractal by another different fractal. In $\mathbb{R}^{2}$, let
$p_{1}=(0,0),p_{2}=(\frac{1}{2},0),p_{3}=(1,0),p_{4}=(1,\frac{1}{2}),$ (9.1)
$p_{5}=(1,1),p_{6}=(\frac{1}{2},1),p_{7}=(0,1),p_{8}=(0,\frac{1}{2}),$
and
$g_{i}(x)=\frac{1}{3}\left(x-p_{i}\right)+p_{i},x\in\mathbb{R}^{2},i=1,\ldots,8.$
Then the Sierpiński carpet is the unique non-empty compact set $K$ in $\mathbb{R}^{2}$ satisfying $K=\cup_{i=1}^{8}g_{i}(K)$, see Figure 2. Let $\alpha=\log 8/\log 3$. By [42, Theorem 1.5.7, Proposition 1.5.8], the Hausdorff dimension of $K$ is $\alpha$ and the $\alpha$-dimensional Hausdorff measure $\mathcal{H}^{\alpha}(K)\in(0,+\infty)$. The construction of Brownian motion on the Sierpiński carpet was given by [6] and [46] using approximation of stochastic processes. See also [33] for another construction using $\Gamma$-convergence of quadratic forms, [9] for the construction on higher dimensional Sierpiński carpets, [17] for the construction on unconstrained Sierpiński carpets. Let $\beta=\log(8\rho)/\log 3\approx 2.09697$, where $\rho\approx 1.25147\in[\frac{7}{6},\frac{3}{2}]$ is some parameter from resistance estimates in [7, 11].
For $a\in\mathbb{R}$, $A\subseteq\mathbb{R}^{2}$, denote $aA=\{ax:x\in A\}$.
Firstly, we consider the unbounded Sierpiński carpet $K^{\infty}=\cup_{n\geq 0}3^{n}K$ as follows, where $\{3^{n}K\}_{n\geq 0}$ is an increasing sequence of subsets of $\mathbb{R}^{2}$. Let $\alpha_{1}=\alpha_{2}=\alpha$, $\beta_{1}=\beta_{2}=\beta$. It is obvious that V($\Phi$) holds. It was proved in [8, Theorem 1.1] that HK($\Psi$) and HHK($\Phi,\Psi$) hold. In a general setting of fractional diffusions introduced in [5, Section 3], it was proved in [1, Theorem 3.7, Lemma 3.4] that wBE($\Phi,\Psi$) and HHK($\Phi,\Psi$) hold. It was proved in [8, Theorem 8.3] that all functions in the domain of the generator of the Dirichlet form are $(\beta-\alpha)$-Hölder continuous. By the main result, Theorem 2.1, we now have the Hölder regularity for harmonic functions as HR($\Phi,\Psi$) and the Hölder estimate $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ with exponential terms for the heat kernel.
Secondly, we consider the Sierpiński carpet cable system. Let $V_{0}=\{p_{1},\ldots,p_{8}\}$ and $V_{n+1}=\cup_{i=1}^{8}g_{i}(V_{n})$ for any $n\geq 0$, then $\{V_{n}\}_{n\geq 0}$ is an increasing sequence of finite subsets of $K$ and the closure of $\cup_{n\geq 0}V_{n}$ with respect to the Euclidean topology is $K$. For any $n\geq 0$, let $V^{(n)}=3^{n}V_{n}$, see Figure 4 for $V^{(0)}$, $V^{(1)}$ and $V^{(2)}$.
(a) $V^{(0)}$
(b) $V^{(1)}$
(c) $V^{(2)}$
Figure 4: $V^{(0)}$, $V^{(1)}$ and $V^{(2)}$
Obviously, $\{V^{(n)}\}_{n\geq 0}$ is an increasing sequence of finite sets. Let $V=\cup_{n\geq 0}V^{(n)}$ and
$E=\left\{\{p,q\}:p,q\in V,|p-q|=\frac{1}{2}\right\},$
then $(V,E)$ is an infinite, locally bounded, connected (undirected) graph, whose corresponding cable system is the Sierpiński carpet cable system, see [25, Section 3] for more detailed construction. Let $\alpha_{1}=1$, $\beta_{1}=2$, $\alpha_{2}=\alpha$, $\beta_{2}=\beta$. It is obvious that V($\Phi$) holds. Using the stability result [10] for heat kernel estimates or parabolic Harnack inequalities, we have HK($\Psi$) also holds, which is similar to the case on the unbounded Sierpiński carpet. By the main result, Theorem 2.1, we have HHK($\Phi,\Psi$), $\text{HHK}_{\text{exp}}(\Phi,\Psi)$, wBE($\Phi,\Psi$) and HR($\Phi,\Psi$) all hold. In particular, the validity of HR($\Phi,\Psi$) implies the validity of GRH($\Phi,\Psi$), which was left open in [25]. Moreover, we also have GHK($\Phi,\Psi$), which can also be given directly by $\text{HHK}_{\text{exp}}(\Phi,\Psi)$.
Thirdly, we consider the blowups of fractals related to the Sierpiński carpet and the Vicsek set. However, we blowup a fractal by another different fractal. We consider the setting in $\mathbb{R}^{2}$.
Let $K,M$ be one of the following fractals in $\mathbb{R}^{2}$: the unit interval $[0,1]\times\{0\}$, the Vicsek set and the Sierpiński carpet. Let $\alpha_{1}$ and $\beta_{1}$ be the Hausdorff dimension and walk dimension of $K$, $\alpha_{2}$ and $\beta_{2}$ the Hausdorff dimension and walk dimension of $M$. Here, $K$ serves as the “cell” fractal and $M$ provides the “model” to blowup. Let $(\mathcal{E}^{K},\mathcal{F}^{K})$ be the strongly local regular Dirichlet form on $L^{2}(K;\mu)$ with two-sided sub-Gaussian heat kernel estimates with walk dimension $\beta_{1}$, where $\mu$ is the normalized $\alpha_{1}$-Hausdorff measure on $K$, see [48, 42, 53] for the case $K$ is the Vicsek set.
Let $\{g^{(1)}_{i}\}_{i=1,\ldots,N^{(1)}}$, $\{g^{(2)}_{i}\}_{i=1,\ldots,N^{(2)}}$ be the collections of contraction mappings generating $K$, $M$, respectively, that is, for $k=1,2$, $N^{(k)},L^{(k)}$ are two positive integers, $p^{(k)}_{1}=(0,0)$, …, $p^{(k)}_{N^{(k)}}\in\mathbb{R}^{2}$,
$g^{(k)}_{i}(x)=\frac{1}{L^{(k)}}x+\frac{L^{(k)}-1}{L^{(k)}}p_{i},x\in\mathbb{R}^{2},i=1,\ldots,N^{(k)},$
and $K$, $M$ are the unique non-empty compact sets in $\mathbb{R}^{2}$ satisfying $K=\cup_{i=1}^{N^{(1)}}g^{(1)}_{i}(K)$, $M=\cup_{i=1}^{N^{(2)}}g^{(2)}_{i}(M)$, respectively.
  • For the unit interval, we have $N^{(k)}=2$, $L^{(k)}=2$ and
    $p^{(k)}_{1}=(0,0),p^{(k)}_{2}=(1,0).$
  • For the Vicsek set, we have $N^{(k)}=5$, $L^{(k)}=3$ and
    $p^{(k)}_{1}=(0,0),p^{(k)}_{2}=(1,0),p^{(k)}_{3}=(1,1),p^{(k)}_{4}=(0,1),p^{(k)}_{5}=\left(\frac{1}{2},\frac{1}{2}\right).$
  • For the Sierpiński carpet, we have $N^{(k)}=8$, $L^{(k)}=3$ and $p^{(k)}_{1},\ldots,p^{(k)}_{8}$ are $p_{1},\ldots,p_{8}$ from Equation (9.1).
We assume that for $i,j=1,\ldots,N^{(2)}$, if $g^{(2)}_{i}\left([0,1]^{2}\right)\cap g^{(2)}_{j}\left([0,1]^{2}\right)\neq\emptyset$, then $g^{(2)}_{i}(K)\cap g^{(2)}_{j}(K)\neq\emptyset$. Let $K_{0}=K$. For any $n\geq 0$, let $K_{n+1}=\cup_{i=1}^{N^{(2)}}g^{(2)}_{i}(K_{n})$, $K^{(n)}=(L^{(2)})^{n}K_{n}$, then
$K^{(n+1)}=(L^{(2)})^{n+1}K_{n+1}=(L^{(2)})^{n+1}\left(\cup_{i=1}^{N^{(2)}}g^{(2)}_{i}(K_{n})\right)$
$=(L^{(2)})^{n+1}\left(g^{(2)}_{1}(K_{n})\cup\bigcup_{i=2}^{N^{(2)}}g^{(2)}_{i}(K_{n})\right)$
$=\left((L^{(2)})^{n}K_{n}\right)\cup\left((L^{(2)})^{n+1}\left(\bigcup_{i=2}^{N^{(2)}}g^{(2)}_{i}(K_{n})\right)\right)$
$\supseteq(L^{(2)})^{n}K_{n}=K^{(n)},$
hence $\{K^{(n)}\}_{n\geq 0}$ is an increasing sequence of subsets of $\mathbb{R}^{2}$.
Let $X=\cup_{n\geq 0}K^{(n)}$. We say that $W\subseteq X$ is a cell if $W$ is a translation of $K^{(0)}=K_{0}=K$. Let $d$ be the Euclidean metric in $\mathbb{R}^{2}$. Let $m$ be the measure on $X$ satisfying that for any cell $W$, $m|_{W}$ is the normalized $\alpha_{1}$-Hausdorff measure on $W$. For any cell $W$, by translation, there exists a strongly local regular Dirichlet form $(\mathcal{E}^{W},\mathcal{F}^{W})$ on $L^{2}(W;m|_{W})$. Let
$\mathcal{K}=\left\{u\in C_{c}(X):u|_{W}\in\mathcal{F}^{W}\text{ for any cell }W\right\}.$
Define
$\mathcal{E}(u,u)$ $=\sum_{W\text{ is a cell}}\mathcal{E}^{W}(u|_{W},u|_{W}),$
$\mathcal{F}$ $=\mathcal{E}_{1}\text{-closure of }\mathcal{K},$
then $(\mathcal{E},\mathcal{F})$ is a strongly local regular Dirichlet form on $L^{2}(X;m)$. Obviously, $(X,d,m,\mathcal{E},\mathcal{F})$ is an unbounded MMD space satisfying V($\Phi$). Since two-sided heat kernel estimates hold on $K,M$, by [30, Theorem 3.14], we have the elliptic Harnack inequalities and the resistance estimates both hold on $K,M$. Combining these two conditions on $K,M$, we obtain easily the corresponding elliptic Harnack inequalities and the resistance estimates on $X$. Then by [30, Theorem 3.14] again, we have HK($\Psi$) holds on $X$. Indeed, in the strongly recurrent setting, only resistance estimates are enough for the above argument, see [12, 45, 37]. By the main result, Theorem 2.1, we have HHK($\Phi,\Psi$), $\text{HHK}_{\text{exp}}(\Phi,\Psi)$, wBE($\Phi,\Psi$) and HR($\Phi,\Psi$) all hold.
We give some examples as follows.
  1. (1)
    Let $K$ be the Sierpiński carpet and $M$ the unit interval, see Figure 5.
    Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
    Figure 5: Blow up the Sierpiński carpet by the unit interval
  2. (2)
    Let $K$ be the Sierpiński carpet and $M$ the Vicsek set, see Figure 6 for $K^{(2)}$.
    Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
    Figure 6: Blow up the Sierpiński carpet by the Vicsek set
  3. (3)
    Let $K$ be the Vicsek set and $M$ the Sierpiński carpet, see Figure 7 for $K^{(2)}$.
    Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
    Figure 7: Blow up the Vicsek set by the Sierpiński carpet
    Since $(\Psi/\Phi)(r)=r$ for $r\in(0,1)$, $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ gives
    $|p_{t}(x,y_{1})-p_{t}(x,y_{2})|$
    $\leq C_{1}\frac{d(y_{1},y_{2})}{t}\left(\exp\left(-\Upsilon\left(C_{2}d(x,y_{1}),t\right)\right)+\exp\left(-\Upsilon\left(C_{2}d(x,y_{2}),t\right)\right)\right)$ (9.2)
    for any $x,y_{1},y_{2}\in X$ with $d(y_{1},y_{2})<1$.
    Let $S_{0}=[(0,0),(1,1)]\cup[(0,1),(1,0)]$ be the union of the diagonals of the unit square $[0,1]^{2}\subseteq\mathbb{R}^{2}$. For any $n\geq 0$, let $S_{n+1}=\cup_{i=1}^{N^{(1)}}g_{i}^{(1)}(S_{n})$. Then $\{S_{n}\}_{n\geq 0}$ is an increasing sequence of subsets of $K$ and the closure of $\cup_{n\geq 0}S_{n}$ is $K$. Let $\mathcal{S}_{0}=\cup_{n\geq 0}S_{n}$, $\nu^{(0)}$ the measure on $\mathcal{S}_{0}$ such that for any $n\geq 0$, $\nu^{(0)}|_{S_{n}}$ is the normalized 1-dimensional Hausdorff measure satisfying
    $\nu^{(0)}\left([(0,0),(1,1)]\right)=\nu^{(0)}\left([(0,1),(1,0)]\right)=1.$
    For any $n\geq 0$, let $\mathcal{S}_{n+1}=\cup_{i=1}^{N^{(2)}}g_{i}^{(2)}(\mathcal{S}_{n})$, $\mathcal{S}^{(n)}=(L^{(2)})^{n}\mathcal{S}_{n}$, then similar to $\{K^{(n)}\}_{n\geq 0}$, $\{\mathcal{S}^{(n)}\}_{n\geq 0}$ is an increasing sequence of subsets of $\mathbb{R}^{2}$, let $\mathcal{S}=\cup_{n\geq 0}\mathcal{S}^{(n)}$. For any cell $W\subseteq X$, we have $\mathcal{S}_{W}:=\mathcal{S}\cap W$ is a translation of $\mathcal{S}^{(0)}=\mathcal{S}_{0}$. For any two cells $W_{1},W_{2}\subseteq X$, we have either the cardinal #$\#(\mathcal{S}_{W_{1}}\cap\mathcal{S}_{W_{2}})\in\{0,1\}$ or $\mathcal{S}_{W_{1}}\cap\mathcal{S}_{W_{2}}$ is countably infinite, hence there exists a unique $\nu$ on $\mathcal{S}$ such that for any cell $W$, $\nu|_{\mathcal{S}_{W}}$ is the push forward of $\nu^{(0)}$ under the translation from $\mathcal{S}^{(0)}=\mathcal{S}_{0}$ to $\mathcal{S}_{W}$. Note that $\nu(B(x_{0}),r)=+\infty$ for any ball $B(x_{0},r)$ and $\nu\perp m$, that is, $\nu$ and $m$ are mutually singular.
    By the Lipschitz estimate, Equation (9.2), we have the weak gradient $\partial p_{t}(x,\cdot)$ exists $\nu$-a.e. for any $t\in(0,+\infty)$, for any $x\in X$, and has the following estimate. For any $t\in(0,+\infty)$, for any $x\in X$, for $\nu$-a.e. $y\in\mathcal{S}$, we have
    $|\partial_{y}p_{t}(x,y)|\leq\frac{C_{1}}{t}\exp\left(-\Upsilon\left(C_{2}d(x,y),t\right)\right)$
    $\leq\begin{cases}\frac{C_{3}}{t}\exp\left(-C_{4}\left(\frac{d(x,y)}{t^{1/\beta_{1}}}\right)^{\frac{\beta_{1}}{\beta_{1}-1}}\right),&\text{ if }t<d(x,y),\\ \frac{C_{3}}{t}\exp\left(-C_{4}\left(\frac{d(x,y)}{t^{1/\beta_{2}}}\right)^{\frac{\beta_{2}}{\beta_{2}-1}}\right),&\text{ if }t\geq d(x,y).\\ \end{cases}$ (9.3)
    Note that $\partial p_{t}(x,\cdot)$ exists and has the above estimate $\nu$-a.e. instead of $m$-a.e.. At small scale, $\beta_{1}=\log 15/\log 3$, the gradient estimate has the sub-Gaussian behavior corresponding to the Vicsek set, while at large scale, $\beta_{2}\approx 2.09697$, the gradient estimate has the sub-Gaussian behavior corresponding to the Sierpiński carpet. See [15, Section 3] for the results about Lipschitz and gradient estimates for heat kernel on unbounded Vicsek set, that is, the blowup of the Vicsek set by the Vicsek set itself.
  4. (4)
    Let $K$ be the Vicsek set and $M$ the unit interval, see Figure 8.
    Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
    Figure 8: Blow up the Vicsek set by the unit interval
    Similar to the blowup of the Vicsek set by the Sierpiński carpet, we also have the Lipschitz estimate, Equation (9.2) and the gradient estimate, Equation (9.3). However, at small scale, $\beta_{1}=\log 15/\log 3$, the gradient estimate has the sub-Gaussian behavior corresponding to the Vicsek set, while at large scale, $\beta_{2}=2$, the gradient estimate has the Gaussian behavior corresponding to the unit interval.

### v3

@@
 Abstract
-We introduce the Hölder regularity condition for harmonic functions on metric measure spaces and prove that under mild volume regular condition and upper heat kernel estimate, the Hölder regularity condition, the weak Bakry-Émery non-negative curvature condition, the heat kernel Hölder continuity with or without exponential terms and the heat kernel near-diagonal lower bound are equivalent.
-As applications, firstly, we prove the validity of the so-called generalized reverse Hölder inequality on the Sierpiński carpet cable system, which was left open by Devyver, Russ, Yang (Int.
+We introduce a Hölder regularity condition for harmonic functions on metric measure spaces and prove that, under a slow volume regular condition and an upper heat kernel estimate, the Hölder regularity condition, the weak Bakry-Émery non-negative curvature condition, Hölder continuity of the heat kernel (with or without exponential terms), and the near-diagonal lower bound for the heat kernel are equivalent.
+As applications, first, we establish the validity of the so-called generalized reverse Hölder inequality on the Sierpiński carpet cable system, resolving an open problem left by Devyver, Russ, Yang (Int.
 Math.
@@
 18, 15537–15583).
-Secondly, we prove that two-sided heat kernel estimates alone imply gradient estimate for the heat kernel on strongly recurrent fractal-like cable systems, which improves the main results of the aforementioned paper.
-Thirdly, we obtain Hölder (Lipschitz) estimate for heat kernel on general metric measure spaces, which extends the classical Li-Yau gradient estimate for heat kernel on Riemannian manifolds.
-Date: August 18, 2024MSC2010: 28A80, 35K08Keywords: Hölder regularity, gradient estimate, harmonic function, heat kernel, Sierpiński carpet.Jin Gao was supported by National Natural Science Foundation of China (12271282).
-Jin Gao acknowledged Jiaxin Hu and Eryan Hu for reading the preliminary draft.
-Part of the work was carried out while Meng Yang was participating a Mini-Symposia in the 9th European Congress of Mathematics (9ECM) in Seville (Spain) between 15 and 19 July 2024, Meng Yang was very grateful to Fabrice Baudoin and Li Chen for the invitation to the symposia.
+Second, we prove that two-sided heat kernel estimates alone imply gradient estimates for the heat kernel on strongly recurrent fractal-like cable systems, improving the main results of the aforementioned paper.
+Third, we obtain Hölder (Lipschitz) estimates for the heat kernel on strongly recurrent metric measure spaces, extending the classical Li-Yau gradient estimate for the heat kernel on Riemannian manifolds.
+Date: MSC2020: 28A80, 35K08Keywords: Hölder regularity, gradient estimate, harmonic function, heat kernel, Sierpiński carpet.Jin Gao was supported by National Natural Science Foundation of China (Grant.
+No.
+12271282), and Zhejiang Provincial Natural Science Foundation of China (Grant.
+No.
+LQN25A010019).
+Jin Gao acknowledges Jiaxin Hu and Eryan Hu for reading the preliminary draft.
+Part of the work was carried out while Meng Yang was participating a Mini-Symposia in the 9th European Congress of Mathematics (9ECM) in Seville (Spain) between 15 and 19 July 2024, Meng Yang is very grateful to Fabrice Baudoin and Li Chen for the invitation to the symposia.
+The authors are grateful to Fabrice Baudoin for discussions on curvature conditions, and warmly thank Naotaka Kajino and Ryosuke Shimizu for pointing out an alternative proof of Proposition 8.12.
+The authors thank the anonymous referees for their helpful comments and suggestions, which have improved the presentation of this paper.
 1 Introduction
@@
 $\lambda|\xi|^{2}\leq\sum_{i,j=1}^{d}a_{ij}(x)\xi_{i}\xi_{j}\leq\Lambda|\xi|^{2}\text{ for any }x\in D,\xi=(\xi_{1},\ldots,\xi_{d})\in\mathbb{R}^{d}.$
-Then any weak solution $u\in W^{1,2}_{loc}(D)$ of $Lu=0$ in $D$, or ($L$-)harmonic function $u\in W^{1,2}_{loc}(D)$, is locally Hölder continuous in $D$ or $u\in C^{\alpha}(D)$, where $\alpha\in(0,1)$ is the Hölder exponent.
-One only needs the measurability of the coefficients $a_{ij}$ to obtain the Hölder regularity of the weak solution $u$.
-To prove this theorem, very powerful iteration techniques were developed, including de Giorgi iteration [24] and Moser iteration [49], and were applied in many areas later, see [18, 36, 21, 32, 13, 29] and references therein for analysis and heat kernel on metric measure spaces.
-However, iteration techniques give the existence of some Hölder exponent $\alpha$ which depends only on $d$, $\lambda$, $\Lambda$, but may not give an explicit value of $\alpha$.
-The main purpose of this paper is to consider Hölder regularity of harmonic functions on metric measure spaces, where certain explicit Hölder exponent can be obtained.
-These metric measure spaces will originate from certain strongly recurrent fractals.
-Fractals are new examples with very different phenomena from Riemannian manifolds.
+Then any weak solution $u\in W^{1,2}_{\mathrm{loc}}(D)$ of $Lu=0$ in $D$, that is, any ($L$-)harmonic function $u\in W^{1,2}_{\mathrm{loc}}(D)$, is locally Hölder continuous in $D$; equivalently, $u\in C^{\alpha}(D)$ for some $\alpha\in(0,1)$, where $\alpha$ is a Hölder exponent.
+One only needs the measurability of the coefficients $a_{ij}$ to obtain Hölder regularity of the weak solution $u$.
+To prove this theorem, very powerful iteration techniques were developed, including the De Giorgi iteration [25] and the Moser iteration [51], and were later applied in many areas; see [19, 38, 22, 34, 14, 30] and the references therein for applications to analysis and heat kernel estimates on metric measure spaces.
+However, these iteration techniques give the existence of some Hölder exponent $\alpha$, which depends only on $d,\lambda,\Lambda$, but generally do not provide an explicit value of $\alpha$.
+The main purpose of this paper is to consider Hölder regularity of harmonic functions on metric measure spaces, where an explicit Hölder exponent can be obtained.
+These metric measure spaces arise from certain strongly recurrent fractals.
+Fractals provide new examples exhibiting phenomena markedly different from those on Riemannian manifolds.
 On a complete non-compact Riemannian manifold, it was independently discovered by Grigor’yan [28] and Saloff-Coste [54, 55] that the following two-sided Gaussian estimate of the heat kernel
@@
 However, one important estimate on fractals is the following two-sided sub-Gaussian estimate of the heat kernel
-$\frac{C_{1}}{V(x,t^{1/\beta})}\exp\left(-C_{2}\left(\frac{d(x,y)}{t^{1/\beta}}\right)^{\frac{\beta}{\beta-1}}\right)\leq p_{t}(x,y)\leq\frac{C_{3}}{V(x,t^{1/\beta})}\exp\left(-C_{4}\left(\frac{d(x,y)}{t^{1/\beta}}\right)^{\frac{\beta}{\beta-1}}\right),$
-where $\beta$ is a new parameter called the walk dimension, which is always strictly greater than 2 on fractals.
-Moreover, we say that a fractal is strongly recurrent if its Hausdorff dimension $\alpha$ is strictly less than its walk dimension $\beta$.
-For example, on the Sierpiński gasket (see Figure 2), we have $\alpha=\log 3/\log 2<\beta=\log 5/\log 2$, see [14, 41], on the Sierpiński carpet (see Figure 2), we have $\alpha=\log 8/\log 3<\beta\approx 2.09697$, see [6, 7, 11, 8, 35, 46].
+$\frac{C_{1}}{V(x,t^{1/\beta})}\exp\left(-C_{2}\left(\frac{d(x,y)}{t^{1/\beta}}\right)^{\frac{\beta}{\beta-1}}\right)\leq p_{t}(x,y)\leq\frac{C_{3}}{V(x,t^{1/\beta})}\exp\left(-C_{4}\left(\frac{d(x,y)}{t^{1/\beta}}\right)^{\frac{\beta}{\beta-1}}\right),$ $\textrm{HK}(\beta)$
+where $\beta$ is a new parameter called the walk dimension, which is typically strictly greater than 2 on fractals.
+Notably, for $\beta=2$, $\textrm{HK}(\beta)$ is indeed HK(2).
+For example, on the Sierpiński gasket (see Figure 2), we have $\beta=\log 5/\log 2$, see [15, 43], while on the Sierpiński carpet (see Figure 2), we have $\beta\approx 2.09697$, see [7, 8, 12, 9, 37, 48].
 Figure 1: The Sierpiński gasket
 Figure 2: The Sierpiński carpet
-Another purpose of this paper is to improve some recent results by Devyver, Russ and the second named author [25], where gradient estimate for heat kernel was considered.
-Gradient estimate for heat kernel plays an important role in the $L^{p}$-boundedness of the Riesz transform for $p>2$, see [3, 22].
-To obtain pointwise upper sub-Gaussian estimate for the gradient of the heat kernel on fractal-like cable systems, they introduced the so-called generalized reverse Hölder inequality and verified this condition on the Vicsek and the Sierpiński gasket cable systems, where the proof can be generalized to the cable systems corresponding to a class of p.c.f.
+Another purpose of this paper is to improve some recent results of Devyver, Russ and the second named author [26], where gradient estimates for the heat kernel were considered.
+Gradient estimates for the heat kernel play an important role in the $L^{p}$-boundedness of the Riesz transform for $p>2$; see [4, 23].
+To obtain pointwise upper sub-Gaussian estimates for the gradient of the heat kernel on fractal-like cable systems, the authors of [26] introduced the so-called generalized reverse Hölder inequality and verified this condition on the Vicsek and Sierpiński gasket cable systems, where the proof can be extended to cable systems corresponding to a class of p.c.f.
 self-similar sets.
-Since the harmonic extension algorithm was intrinsically needed, their proof can not apply on the Sierpiński carpet cable system.
-In this paper, we introduce the Hölder regularity condition, which is stronger than the generalized reverse Hölder inequality, and, a little bit surprisingly, can be ensured by pointwise two-sided heat kernel estimates.
-Therefore, on the one hand, we prove the validity of the generalized reverse Hölder inequality on the Sierpiński carpet cable system, which was left open in [25, Page 15545, Line 2].
-On the other hand, to obtain gradient estimate for the heat kernel, we do not need to assume gradient type conditions, such as the generalized reverse Hölder inequality, but only pointwise two-sided heat kernel estimates and some mild volume condition are enough, thus we improve main results of [25, Theorem 1.1, Corollary 1.3].
-The next purpose of this paper is to generalize the recent result by Baudoin, Chen [15, Theorem 3.3] about Lipschitz estimate for heat kernel on the unbounded Vicsek set.
-Here we obtain Hölder (Lipschitz) estimate for heat kernel on general unbounded metric measure spaces, which may behave differently at small scale and large scale.
-We follow a concise analytical approach, thus avoid to going into the probabilistic framework of fractional diffusions by Barlow [5].
-To end the Introduction, let us state briefly our main result, more detailed statement will be given in Section 2.
-We introduce firstly two scaling functions.
-Let $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\in(0,+\infty)$ satisfy
-$22\leq\beta_{i}\leq\alpha_{i}+1,\alpha_{i}<\beta_{i}\text{ for }i=1,$ 2 .
-Let $\Phi,\Psi:(0,+\infty)\to(0,+\infty)$ be given by
-$\Phi(r)$ $=r^{\alpha_{1}}1_{(0,1)}(r)+r^{\alpha_{2}}1_{[1,+\infty)}(r),$
-$\Psi(r)$ $=r^{\beta_{1}}1_{(0,1)}(r)+r^{\beta_{2}}1_{[1,+\infty)}(r).$
-In an unbounded metric measure Dirichlet space, we introduce the Hölder regularity condition HR($\Phi,\Psi$) and prove that under mild volume regular condition V($\Phi$) and upper heat kernel estimate UHK($\Psi$), the following conditions are indeed equivalent.
-Since we do not need to assume the existence of any gradient operator associated with the metric measure Dirichlet space, in particular, a “carré du champ”, the above equivalence generalizes the corresponding result by Coulhon, Jiang, Koskela, Sikora [22, Theorem 1.2].
-In addition, we also extend the classical result on Riemannian manifolds to metric measure Dirichlet spaces, that is, the conjunction of UHK($\Psi$) and $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ implies NLE($\Psi$) on metric measure Dirichlet spaces, which extends that the conjunction of upper and gradient estimates of heat kernel implies lower estimate of heat kernel on Riemannian manifolds, see Cheeger, Yau [19] and Li, Yau [47].
+Since the harmonic extension algorithm is intrinsically needed, their proof does not apply to the Sierpiński carpet cable system.
+In this paper, we introduce a Hölder regularity condition, which is stronger than the generalized reverse Hölder inequality and, somewhat surprisingly, can be ensured by pointwise two-sided heat kernel estimates.
+As consequences, on the one hand, we establish the validity of the generalized reverse Hölder inequality on the Sierpiński carpet cable system, resolving an open problem left in [26, Page 15545, Line 2].
+On the other hand, to obtain gradient estimates for the heat kernel, we do not need to assume gradient-type conditions such as the generalized reverse Hölder inequality; instead, pointwise two-sided heat kernel estimates together with a slow volume regular condition suffice.
+This leads to an improvement of the main results of [26, Theorem 1.1, Corollary 1.3].
+The next purpose of this paper is to generalize a recent result of Baudoin, Chen [16, Theorem 3.3] about Lipschitz estimates for the heat kernel on the unbounded Vicsek set.
+In this work, we obtain Hölder (Lipschitz) estimates for the heat kernel on strongly recurrent unbounded metric measure spaces, which may exhibit different behaviors at small and large scales.
+We follow a concise analytic approach, thereby avoiding the probabilistic framework of fractional diffusions introduced by Barlow [6].
+We conclude the Introduction by briefly stating our main result; a more detailed statement will be given in Section 2.
+Let $(X,d,m)$ be a metric measure space and let $(\mathcal{E},\mathcal{F})$ be a strongly local regular Dirichlet form on $L^{2}(X;m)$.
+A simplified consequence of our results is that, if $X$ is $\alpha$-regular for some $\alpha>0$, that is, for any metric ball $B(x,r)$, we have
+$V(x,r)=m(B(x,r))\asymp r^{\alpha},$
+and if $(\mathcal{E},\mathcal{F})$ admits a heat kernel $p_{t}(x,y)$ satisfying $\textrm{HK}(\beta)$ with walk dimension $\beta>\alpha$, then all harmonic functions associated with $(\mathcal{E},\mathcal{F})$ are $(\beta-\alpha)$-Hölder continuous.
+Moreover, the same Hölder regularity holds for the heat kernel.
+This result is motivated by classical examples on fractals, including the Sierpiński gasket and the Sierpiński carpet; see [15, 9].
+In fact, we consider more general and natural assumptions. We assume the volume regular condition
+$V(x,r)\asymp\Phi(r),$
+where
+$\Phi(r)=r^{\alpha_{1}}1_{(0,1)}(r)+r^{\alpha_{2}}1_{[1,+\infty)}(r),$ (1.1)
+and $\alpha_{1},\alpha_{2}>0$. $\textrm{HK}(\beta)$ is also generalized by introducing a time-distance scaling function $\Psi$ of the form
+$\Psi(r)=r^{\beta_{1}}1_{(0,1)}(r)+r^{\beta_{2}}1_{[1,+\infty)}(r),$ (1.2)
+which involves two walk dimensions: $\beta_{1}$ at small scales and $\beta_{2}$ at large scales. We further assume the strongly recurrent condition
+$\alpha_{i}<\beta_{i}\text{ for }i=1,2.$ (1.3)
+Under these assumptions, the Hölder regularity result is stated with two Hölder exponents: $\beta_{1}-\alpha_{1}$ at small scales and $\beta_{2}-\alpha_{2}$ at large scales.
+More precisely, in an unbounded metric measure Dirichlet space, we introduce the Hölder regularity condition HR($\Phi,\Psi$) and prove that, assuming (1.3), the volume regular condition V($\Phi$), and the upper heat kernel estimate UHK($\Psi$), the following conditions are equivalent:
+Since we do not assume the existence of any gradient operator associated with the metric measure Dirichlet space, in particular, a “carré du champ”, the above equivalence generalizes the corresponding result of Coulhon, Jiang, Koskela, Sikora [23, Theorem 1.2].
+In addition, we extend a classical result from the setting of Riemannian manifolds to metric measure Dirichlet spaces.
+We prove that the conjunction of UHK($\Psi$) and $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ implies NLE($\Psi$) on metric measure Dirichlet spaces.
+This generalizes the fact that, on Riemannian manifolds, the conjunction of upper heat kernel bounds and gradient estimates implies a lower heat kernel bound; see Cheeger, Yau [20] and Li, Yau [49].
 Throughout the paper, the letters $C,C_{1},C_{2},C_{A},C_{B}$ will always refer to some positive constants and may change at each occurrence.
@@
 2 Statement of the main results
-Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded metric measure Dirichlet (MMD) space, that is, $(X,d)$ is a locally compact separable unbounded metric space, $m$ is a positive Radon measure on $X$ with full support, $(\mathcal{E},\mathcal{F})$ is a strongly local regular Dirichlet form on $L^{2}(X;m)$.
-Throughout this paper, we always assume that all metric balls are relatively compact and $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is conservative.
-For any $x\in X$, for any $r\in(0,+\infty)$, denote the (metric) ball $B(x,r)=\{y\in X:d(x,y)<r\}$, denote $V(x,r)=m(B(x,r))$.
+Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded metric measure Dirichlet (MMD) space, that is, $(X,d)$ is a locally compact separable unbounded metric space, $m$ is a positive Radon measure on $X$ with full support, and $(\mathcal{E},\mathcal{F})$ is a strongly local regular Dirichlet form on $L^{2}(X;m)$.
+Although the results of this paper also apply to bounded spaces, we focus on the unbounded case, where the space may exhibit different behaviors at small and large scales.
+Throughout this paper, we always assume that all metric balls are relatively compact, and that $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is conservative.
+For any $x\in X$ and any $r\in(0,+\infty)$, denote the (metric) ball $B(x,r)=\{y\in X:d(x,y)<r\}$, denote $V(x,r)=m(B(x,r))$.
 If $B=B(x,r)$, then we denote $\delta B=B(x,\delta r)$ for any $\delta\in(0,+\infty)$.
 Let $C(X)$ denote the space of all real-valued continuous functions on $X$ and let $C_{c}(X)$ denote the space of all real-valued continuous functions on $X$ with compact support.
+Denote $\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{A}=\frac{1}{m(A)}\int_{A}$ and $u_{A}=\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{A}u\mathrm{d}m$ for any measurable set $A$ with $m(A)\in(0,+\infty)$ and any function $u$ such that the integral $\int_{A}u\mathrm{d}m$ is well-defined.
 Consider the strongly local regular Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$.
@@
 Let $\Gamma$ be the corresponding energy measure.
-Denote $\mathcal{E}_{1}(\cdot,\cdot)=\mathcal{E}(\cdot,\cdot)+(\cdot,\cdot)$, where (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) is the inner product in $L^{2}(X;m)$.
+Denote $\mathcal{E}_{1}(\cdot,\cdot)=\mathcal{E}(\cdot,\cdot)+(\cdot,\cdot)$, where $(\cdot,\cdot)$ is the inner product in $L^{2}(X;m)$.
 We refer to [27] for related results about Dirichlet forms.
-We need to introduce some conditions to state our main results.
+We introduce several conditions in order to state our main results.
 We say that the volume doubling condition VD holds if there exists $C_{VD}\in(0,+\infty)$ such that
@@
 See [34] for more details.
-We say that the heat kernel upper bound UHK($\Psi$) holds if there exist a properly exceptional set $\mathcal{N}$ and $C_{1}$, $C_{2}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y\in X\backslash\mathcal{N}$, we have
+We say that the upper heat kernel estimate UHK($\Psi$) holds if there exist a properly exceptional set $\mathcal{N}$ and $C_{1}$, $C_{2}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y\in X\backslash\mathcal{N}$, we have
 $p_{t}(x,y)\leq\frac{C_{1}}{V\left(x,\Psi^{-1}(t)\right)}\exp\left(-\Upsilon\left(C_{2}d(x,y),t\right)\right),$ UHK($\Psi$)
@@
 $p_{t}(x,y)\leq\begin{cases}\frac{C_{1}}{V\left(x,\Psi^{-1}(t)\right)}\exp\left(-C_{2}\left(\frac{d(x,y)}{t^{1/{\beta_{1}}}}\right)^{\frac{\beta_{1}}{\beta_{1}-1}}\right),&\text{if }t<d(x,y),\\ \frac{C_{1}}{V\left(x,\Psi^{-1}(t)\right)}\exp\left(-C_{2}\left(\frac{d(x,y)}{t^{1/{\beta_{2}}}}\right)^{\frac{\beta_{2}}{\beta_{2}-1}}\right),&\text{if }t\geq d(x,y).\end{cases}$
-If a lower bound, similar to UHK($\Psi$), with different constants $C_{i}$ also holds, then we say that HK($\Psi$) holds.
-We say that the heat kernel near-diagonal lower bound NLE($\Psi$) holds if there exist a properly exceptional set $\mathcal{N}$ and $C,\varepsilon\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y\in X\backslash\mathcal{N}$ with $d(x,y)<\varepsilon\Psi^{-1}(t)$, we have
+If a corresponding lower bound, analogous to UHK($\Psi$) but possibly with different constants $C_{i}$, also holds, then we say that HK($\Psi$) holds.
+We say that the near-diagonal lower bound for the heat kernel NLE($\Psi$) holds if there exist a properly exceptional set $\mathcal{N}$ and $C,\varepsilon\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y\in X\backslash\mathcal{N}$ with $d(x,y)<\varepsilon\Psi^{-1}(t)$, we have
 $p_{t}(x,y)\geq\frac{C}{V(x,\Psi^{-1}(t))}.$ NLE($\Psi$)
-If the metric $d$ is furthermore assumed to satisfy the chain condition, then the conjunction of UHK($\Psi$) and NLE($\Psi$) is equivalent to HK($\Psi$), see [32, Section 6].
-Let $D$ be an open subset of $X$. Let
+If the metric $d$ is furthermore assumed to satisfy the chain condition, then the conjunction of UHK($\Psi$) and NLE($\Psi$) is equivalent to HK($\Psi$), see [34, Section 6].
+Moreover, assuming HK($\Psi$), $d$ satisfies the chain condition; see [52, Theorem 2.11].
+If, in addition, V($\Phi$) holds, then $2\leq\beta_{i}\leq\alpha_{i}+1$ for $i=1,2$, see [53, THEOREM 2.1] and [32, Corollary 4.9].
+Let $D$ be an open subset of $X$ and
 $\mathcal{F}_{D}=\text{ the }\mathcal{E}_{1}\text{-closure of }\mathcal{F}\cap C_{c}(D).$
-Let $f\in L^{1}_{loc}(D)$. We say that $u\in\mathcal{F}$ is a solution of the Poisson equation $\Delta u=f$ in $D$ if
+Let $f\in L^{1}_{\mathrm{loc}}(D)$. We say that $u\in\mathcal{F}$ is a solution of the Poisson equation $\Delta u=f$ in $D$ if
 $\mathcal{E}(u,\varphi)=\int_{D}f\varphi\mathrm{d}m\text{ for any }\varphi\in\mathcal{F}\cap C_{c}(D).$
@@
 $x,y\in B$ with $x\neq y$, we have
-$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u(x)-u(y)|\leq C_{H}\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$ HR($\Phi,\Psi$)
-The above condition is some kind of generalization of the so-called generalized reversed Hölder inequality GRH($\Phi,\Psi$) introduced in [25].
-We say that the Hölder estimate HHK($\Phi,\Psi$) for the heat kernel holds if there exists $C_{HHK}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x_{1},x_{2},y_{1},y_{2}\in X$, we have
+$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u(x)-u(y)|\leq C_{H}\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m.$ HR($\Phi,\Psi$)
+The above condition can be viewed as a generalization of the so-called generalized reversed Hölder inequality GRH($\Phi,\Psi$) introduced in [26].
+We say that the Hölder estimate for the heat kernel HHK($\Phi,\Psi$) holds if there exists $C_{HHK}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x_{1},x_{2},y_{1},y_{2}\in X$, we have
 $|p_{t}(x_{1},y_{1})-p_{t}(x_{2},y_{2})|\leq C_{HHK}\frac{\left(\frac{\Psi}{\Phi}\right)(d(x_{1},x_{2}))+\left(\frac{\Psi}{\Phi}\right)(d(y_{1},y_{2}))}{t}.$ HHK($\Phi,\Psi$)
-The above condition implies that for any $t\in(0,+\infty)$, for any $x\in X$, $p_{t}(x,\cdot)$ is $\gamma_{1}$-Hölder continuous in the small scale and $\gamma_{2}$-Hölder continuous in the large scale.
-We say that the Hölder estimate $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ with exponential terms for the heat kernel holds if there exist $C_{1},C_{2}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y_{1},y_{2}\in X$, we have
+The above condition implies that for any $t\in(0,+\infty)$, for any $x\in X$, $p_{t}(x,\cdot)$ is $\gamma_{1}$-Hölder continuous at small scales and $\gamma_{2}$-Hölder continuous at large scales.
+We say that the Hölder estimate with exponential terms for the heat kernel $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ holds if there exist $C_{1},C_{2}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y_{1},y_{2}\in X$, we have
 $|p_{t}(x,y_{1})-p_{t}(x,y_{2})|$
@@
 $\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|P_{t}f(x)-P_{t}f(y)|\leq C_{wBE}\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\lVert{f}\rVert_{L^{\infty}(X;m)}.$ wBE($\Phi,\Psi$)
-The above condition is some kind of generalization of the weak Bakry-Émery curvature condition introduced in [1, Definition 3.1].
+The above condition can be viewed as a generalization of the weak Bakry-Émery non-negative curvature condition introduced in [2, Definition 3.1].
+Roughly speaking, letting $t\to 0+$ in wBE($\Phi,\Psi$) suggests that the space is strongly recurrent.
 The main result of this paper is as follows.
@@
 Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$) and
-UHK($\Psi$). Then the followings are equivalent.
+UHK($\Psi$), where $\Phi$ and $\Psi$ are given by (1.1) and (1.2), respectively, with parameters $\alpha_{i},\beta_{i}$, $i=1,2$, satisfying the strongly recurrent condition (1.3).
+Then the followings are equivalent.
   1. (1)
@@
   5. (5)
-The key novality of the above result is that the pointwise two-sided heat kernel estimates are enough to give some proper explicit Hölder regularity of harmonic functions.
-In previous literatures (for example [36, Section 5.3], [21, Proposition 4.5], [32, Section 5.3], [13, Section 4.2.1]), using Moser iteration technique or de Giorgi iteration technique and oscillation inequality, one can obtain only some Hölder exponent depending implicitly on other constants.
-Our proof will follow the following diagrams.
+The key novelty of the above result is that pointwise two-sided heat kernel estimates suffice to give explicit Hölder regularity for harmonic functions.
+In the existing literature (see, for example, [38, Section 5.3], [22, Proposition 4.5], [34, Section 5.3], and [14, Section 4.2.1]), approaches based on the Moser or De Giorgi iteration, together with oscillation inequalities, typically give only Hölder continuity with an exponent that depends implicitly on other constants.
+Our proof will follow the implication diagrams below:
 […]
@@
 Remark 2.2.
-In the setting of fractional metric spaces with fractional diffusions introduced in [5, Section 3], when the dimension $\alpha$ of the state space is smaller than the walk dimension $\beta$, or $\alpha_{1}=\alpha_{2}=\alpha<\beta_{1}=\beta_{2}=\beta$ in our setting, it was proved in [1, Lemma 3.4, Remark 3.6, Theorem 3.7] that
+In the setting of fractional metric spaces with fractional diffusions introduced in [6, Section 3], when the dimension $\alpha$ of the state space is strictly smaller than the walk dimension $\beta$, that is, when $\alpha_{1}=\alpha_{2}=\alpha<\beta_{1}=\beta_{2}=\beta$ in our notation, it was proved in [2, Lemma 3.4, Remark 3.6, Theorem 3.7] that
 […]
@@
 Remark 2.3.
-In the fractal setting, HHK($\Phi,\Psi$) was also proved on the unbounded Sierpiński gasket [14, Theorem 1.5 (c)] and on the unbounded Sierpiński carpet [8, Theorem 1.1 (e)].
+In the fractal setting, HHK($\Phi,\Psi$) was also established on the unbounded Sierpiński gasket (see [15, Theorem 1.5 (c)]) and on the unbounded Sierpiński carpet (see [9, Theorem 1.1 (e)]).
 Remark 2.4.
-The above equivalence can be regarded as some “Hölder” version generalization of the “gradient” version result [22, Theorem 1.2].
-It was proved that in a non-compact doubling Dirichlet metric measure space $(X,d,\mu,\mathscr{E})$ endowed with a “carré du champ”.
-Assume the upper Gaussian bound $(UE)$ for the heat kernel and a local $L^{\infty}$-Poincaré inequality $(P_{\infty,\text{loc}})$.
-Then the following conditions are equivalent.
-  • The quantitative reverse $L^{\infty}$-Hölder inequality ($\text{RH}_{\infty}$) for gradients of harmonic functions: There exists $C\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $u$ which is harmonic in $2B$, we have
-    $\lVert|\nabla u|\rVert_{L^{\infty}(B)}\leq\frac{C}{r}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}\mu.$ ($\text{RH}_{\infty}$)
-  • The pointwise Li-Yau gradient estimate ($\text{GLY}_{\infty}$) for heat kernel: There exist $C,c\in(0,+\infty)$ such that
+The above equivalence can be regarded as a Hölder-type analogue of the gradient-type result in [23, Theorem 1.2].
+In that work, it was proved that on a non-compact doubling Dirichlet metric measure space $(X,d,\mu,\mathscr{E})$ endowed with a “carré du champ”, assuming the upper Gaussian bound $(UE)$ for the heat kernel and a local $L^{\infty}$-Poincaré inequality $(P_{\infty,\text{loc}})$, the following conditions are equivalent.
+  • ($\text{RH}_{\infty}$) Quantitative reverse $L^{\infty}$-Hölder inequality for gradients of harmonic functions: There exists $C\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $u$ which is harmonic in $2B$, we have
+    $\lVert{|\nabla u|}\rVert_{L^{\infty}(B)}\leq\frac{C}{r}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{2B}|u|\mathrm{d}\mu.$ ($\text{RH}_{\infty}$)
+  • ($\text{GLY}_{\infty}$) Pointwise Li-Yau gradient estimate for the heat kernel: There exist $C,c\in(0,+\infty)$ such that
     $|\nabla_{x}h_{t}(x,y)|\leq\frac{C}{\sqrt{t}V(y,\sqrt{t})}\exp\left\{-c\frac{d(x,y)^{2}}{t}\right\}$ ($\text{GLY}_{\infty}$)
     for any $t\in(0,+\infty)$, for $\mu$-a.e. $x,y\in X$.
-  • The $L^{\infty}$-boundedness ($\text{G}_{\infty}$) of the gradient of the heat semi-group $|\nabla H_{t}|$: There exists $C\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, we have
+  • ($\text{G}_{\infty}$) $L^{\infty}$-boundedness of the gradient of the heat semi-group $|\nabla H_{t}|$: There exists $C\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, we have
     $\lVert{|\nabla H_{t}|}\rVert_{\infty\to\infty}\leq\frac{C}{\sqrt{t}},$
@@
     $\lVert{|\nabla H_{t}f|}\rVert_{L^{\infty}(X)}\leq\frac{C}{\sqrt{t}}\lVert{f}\rVert_{L^{\infty}(X)}.$ ($\text{G}_{\infty}$)
-Our conditions are “Hölder” version of the above “gradient” conditions as in Figure 3.
-“Hölder” conditions “gradient” conditions
+  • (GBE) A generalized Bakry-Émery condition: There exist $C,c\in(0,+\infty)$ such that for any $f\in W^{1,2}(X)$, for any $t\in(0,+\infty)$ and $\mu$-a.e. $x\in X$, we have
+    $\lvert\nabla H_{t}f(x)\rvert^{2}\leq CH_{ct}(\lvert\nabla f\rvert^{2})(x).$ (GBE)
+Our conditions can be viewed as Hölder-type analogues of the above gradient conditions, as illustrated in Figure 3.
+Hölder-type conditions gradient-type conditions
 HR($\Phi,\Psi$) ($\text{RH}_{\infty}$)
@@
 wBE($\Phi,\Psi$) ($\text{G}_{\infty}$)
-Figure 3: The correspondence between “Hölder” and “gradient” conditions
-In their proof to obtain ($\text{RH}_{\infty}$), they established a reproducing formula for harmonic functions using the finite propagation speed property, which may not be available on general MMD spaces.
-We leave it as an open question to generalize the reproducing formula technique to prove HR($\Phi,\Psi$) on MMD spaces or the following GRH($\Phi,\Psi$) on cable systems, in particular, to develop the corresponding finite propagation speed theory.
+Figure 3: The correspondence between Hölder-type and gradient-type conditions
+In their proof of ($\text{RH}_{\infty}$), they established a reproducing formula for harmonic functions using the finite propagation speed property, which may not be available on general MMD spaces.
+We leave it as an open question to extend the reproducing formula technique to prove HR($\Phi,\Psi$) on MMD spaces or the following GRH($\Phi,\Psi$) on cable systems, and in particular to develop a corresponding finite propagation speed theory in these settings.
 Remark 2.5.
-Let us compare the stability of “gradient” and “Hölder” conditions as follows.
-By a standard technique, it is easy to see that under $(UE)$, any one of the equivalent “gradient” conditions implies the heat kernel near-diagonal lower Gaussian bound $(NLE)$.
-In particular, under $(UE)$, we have ($\text{GLY}_{\infty}$) implies $(NLE)$.
+Let us compare the stability of the gradient-type and Hölder-type conditions as follows.
+By a standard argument, under $(UE)$, any of the equivalent gradient-type conditions implies the near-diagonal Gaussian lower bound $(NLE)$ for the heat kernel.
+In particular, assuming $(UE)$, we have ($\text{GLY}_{\infty}$) implies $(NLE)$.
 However, the converse is usually not true.
-Indeed, the conjunction of $(UE)$ and $(NLE)$, or equivalently, HK(2), is equivalent to the conjunction of the volume doubling condition and the scale-invariant $L^{2}$-Poincaré inequality on balls, hence is quasi-isometry stable.
-However, to obtain ($\text{GLY}_{\infty}$), some curvature assumptions are usually needed, see [47, 16], hence ($\text{GLY}_{\infty}$) is quasi-isometry unstable.
-For comparison, by our main result, Theorem 2.1, under UHK($\Psi$), all the “Hölder” conditions are equivalent to NLE($\Psi$).
-By the quasi-isometry stability of UHK($\Psi$) (see [2]) and the conjunction of UHK($\Psi$) and NLE($\Psi$) (see [31]), our “Hölder” conditions are quasi-isometry stable.
-Roughly speaking, the stability originates from the strongly recurrent condition, which requires our MMD spaces to behave like $\mathbb{R}^{1}$ in the classical Riemannian setting.
+Indeed, the conjunction of $(UE)$ and $(NLE)$, or equivalently, HK(2), is equivalent to the conjunction of the volume doubling condition and the scale-invariant $L^{2}$-Poincaré inequality on balls, and is therefore quasi-isometry stable.
+By contrast, to obtain ($\text{GLY}_{\infty}$), some curvature assumptions are usually needed, see [49, 17], and hence ($\text{GLY}_{\infty}$) is quasi-isometry unstable.
+For comparison, by our main result (Theorem 2.1), under UHK($\Psi$), all the Hölder-type conditions are equivalent to NLE($\Psi$).
+By the quasi-isometry stability of UHK($\Psi$) (see [3]), as well as the quasi-isometry stability of the conjunction of UHK($\Psi$) and NLE($\Psi$) (see [33]), our Hölder-type conditions are quasi-isometry stable.
+Roughly speaking, this stability stems from the strongly recurrent condition, which forces our MMD spaces to behave analogously to $\mathbb{R}^{1}$ in the classical Riemannian setting.
 Let $(V,E)$ be an infinite, locally bounded, connected (undirected) graph. Let
-$(X,d,m,\mathcal{E},\mathcal{F})$ be the corresponding unbounded cable system with an elementary gradient operator $\nabla$, see [25, Section 3] for more details.
+$(X,d,m,\mathcal{E},\mathcal{F})$ be the corresponding unbounded cable system, equipped with the elementary gradient operator $\nabla$, see [26, Section 3] for more details.
 Let $\alpha_{1}=1$, $\beta_{1}=2$, $\alpha_{2}=\alpha$, $\beta_{2}=\beta$.
 We say that the generalized reverse Hölder inequality GRH($\Phi,\Psi$) holds if there exists $C_{H}\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $u\in\mathcal{F}$ which is harmonic in $2B$, we have
-$\lVert\lvert\nabla u\rvert\rVert_{L^{\infty}(B)}\leq C_{H}\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$ GRH($\Phi,\Psi$)
-It is obvious that on an unbounded cable system, HR($\Phi,\Psi$) implies GRH($\Phi,\Psi$).
-Indeed, assume HR($\Phi,\Psi$), then the LHS of HR($\Phi,\Psi$) becomes $|u(x)-u(y)|/d(x,y)$ for $d(x,y)\in(0,1)$.
+$\lVert{\lvert\nabla u\rvert}\rVert_{L^{\infty}(B)}\leq C_{H}\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m.$ GRH($\Phi,\Psi$)
+It is obvious that, on an unbounded cable system, HR($\Phi,\Psi$) implies GRH($\Phi,\Psi$).
+Indeed, assuming HR($\Phi,\Psi$), the LHS of HR($\Phi,\Psi$) reduces to |u(x)u(y)|/d(x,y)|u(x)-u(y)|/d(x,y) for $d(x,y)\in(0,1)$.
 Letting $y\to x$ and using the Lebesgue density theorem, we obtain the limit $|\nabla u(x)|$.
-Taking essential supremum with respect to $x\in B$, we have GRH($\Phi,\Psi$).
-Roughly speaking, HR($\Phi,\Psi$) encodes more Hölder continuity information than GRH($\Phi,\Psi$) about the behavior of harmonic functions at large scale.
-We give some results as important applications of our main result, Theorem 2.1, on unbounded cable systems.
-Firstly, due to the fact that two-sided heat kernel estimates have already obtained on a large class of fractals (see [5, 34, 42, 43]), including the Sierpiński carpet (see [8, 9, 35]), we have GRH($\Phi,\Psi$) holds on a large class of fractal-like cable systems, including the Sierpiński carpet cable system, whose validity was left open in [25].
+Taking the essential supremum over $x\in B$, we have GRH($\Phi,\Psi$).
+Roughly speaking, HR($\Phi,\Psi$) encodes more Hölder continuity information than GRH($\Phi,\Psi$) about the behavior of harmonic functions at large scales.
+We give some results as important applications of our main result, Theorem 2.1, to unbounded cable systems.
+First, since two-sided heat kernel estimates have already been established for a large class of fractals (see [6, 36, 44, 45]), including the Sierpiński carpet (see [9, 10, 37]), we have GRH($\Phi,\Psi$) holds on a large class of fractal-like cable systems, including the Sierpiński carpet cable system, a case whose validity was left open in [26].
 Corollary 2.6.
 Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded cable system satisfying V($\Phi$) and
-HK($\Psi$) with $\alpha<\beta$. Then GRH($\Phi,\Psi$) holds. In particular, GRH($\Phi,\Psi$) holds on the Sierpiński carpet cable system with $\alpha=\log 8/\log 3<2<\beta\approx 2.09697$.
+HK($\Psi$) with $\alpha_{1}=1$, $\beta_{1}=2$, and $\alpha_{2}=\alpha<\beta=\beta_{2}$.
+Then GRH($\Phi,\Psi$) holds.
+In particular, GRH($\Phi,\Psi$) holds on the Sierpiński carpet cable system with $\alpha=\log 8/\log 3<2<\beta\approx 2.09697$.
 Remark 2.7.
-In [25], the authors verified directly GRH($\Phi,\Psi$) on the Vicsek cable system and the Sierpiński gasket cable system using certain Hölder regularity  [^1] The Hölder regularity follows indeed from certain extension algorithm for harmonic functions.
+In [26], the authors verified GRH($\Phi,\Psi$) directly on the Vicsek cable system and the Sierpiński gasket cable system, using certain Hölder regularity [^1] The Hölder regularity follows indeed from a certain extension algorithm for harmonic functions.
 of harmonic functions on the corresponding fractals [56], which is also available on some class of p.c.f.
-self-similar sets [54] without any technical difficulty, but not available on general fractals, especially on the Sierpiński carpet.
-Secondly, we improve the results about gradient estimate for heat kernel as follows.
+self-similar sets, but not available on general fractals, especially on the Sierpiński carpet.
+Second, we improve the results about gradient estimates for the heat kernel as follows.
 Corollary 2.8.
 Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded cable system satisfying V($\Phi$) and
-HK($\Psi$) with $\alpha<\beta$.
+HK($\Psi$) with $\alpha_{1}=1$, $\beta_{1}=2$, and $\alpha_{2}=\alpha<\beta=\beta_{2}$.
   1. (1)
@@
 Remark 2.9.
-(1) is certain improvement of [25, Theorem 1.1] and (2) is certain improvement of [25, Corollary 1.3], where GRH($\Phi,\Psi$) was needed.
-Now we can obtain gradient estimate result GHK($\Phi,\Psi$) without assuming gradient type conditions like GRH($\Phi,\Psi$).
-As a direct consequence of $\text{HHK}_{\text{exp}}(\Phi,\Psi)$, we also obtain the following Hölder estimate for the heat kernel on unbounded MMD spaces.
+(1) constitutes an improvement of [26, Theorem 1.1], while (2) improves upon [26, Corollary 1.3], where GRH($\Phi,\Psi$) was assumed.
+In the present work, we obtain the gradient estimate GHK($\Phi,\Psi$) without imposing gradient-type conditions such as GRH($\Phi,\Psi$).
+As a direct consequence of $\text{HHK}_{\text{exp}}(\Phi,\Psi)$, we obtain the following Hölder estimate for the heat kernel on unbounded MMD spaces.
 Corollary 2.10.
@@
 Remark 2.11.
-If $\beta_{1}=\alpha_{1}+1$, then the above Hölder estimate gives a local Lipschitz estimate, which will imply the existence and some estimate of the gradient for the heat kernel.
+If $\beta_{1}=\alpha_{1}+1$, then the above Hölder estimate gives a local Lipschitz estimate, which will imply the existence and some estimate of the gradient of the heat kernel.
 We give some examples as follows.
-  • Unbounded cable systems satisfying V($\Phi$) and HK($\Psi$). We have $\alpha_{1}=1$ and $\beta_{1}=2$ and this gives indeed also Corollary 2.8.
+  • Unbounded cable systems satisfying V($\Phi$) and HK($\Psi$). In this case, $\alpha_{1}=1$ and $\beta_{1}=2$. This indeed recovers Corollary 2.8.
   • The unbounded Vicsek set.
-We have $\alpha_{1}=\alpha_{2}=\log 5/\log 3$, $\beta_{1}=\beta_{2}=\log 15/\log 3$ and the above Hölder estimate is indeed a Lipschitz estimate, which was also given by Baudoin, Chen [15, Theorem 3.3].
-However, a resolvent kernel estimate [5, Theorem 3.40] was intrinsically needed in their proof.
-  • The blowups of the Vicsek set by another fractal.
-We have $\alpha_{1}=\log 5/\log 3$ and $\beta_{1}=\log 15/\log 3$.
-$\alpha_{2}$ and $\beta_{2}$ will be the Hausdorff dimension and walk dimension of another fractal, say the Sierpiński carpet or the unit interval, see Figure 7 or Figure 8, respectively.
-The estimates of the gradient for heat kernel can behave differently at small scale and large scale.
-We will give a detailed discussion in Section 9 about these two examples.
+Here $\alpha_{1}=\alpha_{2}=\log 5/\log 3$ and $\beta_{1}=\beta_{2}=\log 15/\log 3$, hence the above Hölder estimate is indeed a Lipschitz estimate, which was also given by Baudoin, Chen [16, Theorem 3.3].
+However, a resolvent kernel estimate [6, Theorem 3.40] was intrinsically needed in their proof.
+  • Blowups of the Vicsek set by another fractal.
+In this case, $\alpha_{1}=\log 5/\log 3$ and $\beta_{1}=\log 15/\log 3$, while $\alpha_{2}$ and $\beta_{2}$ will be given by the Hausdorff dimension and the walk dimension of another fractal, respectively (for instance, the Sierpiński carpet or the unit interval, see Figure 7 and Figure 8).
+The resulting gradient estimates for the heat kernel can behave differently at small and large scales.
+We will give a detailed discussion of these two examples in Section 9.
+In this paper, we assume a-priori the strongly recurrent condition (1.3), which singles out specific choices of $\Phi,\Psi$.
+This framework contains many classical fractals, including the Sierpiński gasket, the Sierpiński carpet, nested fractals, and p.c.f.
+self-similar sets.
+Starting from this setting, further examples can be constructed via tensor products.
+For instance, let $X$ be a space with Hausdorff dimension $\alpha$ and walk dimension $\beta>\alpha$.
+By taking the $n$-fold tensor product $X^{\otimes n}$, one obtains a space with Hausdorff dimension $n\alpha$ and walk dimension $\beta$.
+For $n$ sufficiently large, we have $n\alpha\geq\beta$, while the Hölder exponent of harmonic functions on $X^{\otimes n}$ remains equal to $\beta-\alpha$, the same as that on $X$; see [2, Proposition 3.8].
+For the general case in which that each factor space is equipped with scaling functions $\Phi,\Psi$, see [1, Proposition 3.7].
 […]
@@
 Let $D$ be an open subset of $X$. Denote by $\lambda_{1}(D)$ the smallest Dirichlet eigenvalue for $D$, that is,
-$\lambda_{1}(D)=\inf\left\{\frac{\mathcal{E}(u,u)}{\lVert u\rVert_{2}^{2}}:u\in\mathcal{F}_{D}\backslash\left\{0\right\}\right\}.$
+[…]
 We say that the relative Faber-Krahn inequality FK($\Psi$) holds if there exist $C_{F}\in(0,+\infty)$, $\nu\in(0,1)$ such that for any ball $B=B(x,r)$, for any open subset $D$ of $B$, we have
@@
 We say that the Poincaré inequality PI($\Psi$) holds if there exists $C_{P}\in(0,+\infty)$ such that for any ball $B=B(x,r)$, for any $u\in\mathcal{F}$, we have
-$\int_{B}|u-u_{B}|^{2}\mathrm{d}m\leq C_{P}\Psi(r)\int_{2B}\mathrm{d}\Gamma(u,u),$ PI($\Psi$)
-where $u_{A}$ is the mean value of $u$ on a measurable set $A$ with $m(A)\in(0,+\infty)$, that is,
-$u_{A}=\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{A}u\mathrm{d}m=\frac{1}{m(A)}\int_{A}u\mathrm{d}m.$
+$\int_{B}|u-u_{B}|^{2}\mathrm{d}m\leq C_{P}\Psi(r)\int_{2B}\mathrm{d}\Gamma(u,u).$ PI($\Psi$)
 Let $U,V$ be two open subsets of $X$ satisfying $U\subseteq\overline{{U}}\subseteq V$.
@@
 We say that the cutoff Sobolev inequality CS($\Psi$) holds if there exists $C_{S}\in(0,+\infty)$ such that for any $x\in X$, for any $R,r\in(0,+\infty)$, there exists a cutoff function $\varphi\in\mathcal{F}$ for $B(x,R)\subseteq B(x,R+r)$ such that for any $f\in\mathcal{F}$, we have
-$\int_{B(x,R+r)\backslash\overline{B(x,R)}}f^{2}\mathrm{d}\Gamma(\varphi,\varphi)$
-$\leq\frac{1}{8}\int_{B(x,R+r)\backslash\overline{B(x,R)}}\varphi^{2}\mathrm{d}\Gamma(f,f)+\frac{C_{S}}{\Psi(r)}\int_{B(x,R+r)\backslash\overline{B(x,R)}}f^{2}\mathrm{d}m.$ CS($\Psi$)
-We have the characterisations of heat kernel estimates as follows.
-Proposition 3.1.
-([2, Theorem 1.12]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD. Then the followings are equivalent.
+[…]
+[…]
+We have the characterizations of heat kernel estimates as follows.
+Proposition 3.1 ([3, Theorem 1.12]).
+Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD. Then the followings are equivalent.
   1. (1)
   2. (2)
-Proposition 3.2.
-([31, THEOREM 1.2]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$). Then the followings are equivalent.
+Proposition 3.2 ([33, THEOREM 1.2]).
+Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$). Then the followings are equivalent.
   1. (1)
@@
 […]
-First, we have the existence, the uniqueness and the regularity for the solutions of Poisson equation as follows.
-Lemma 4.1.
-([25, Lemma 2.2, Lemma 2.6]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying FK($\Psi,\nu$).
+First, we have existence, uniqueness, and regularity for solutions of the Poisson equation as follows.
+Lemma 4.1 ([26, Lemma 2.2, Lemma 2.6]).
+Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying FK($\Psi,\nu$).
 Let $q=\frac{2}{1-\nu}$.
@@
 There exists $C\in(0,+\infty)$ depending only on $\nu$, $C_{F}$ such that
-[…]
-Second, we have the pointwise estimate for the solutions of Poisson equation as follows.
-Lemma 4.2.
-([25, Lemma 2.7]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD, FK($\Psi,\nu$) and CS($\Psi$).
+$\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{B}|u|\mathrm{d}m\leq C\sqrt{\Psi(r)}\left(\frac{1}{m(B)}\mathcal{E}(u,u)\right)^{1/2}\leq C\Psi(r)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{B}|f|^{p}\mathrm{d}m\right)^{1/p}.$
+Second, we have pointwise estimates for solutions of the Poisson equation as follows.
+Lemma 4.2 ([26, Lemma 2.7]).
+Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD, FK($\Psi,\nu$) and CS($\Psi$).
 Let $q=\frac{2}{1-\nu}$.
@@
 $x\in B$, we have
-$|u(x)|\leq C\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+F_{1}(x)\right),$
+$|u(x)|\leq C\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{2B}|u|\mathrm{d}m+F_{1}(x)\right),$
 where
-$F_{1}(x)=\sum_{j\leq[\log_{2}r]}\Psi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
-The proof of the above two results in [25] was given for the case $\beta_{1}=2$. It is easy to see that the same proof holds for more general $\Psi$, including the case in our paper.
-Third, we have Hölder estimate for the solutions of Poisson equation as follows.
-The proof is similar to [25, Proposition 5.1] and [22, Theorem 3.2] to do gradient estimate.
-Some technical modifications will be needed.
+$F_{1}(x)=\sum_{j\leq[\log_{2}r]}\Psi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
+The proofs of the above two results in [26] were given for the case $\beta_{1}=2$. It is easy to see that the same proofs remain valid for more general $\Psi$, including the case in our paper.
+Third, we have Hölder estimates for solutions of the Poisson equation as follows.
+The proof is similar to that of [26, Proposition 5.1] and [23, Theorem 3.2] for gradient estimates, with some technical modifications.
 Proposition 4.3.
@@
 $x,y\in\frac{1}{16}B$ with $x\neq y$, we have
-$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u(x)-u(y)|\leq C\left(\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+F_{2}(x)+F_{2}(y)\right),$
+$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u(x)-u(y)|\leq C\left(\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{2B}|u|\mathrm{d}m+F_{2}(x)+F_{2}(y)\right),$
 where
-$F_{2}(x)=\sum_{j\leq[\log_{2}r]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
+$F_{2}(x)=\sum_{j\leq[\log_{2}r]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
 Proof.
@@
 For any $k=k_{0}+3,\ldots,k_{1}$, by Lemma 4.1, there exists a unique $u_{k}\in\mathcal{F}_{B(x,2^{k})}$ such that $\Delta u_{k}=f$ in $B(x,2^{k})$ and
-[…]
+$\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k-1})}|u_{k}|\mathrm{d}m\lesssim\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|u_{k}|\mathrm{d}m\lesssim\Psi\left(2^{k}\right)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
 Then
@@
 $\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|(u-u_{k_{1}})(x)-(u-u_{k_{1}})(y)|$
-$\lesssim\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|u-u_{k_{1}}|\mathrm{d}m$
-$\leq\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|u|\mathrm{d}m+\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|u_{k_{1}}|\mathrm{d}m\right)$
-$\lesssim\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+\Psi\left(2^{k_{1}}\right)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
-$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+\Phi(2^{k_{1}})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
+$\lesssim\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{1}})}|u-u_{k_{1}}|\mathrm{d}m$
+$\leq\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{1}})}|u|\mathrm{d}m+\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{1}})}|u_{k_{1}}|\mathrm{d}m\right)$
+$\lesssim\frac{\Phi(2^{k_{1}-1})}{\Psi(2^{k_{1}-1})}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m+\Psi\left(2^{k_{1}}\right)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{1}})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
+$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m+\Phi(2^{k_{1}})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{1}})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
 Similarly, for any $k=k_{0}+4,\ldots,k_{1}$, since $\Delta(u_{k}-u_{k-1})=0$ in $B(x,2^{k-1})$, by HR($\Phi,\Psi$), we have
 $\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|(u_{k}-u_{k-1})(x)-(u_{k}-u_{k-1})(y)|$
-$\lesssim\frac{\Phi(2^{k-2})}{\Psi(2^{k-2})}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k-1})}|u_{k}-u_{k-1}|\mathrm{d}m$
-$\leq\frac{\Phi(2^{k-2})}{\Psi(2^{k-2})}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k-1})}|u_{k}|\mathrm{d}m+\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k-1})}|u_{k-1}|\mathrm{d}m\right)$
-$\lesssim\frac{\Phi(2^{k-2})}{\Psi(2^{k-2})}\left(\Psi\left(2^{k}\right)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\Psi\left(2^{k-1}\right)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k-1})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
-$\lesssim\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\Phi(2^{k-1})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k-1})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
+$\lesssim\frac{\Phi(2^{k-2})}{\Psi(2^{k-2})}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k-1})}|u_{k}-u_{k-1}|\mathrm{d}m$
+$\leq\frac{\Phi(2^{k-2})}{\Psi(2^{k-2})}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k-1})}|u_{k}|\mathrm{d}m+\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k-1})}|u_{k-1}|\mathrm{d}m\right)$
+$\lesssim\frac{\Phi(2^{k-2})}{\Psi(2^{k-2})}\left(\Psi\left(2^{k}\right)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\Psi\left(2^{k-1}\right)\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k-1})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
+$\lesssim\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\Phi(2^{k-1})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k-1})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
 Since $\Delta u_{k_{0}+3}=f$ in $B(x,2^{k_{0}+3})$, by Lemma 4.2, we have
 $\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u_{k_{0}+3}(x)|$
-$\lesssim\frac{\Phi(d(x,y))}{\Psi(d(x,y))}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|u_{k_{0}+3}|\mathrm{d}m+\sum_{j\leq k_{0}+2}\Psi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
-$\lesssim\frac{\Phi(d(x,y))}{\Psi(d(x,y))}\left(\Psi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Psi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
-$\lesssim\frac{\Phi(2^{k_{0}+3})}{\Psi(2^{k_{0}+3})}\Psi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}$
-$+\sum_{j\leq k_{0}+2}\frac{\Phi(2^{k_{0}+2})}{\Psi(2^{k_{0}+2})}\frac{\Psi(2^{j})}{\Phi(2^{j})}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}$
-$\leq\Phi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p},$
+$\lesssim\frac{\Phi(d(x,y))}{\Psi(d(x,y))}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{0}+3})}|u_{k_{0}+3}|\mathrm{d}m+\sum_{j\leq k_{0}+2}\Psi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
+$\lesssim\frac{\Phi(d(x,y))}{\Psi(d(x,y))}\left(\Psi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Psi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
+$\lesssim\frac{\Phi(2^{k_{0}+3})}{\Psi(2^{k_{0}+3})}\Psi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}$
+$+\sum_{j\leq k_{0}+2}\frac{\Phi(2^{k_{0}+2})}{\Psi(2^{k_{0}+2})}\frac{\Psi(2^{j})}{\Phi(2^{j})}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}$
+$\leq\Phi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p},$
 where in the last inequality, we use the fact that $\gamma_{1},\gamma_{2}>0$, which implies that
@@
 $\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u_{k_{0}+3}(y)|$
-$\lesssim\frac{\Phi(d(x,y))}{\Psi(d(x,y))}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|u_{k_{0}+3}|\mathrm{d}m+\sum_{j\leq k_{0}+2}\Psi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
-$\lesssim\Phi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
+$\lesssim\frac{\Phi(d(x,y))}{\Psi(d(x,y))}\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{0}+3})}|u_{k_{0}+3}|\mathrm{d}m+\sum_{j\leq k_{0}+2}\Psi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
+$\lesssim\Phi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
 In summary, we have
 $\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u(x)-u(y)|$
-$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+\Phi(2^{k_{1}})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{1}})}|f|^{p}\mathrm{d}m\right)^{1/p}$
-$+\sum_{k=k_{0}+4}^{k_{1}}\left(\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\Phi(2^{k-1})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k-1})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
-$+\Phi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}$
-$+\Phi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}$
-$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+\sum_{k=k_{0}+3}^{k_{1}}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}$
-$+\sum_{k\leq k_{0}+3}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{k\leq k_{0}+2}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}$
-$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+\sum_{k\leq k_{1}}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{k\leq k_{1}}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
+$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m+\Phi(2^{k_{1}})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{1}})}|f|^{p}\mathrm{d}m\right)^{1/p}$
+$+\sum_{k=k_{0}+4}^{k_{1}}\left(\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\Phi(2^{k-1})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k-1})}|f|^{p}\mathrm{d}m\right)^{1/p}\right)$
+$+\Phi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}$
+$+\Phi(2^{k_{0}+3})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k_{0}+3})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{j\leq k_{0}+2}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y,2^{j})}|f|^{p}\mathrm{d}m\right)^{1/p}$
+$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m+\sum_{k=k_{0}+3}^{k_{1}}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}$
+$+\sum_{k\leq k_{0}+3}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{k\leq k_{0}+2}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}$
+$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m+\sum_{k\leq k_{1}}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{k\leq k_{1}}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
 […]
-The idea is to consider the heat equation “$\Delta(P_{t}f)=\frac{\partial}{\partial t}(P_{t}f)$” for fixed $t$ as a Poisson equation.
+The idea is to consider the heat equation $\Delta(P_{t}f)=\frac{\partial}{\partial t}(P_{t}f)$” for fixed $t$ as a Poisson equation.
 […]
@@
 $\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|P_{t}f(x)-P_{t}f(y)|$
-$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,32\Psi^{-1}(t))}|P_{t}f|\mathrm{d}m$
-$+\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}$
-$+\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}.$
+$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,32\Psi^{-1}(t))}|P_{t}f|\mathrm{d}m$
+$+\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}$
+$+\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}.$
 For the first term, we have
-$\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,32\Psi^{-1}(t))}|P_{t}f|\mathrm{d}m$
-$\lesssim\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,32\Psi^{-1}(t))}\lVert P_{t}f\rVert_{L^{\infty}(X;m)}\mathrm{d}m$
+$\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,32\Psi^{-1}(t))}|P_{t}f|\mathrm{d}m$
+$\lesssim\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,32\Psi^{-1}(t))}\lVert{P_{t}f}\rVert_{L^{\infty}(X;m)}\mathrm{d}m$
 $\leq\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\lVert{f}\rVert_{L^{\infty}(X;m)}.$
-For the second term, by UHK($\Psi$) and [23, THEOREM 4], we have the following estimate of the time derivative of the heat kernel
+For the second term, by UHK($\Psi$) and [24, THEOREM 4], we have the following estimate of the time derivative of the heat kernel
 $\left|\frac{\partial}{\partial t}p_{t}(z,w)\right|\leq\frac{C_{1}}{tV\left(z,\Psi^{-1}(t)\right)}\exp\left(-\Upsilon\left(C_{2}d(z,w),t\right)\right),$
@@
 By writing
-$\int_{X}=\int_{B(z,\Psi^{-1}(t))}+\sum_{n=0}^{\infty}\int_{B(z,2^{n+1}\Psi^{-1}(t))\backslash B(z,2^{n}\Psi^{-1}(t))},$
+[…]
 and using V($\Phi$), we have the integral
 $\int_{X}\frac{1}{V\left(z,\Psi^{-1}(t)\right)}\exp\left(-\Upsilon\left(C_{2}d(z,w),t\right)\right)m(\mathrm{d}w)$
-is bounded by some constant depending only on $C_{2}$ and $\Phi$, see also [3, Page 944, Line -5], [20, Equation (2.5)], hence
+is bounded by some constant depending only on $C_{2}$ and $\Phi$, see also [4, Page 944, Line -5], [21, Equation (2.5)], hence
 $|\frac{\partial}{\partial t}(P_{t}f)(z)|\lesssim\frac{\lVert{f}\rVert_{L^{\infty}(X;m)}}{t}.$
 Hence
-$\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}$
-$\lesssim\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}\left(\frac{\lVert f\rVert_{L^{\infty}(X;m)}}{t}\right)^{p}\mathrm{d}m\right)^{1/p}$
+$\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}$
+$\lesssim\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}\left(\frac{\lVert{f}\rVert_{L^{\infty}(X;m)}}{t}\right)^{p}\mathrm{d}m\right)^{1/p}$
 $=\frac{\lVert{f}\rVert_{L^{\infty}(X;m)}}{t}\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\frac{\Phi(2^{k})}{\Phi(16\Psi^{-1}(t))}\frac{\Phi(16\Psi^{-1}(t))}{\Phi(\Psi^{-1}(t))}\Phi(\Psi^{-1}(t))$
@@
 For the third term, similarly, we have
-$\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}\lesssim\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\lVert f\rVert_{L^{\infty}(X;m)}.$
+$\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}\lesssim\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\lVert{f}\rVert_{L^{\infty}(X;m)}.$
 In summary, we have
@@
 […]
-By UHK($\Psi$) and [23, THEOREM 4], we have the following estimates of the heat kernel
+By UHK($\Psi$) and [24, THEOREM 4], we have the following estimates of the heat kernel
 $p_{t}(x,y)\leq\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,y),t)\right),$
@@
 $\left|\frac{\partial}{\partial t}p_{t}(z,w)\right|\leq\frac{C_{1}}{tV\left(z,\Psi^{-1}(t)\right)}\exp\left(-\Upsilon\left(C_{2}d(z,w),t\right)\right).$
-Take any $t\in(0,+\infty)$, $x,y_{1},y_{2}\in X$. We now distinguish the following four cases.
+Take any $t\in(0,+\infty)$, $x,y_{1},y_{2}\in X$. We now distinguish the following four cases:
   1. (i)
@@
     $d(y_{1},y_{2})<\Psi^{-1}(t)$ and $d(x,y_{2})\geq 128\Psi^{-1}(t)$.
-For the case (i), we obtain the result directly from UHK($\Psi$) as follows.
+For case (i), we obtain the result directly from UHK($\Psi$) as follows.
 $|p_{t}(x,y_{1})-p_{t}(x,y_{2})|\leq p_{t}(x,y_{1})+p_{t}(x,y_{2})$
@@
 $\frac{\Phi(d(y_{1},y_{2}))}{\Psi(d(y_{1},y_{2}))}|p_{t}(x,y_{1})-p_{t}(x,y_{2})|$
-$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}p_{t}(x,z)m(\mathrm{d}z)$
-$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},2^{j})}\left|\frac{\partial}{\partial t}p_{t}(x,z)\right|^{p}m(\mathrm{d}z)\right)^{1/p}$
-$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{2},2^{j})}\left|\frac{\partial}{\partial t}p_{t}(x,z)\right|^{p}m(\mathrm{d}z)\right)^{1/p}.$ (6.1)
-For the case (ii), we have $d(x,y_{1})<128\Psi^{-1}(t)$ and $d(x,y_{2})<128\Psi^{-1}(t)$. By Equation (6.1), we have
+$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}p_{t}(x,z)m(\mathrm{d}z)$
+$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},2^{j})}\left|\frac{\partial}{\partial t}p_{t}(x,z)\right|^{p}m(\mathrm{d}z)\right)^{1/p}$
+$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{2},2^{j})}\left|\frac{\partial}{\partial t}p_{t}(x,z)\right|^{p}m(\mathrm{d}z)\right)^{1/p}.$ (6.1)
+For case (ii), we have $d(x,y_{1})<128\Psi^{-1}(t)$ and $d(x,y_{2})<128\Psi^{-1}(t)$. By Equation (6.1), we have
 $\frac{\Phi(d(y_{1},y_{2}))}{\Psi(d(y_{1},y_{2}))}|p_{t}(x,y_{1})-p_{t}(x,y_{2})|$
-$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}m(\mathrm{d}z)$
-$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
-$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
+$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}m(\mathrm{d}z)$
+$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
+$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
 $=\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}+2\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\frac{C_{1}}{tV(x,\Psi^{-1}(t))}$
@@
 $\leq\frac{e^{C_{3}}}{t}\left(\exp(-\Upsilon(C_{2}d(x,y_{1}),t))+\exp(-\Upsilon(C_{2}d(x,y_{2}),t))\right).$
-For the case (iii), we have $d(x,y_{1})\geq 128\Psi^{-1}(t)$. Then in Equation (6.1), for any $z\in B(y_{1},32\Psi^{-1}(t))$, we have
+For case (iii), we have $d(x,y_{1})\geq 128\Psi^{-1}(t)$. Then in Equation (6.1), for any $z\in B(y_{1},32\Psi^{-1}(t))$, we have
 $d(x,z)\geq d(x,y_{1})-d(y_{1},z)\geq d(x,y_{1})-32\Psi^{-1}(t)$
@@
 $\frac{\Phi(d(y_{1},y_{2}))}{\Psi(d(y_{1},y_{2}))}|p_{t}(x,y_{1})-p_{t}(x,y_{2})|$
-$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)m(\mathrm{d}z)$
-$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
-$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
-$\leq\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)m(\mathrm{d}z)$
-$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
-$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
+$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)m(\mathrm{d}z)$
+$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
+$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
+$\leq\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)m(\mathrm{d}z)$
+$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
+$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
 $=\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)$
@@
 $\asymp\frac{1}{t}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right).$
-For the case (iv), similar to the case (iii), we also have
+For case (iv), similar to case (iii), we also have
 $\frac{\Phi(d(y_{1},y_{2}))}{\Psi(d(y_{1},y_{2}))}|p_{t}(x,y_{1})-p_{t}(x,y_{2})|\lesssim\frac{1}{t}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{2}),t)\right).$
@@
 $\lesssim\frac{1}{t}\frac{\Psi(d(y_{1},y_{2}))}{\Phi(d(y_{1},y_{2}))}\left(\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)+\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{2}),t)\right)\right).$
-Lemma 6.1.
-([25, Lemma 5.2]) Let $A\in(0,+\infty)$. Then there exists $C\in(0,+\infty)$ depending only on $A$, $\beta_{1}$, $\beta_{2}$ such that
+Lemma 6.1 ([26, Lemma 5.2]).
+Let $A\in(0,+\infty)$. Then there exists $C\in(0,+\infty)$ depending only on $A$, $\beta_{1}$, $\beta_{2}$ such that
 $\sup_{t,s\in(0,+\infty)}\left(A\frac{\Psi^{-1}(t)}{\Psi^{-1}(s)}-\frac{t}{s}\right)\leq C.$
@@
 $=\sum_{n=N}^{\infty}\int_{B(x,2^{n+1}\Psi^{-1}(t))\backslash B(x,2^{n}\Psi^{-1}(t))}p_{t}(x,y)m(\mathrm{d}y)$
-$\leq\sum_{n=N}^{\infty}\int_{B(x,2^{n+1}\Psi^{-1}(t))\backslash B(x,2^{n}\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,y),t)\right)m(\mathrm{d}y)$
+[…]
 $\leq\sum_{n=N}^{\infty}C_{1}\frac{V(x,2^{n+1}\Psi^{-1}(t))}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}2^{n}\Psi^{-1}(t),t)\right).$
@@
 […]
-Recall that a Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is recurrent if for any non-negative function $f\in L^{1}(X;m)$, we have either $Gf=+\infty$ or $Gf=0$ $m$-a.e., where $G$ is the Green operator, see [26, Equation (1.6.2)].
+In this section, the strongly recurrent condition (1.3) is intrinsically required to ensure that the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ indeed induces a resistance form, as will be shown in Lemma 8.6, after which the arguments rely heavily on Kigami’s resistance form theory [46].
+Recall that a Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is recurrent if for any non-negative function $f\in L^{1}(X;m)$, we have either $Gf=+\infty$ or $Gf=0$ mm-a.e., where $G$ is the Green operator, see [27, Equation (1.6.2)].
 In our setting, the Dirichlet form is recurrent as follows.
@@
 for $m$-a.e. $x\in X$. Hence $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is recurrent. ∎
-We use some techniques from resistance forms.
-The theory of resistance forms was developed by Kigami [44].
-To begin with, let us introduce effective resistances.
-Let $A,B$ be two non-empty subsets of $X$.
-The effective resistance between $A$ and $B$ is defined as follows.
+To begin with, let us introduce effective resistances. Let $A,B$ be two non-empty subsets of $X$. The effective resistance between $A$ and $B$ is defined as follows.
 $R(A,B)=\left(\inf\left\{\mathcal{E}(u,u):u\in\mathcal{F},u=0\text{ on }A,u=1\text{ on }B\right\}\right)^{-1},$
@@
 In the strongly recurrent setting, two-sided heat kernel estimates and two-sided resistance estimates are indeed equivalent.
-This equivalence was proved in [12] on graphs, in [45] on graphs and resistance forms and in [37] on metric spaces using an analytical approach.
+This equivalence was proved in [13] for graphs, in [47] for graphs and resistance forms, and in [39] for metric spaces using an analytical approach.
 Under the assumptions of the above result, by Proposition 3.2, we have PI($\Psi$) and CS($\Psi$).
 The following result will play an important role in this section.
-Lemma 8.4.
-(Morrey-Sobolev inequality) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$) and PI($\Psi$).
+Lemma 8.4 (Morrey-Sobolev inequality).
+Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$) and PI($\Psi$).
 Then we have the following Morrey-Sobolev inequality MS($\Phi,\Psi$).
@@
 Hence any function in $\mathcal{F}$ has a continuous version, or equivalently, $\mathcal{F}\subseteq C(X)$.
-The proof is standard using telescopic technique, see also [12, Page 1654].
+The proof is standard using a telescopic technique, see also [13, Page 1654].
 Proof.
@@
 For any $n\in\mathbb{Z}$, denote $B_{n}=B(x,2^{-n}r)$. By V($\Phi$) and PI($\Psi$), we have
-$|u_{B_{n+1}}-u_{B_{n}}|\leq\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B_{n+1}}|u-u_{B_{n}}|\mathrm{d}m\lesssim\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B_{n}}|u-u_{B_{n}}|\mathrm{d}m\leq\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B_{n}}|u-u_{B_{n}}|^{2}\mathrm{d}m\right)^{1/2}$
+$|u_{B_{n+1}}-u_{B_{n}}|\leq\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B_{n+1}}|u-u_{B_{n}}|\mathrm{d}m\lesssim\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B_{n}}|u-u_{B_{n}}|\mathrm{d}m\leq\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B_{n}}|u-u_{B_{n}}|^{2}\mathrm{d}m\right)^{1/2}$
 $\lesssim\left(\frac{1}{\Phi(2^{-n}r)}\Psi(2^{-n}r)\int_{B_{n-1}}\mathrm{d}\Gamma(u,u)\right)^{1/2}\leq\left(\frac{\Psi(2^{-n}r)}{\Phi(2^{-n}r)}\mathcal{E}(u,u)\right)^{1/2},$
@@
 Note that
-$|u_{B(x,r)}-u_{B(y,r)}|\leq\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,r)}|u(z_{1})-u(z_{2})|m(\mathrm{d}z_{1})m(\mathrm{d}z_{2})$
-[…]
+$|u_{B(x,r)}-u_{B(y,r)}|\leq\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y,r)}|u(z_{1})-u(z_{2})|m(\mathrm{d}z_{1})m(\mathrm{d}z_{2})$
+$\leq\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y,r)}|u(z_{1})-u(z_{2})|^{2}m(\mathrm{d}z_{1})m(\mathrm{d}z_{2})\right)^{1/2}$
 $\asymp\frac{1}{\Phi(r)}\left(\int_{B(x,r)}\int_{B(y,r)}|u(z_{1})-u(z_{2})|^{2}m(\mathrm{d}z_{1})m(\mathrm{d}z_{2})\right)^{1/2}$
@@
 Proof of Proposition 8.2.
-The upper bound is obvious by MS($\Phi,\Psi$). We use the idea of an analytical proof of [37, Theorem 3.4] to give the lower bound. By the spectral calculus, it is easy to see that
+The upper bound is obvious by MS($\Phi,\Psi$). We use the idea of an analytical proof of [39, Theorem 3.4] to give the lower bound. By the spectral calculus, it is easy to see that
 $\mathcal{E}(P_{t}u,P_{t}u)\leq\frac{1}{2et}(u,u)\text{ for any }t\in(0,+\infty),u\in L^{2}(X;m),$
@@
 Let us recall the definition of resistance forms as follows.
-Definition 8.5.
-([44, DEFINITION 3.1]) Let $X$ be a set. A pair $(\mathcal{E},\mathcal{G})$ is called a resistance form on $X$ if the following five conditions are satisfied.
+Definition 8.5 ([46, DEFINITION 3.1]).
+Let $X$ be a set. A pair $(\mathcal{E},\mathcal{G})$ is called a resistance form on $X$ if the following five conditions are satisfied.
   1. (RF1)
@@
   2. (RF2)
-    Let $\sim$ be the equivalence relation on $\mathcal{G}$ given by $u,v\in\mathcal{G}$, $u\sim v$ if and only if $u-v$ is a constant function.
+    Let $\sim$ be the equivalence relation on $\mathcal{G}$ given by $\in\mathcal{G}, uvu\sim v$ if and only if uvu-v is a constant function.
 Then $(\mathcal{G}/\!\sim,\mathcal{E})$ is a Hilbert space.
@@
 We can extend $\mathcal{E}$ to $\mathcal{F}_{e}$ and $\mathcal{F}=\mathcal{F}_{e}\cap L^{2}(X;m)$.
-See [26, Section 1.5] for more details.
+See [27, Section 1.5] for more details.
 We can also define the harmonicity of a function in the extended Dirichlet space as follows.
@@
   4. (4)
-    Let $u\in\mathcal{F}_{e}$ be harmonic in a non-empty open subset $D$ of $X$. Then $\mathcal{E}(u,\varphi)=0$ for any $\varphi\in\mathcal{F}_{e}\subseteq C(X)$ with $\varphi=0$ on $X\backslash D$.
+    Let $u\in\mathcal{F}_{e}$ be harmonic in a non-empty open subset $D$ of $X$. Then $\mathcal{E}(u,\varphi)=0$ for any $\varphi\in\mathcal{F}_{e,D}$, where
+    $\mathcal{F}_{e,D}=\left\{u\in\mathcal{F}_{e}\subseteq C(X):u=0\text{ on }X\backslash D\right\}.$ (8.3)
 Proof.
 (1) It is obvious by the definition of $\mathcal{F}_{e}$ and Lemma 8.4.
-(2) Since the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is recurrent, by [26, Theorem 1.6.3], we have $1\in\mathcal{F}_{e}$ and $\mathcal{E}(1,1)=0$.
+(2) Since the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is recurrent, by [27, Theorem 1.6.3], we have $1\in\mathcal{F}_{e}$ and $\mathcal{E}(1,1)=0$.
 For $u\in\mathcal{F}_{e}$, if $\mathcal{E}(u,u)=0$, then by MS($\Phi,\Psi$), we have $u$ is a constant function.
@@
 By the definition of $\mathcal{F}_{e}$, we have $u\in\mathcal{F}_{e}$ and $\{u_{n}\}$ is $\mathcal{E}$-convergent to $u$.
-Hence $(\mathcal{F}_{e}/\sim,\mathcal{E})$ is complete, hence we have (RF2).
+Hence $(\mathcal{F}_{e}/\!\sim,\mathcal{E})$ is complete, then we have (RF2).
 (RF3) follows from the regularity of the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$.
-(RF5) follows from [26, Corollary 1.6.3].
+(RF4) follows from MS($\Phi,\Psi$).
+(RF5) follows from [27, Corollary 1.6.3].
 (3) We need to prove that
@@
 Taking supremum with respect to $u\in\mathcal{F}_{e}$ with $\mathcal{E}(u,u)>0$, we have $\geq$”.
-(4) Since $(\mathcal{E},\mathcal{F}_{e})$ is a resistance form on $X$, by [44, PROPOSITION 9.13], we have the continuity and the quasi-continuity coincide.
-Then $\mathcal{F}_{e,D}$ given by [26, Equation (2.3.14)] can be re-written as
-$\mathcal{F}_{e,D}:=\left\{u\in\mathcal{F}_{e}:\widetilde{u}=0\text{ q.e. on }X\backslash D\right\}=\left\{u\in\mathcal{F}_{e}\subseteq C(X):u=0\text{ on }X\backslash D\right\},$
+(4) Since $(\mathcal{E},\mathcal{F}_{e})$ is a resistance form on $X$, by [46, PROPOSITION 9.13], we have the continuity and the quasi-continuity coincide. By [27, Equation (2.3.14)], we have
+$\mathcal{F}_{e,D}=\left\{u\in\mathcal{F}_{e}:\widetilde{{u}}=0\text{ q.e. on }X\backslash D\right\},$
 where $\widetilde{{u}}$ is a quasi-continuous modification of $u\in\mathcal{F}_{e}$.
-Since the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is regular, by [26, Exercise 1.4.1], we have $\mathcal{F}\cap C_{c}(X)$ is a special standard core of $\mathcal{E}$.
-Then (4) follows from the equivalence in [26, Theorem 2.3.3 (ii)].
+Since the Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is regular, by [27, Exercise 1.4.1], we have $\mathcal{F}\cap C_{c}(X)$ is a special standard core of $\mathcal{E}$.
+Then (4) follows from the equivalence in [27, Theorem 2.3.3 (ii)].
 ∎
-By the resistance form theory, we can solve “boundary value problems” or take a trace of resistance forms as follows.
-Lemma 8.7.
-([44, Section 8]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$), UHK($\Psi$) and NLE($\Psi$).
-Let $Y\subsetneqq X$ be a non-empty closed subset and $\mathcal{F}_{e}|_{Y}=\{u|_{Y}:u\in\mathcal{F}_{e}\}$.
-Then for any $u\in\mathcal{F}_{e}$, there exists a unique $v\in\mathcal{F}_{e}$ which is harmonic in $X\backslash Y$ such that $v=u$ on $Y$.
-Moreover, $v$ is the unique function in $\mathcal{F}_{e}$ that minimizes the following variational problem
+By resistance form theory, we can solve “boundary value problems” and take traces of resistance forms as follows.
+Lemma 8.7 ([46, Section 8]).
+[…]
 $\inf\left\{\mathcal{E}(v,v):v\in\mathcal{F}_{e},v=u\text{ on }Y\right\}.$
@@
 We have the following maximum principle.
-Corollary 8.8.
-(Maximum principle) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying V($\Phi$), UHK($\Psi$) and NLE($\Psi$).
-Let $D\subsetneqq X$ be a non-empty open subset of $X$.
-Let $u,v\in\mathcal{F}_{e}$ be harmonic in $D$ and satisfy $u\leq v$ on $X\backslash D$, then $u\leq v$ in $D$.
-In particular, $\inf_{X\backslash D}u\leq u\leq\sup_{X\backslash D}u$.
+Corollary 8.8 (Maximum principle).
+[…]
 Proof.
-We only need to show that if $u\in\mathcal{F}_{e}$ is harmonic in $D$ and satisfies $u\leq 0$ on $X\backslash D$, then $u\leq 0$ in $D$.
-Indeed, $u\wedge 0\in\mathcal{F}_{e}$ satisfies $u\wedge 0=u$ on $X\backslash D$, by Lemma 8.7, we have
+[…]
 $\mathcal{E}(u,u)\geq\mathcal{E}(u\wedge 0,u\wedge 0)\geq\inf\left\{\mathcal{E}(v,v):v\in\mathcal{F}_{e},v=u\text{ on }X\backslash D\right\}=\mathcal{E}(u,u).$
 Hence the above inequalities are indeed equalities. By the uniqueness of $u\in\mathcal{F}_{e}$ from Lemma 8.7, we have $u\wedge 0=u$, hence $u\leq 0$ in $D$. ∎
-The effective resistance between certain two sets can be achieved by the energy of some functions as follows.
+For certain pairs of sets, the effective resistance can be realized by the energy of a suitable function, as follows.
 Corollary 8.9.
@@
 Let $A,B$ be two non-empty closed subsets of $X$ satisfying $A\cap B=\emptyset$ and $A\cup B\subsetneqq X$.
-If one of the following conditions holds.
+Assume that one of the following conditions holds:
   1. (a)
-    $X\backslash A$ is relatively compact and $B$ is compact.
+[…]
   2. (b)
@@
 Proof.
-Since $B$ is compact, $X\backslash A$ is open and $B\subseteq X\backslash A$, by the regularity of $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$, there exists $u\in\mathcal{F}\cap C_{c}(X)\subseteq\mathcal{F}_{e}$ such that $0\leq u\leq 1$ in $X$, $u=1$ on $B$ and $\mathrm{supp}(u)\subseteq X\backslash A$.
-Hence $u|_{A\cup B}=1_{B}\in\mathcal{F}_{e}|_{A\cup B}$, $(1-u)|_{A\cup B}=1_{A}\in\mathcal{F}_{e}|_{A\cup B}$.
+[…]
 Since $A\cup B\subsetneqq X$ is a non-empty closed subset, by Lemma 8.7, there exists a unique $v=h_{A\cup B}(u|_{A\cup B})=h_{A\cup B}(1_{B})\in\mathcal{F}_{e}$ which is harmonic in $X\backslash(A\cup B)$ and $v=u$ on $A\cup B$, that is, $v|_{A\cup B}=1_{B}$, or $v=0$ on $A$ and $v=1$ on $B$.
@@
 Since $\mathcal{F}\subseteq\mathcal{F}_{e}$, we only need to show $\geq$”.
-By the maximum principle Corollary 8.8, we have $0\leq v\leq 1$ in $X$.
-For (a), since $0\leq v\leq 1$ in $X$, $v=0$ on $A$ and $X\backslash A$ is a relatively compact open set, we have $v\in\mathcal{F}_{e}\cap L^{2}(X;m)=\mathcal{F}$, $\geq$” is obvious.
-For (b), since $A,B$ are both compact, by a similar proof to Lemma 8.6 (2), there exists an approximation sequence $\{v_{n}\}\subseteq\mathcal{F}\subseteq C(X)$ of $v\in\mathcal{F}_{e}$ that converges uniformly to $v$ on $A\cup B$.
-For any $\varepsilon\in(0,1/4)$, there exists $N\geq 1$ such that for any $n\geq N$, we have $|v_{n}-v|<\varepsilon$ on $A\cup B$, hence $-\varepsilon<v_{n}<\varepsilon$ on $A$ and $1-\varepsilon<v_{n}<1+\varepsilon$ on $B$.
+By the maximum principle (Corollary 8.8), we have $0\leq v\leq 1$ in $X$.
+[…]
+Assuming (b), since $A,B$ are both compact, by a similar proof to Lemma 8.6 (2), there exists an approximation sequence $\{v_{n}\}\subseteq\mathcal{F}\subseteq C(X)$ of $v\in\mathcal{F}_{e}$ that converges uniformly to $v$ on $A\cup B$.
+For any $\varepsilon\in(0,1/4)$, there exists $N\geq 1$ such that for any $n\geq N$, we have $|v_{n}-v|<\varepsilon$ on $A\cup B$, hence $\varepsilon<v_{n}<\varepsilon$ on $A$ and $1-\varepsilon<v_{n}<1+\varepsilon$ on $B$.
 Let $\varphi:\mathbb{R}\to\mathbb{R}$ be given by
 $\varphi(t)=\frac{\frac{1}{2}}{\frac{1}{2}-\varepsilon}\left((t\wedge(1-\varepsilon))\vee\varepsilon-\frac{1}{2}\right)+\frac{1}{2},t\in\mathbb{R},$
-that is, $\varphi\in C(\mathbb{R})$ satisfies $\varphi=0$ on $(-\infty,\varepsilon]$, $\varphi=1$ on $[1-\varepsilon,+\infty)$ and $\varphi$ is linear on $[\varepsilon,1-\varepsilon]$.
+that is, $\varphi\in C(\mathbb{R})$ satisfies $\varphi=0$ on $(-\infty,\varepsilon]$, $\varphi=1$ on $1-\varepsilon,+\infty)$ and $\varphi$ is linear on $\varepsilon,1-\varepsilon]$.
 Then $\varphi(0)=0$ and $|\varphi(t)-\varphi(s)|\leq\frac{\frac{1}{2}}{\frac{1}{2}-\varepsilon}|t-s|$ for any $t,s\in\mathbb{R}$.
@@
 Remark 8.10.
-We will be interested in the case $A=X\backslash B(x,r)$, $B=\{y\}$ for (a), and the case $A=\{x\}$, $B=\{y\}$ for (b), where $x,y\in X$, $r\in(0,+\infty)$.
+[…]
 Lemma 8.11.
@@
 $R(x_{0},X\backslash B(x_{0},r))\geq C\frac{\Psi(r)}{\Phi(r)}.$
-The proof is essentially the same as the proof of [12, LEMMA 2.4], [45, Lemma 4.1] and [37, Proposition 5.3].
+The proof is essentially the same as the proof of [13, LEMMA 2.4], [47, Lemma 4.1] and [39, Proposition 5.3].
 Proof.
@@
 $|h_{x}(y)|\leq C_{R}(2\varepsilon)^{\frac{{\gamma_{1}\wedge\gamma_{2}}}{2}}\leq\frac{1}{2}\text{ for any }y\in B(x,\varepsilon r).$
-By V($\Phi$), there exists $N\geq 1$ depending only on $C_{VR},\alpha_{1},\alpha_{2},\varepsilon$, hence only on $C_{VR}$, $C_{R}$, $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, $\beta_{2}$, such that $B\backslash\left(\frac{1}{2}B\right)$ can ben covered by at most $N$ balls with center in $B\backslash\left(\frac{1}{2}B\right)$ and radius $\varepsilon r$.
+By V($\Phi$), there exists $N\geq 1$ depending only on $C_{VR},\alpha_{1},\alpha_{2},\varepsilon$, hence only on $C_{VR}$, $C_{R}$, $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, $\beta_{2}$, such that $B\backslash\left(\frac{1}{2}B\right)$ can be covered by at most $N$ balls with center in $B\backslash\left(\frac{1}{2}B\right)$ and radius $\varepsilon r$.
 Take $x_{1},\ldots,x_{N}\in B\backslash\left(\frac{1}{2}B\right)$ such that $B\backslash\left(\frac{1}{2}B\right)\subseteq\cup_{1\leq i\leq N}B(x_{i},\varepsilon r)$, let $h=\min_{1\leq i\leq N}h_{x_{i}}$.
@@
 $\mathrm{osc}_{A}u=\sup\nolimits_{A}u-\inf\nolimits_{A}u$
-is the oscillation of the function $u\in l(A)$ on the set $A$.
-The idea of the proof is from a recent result [40, Theorem 6.27], where a general result in the setting of $p$-energy forms was given to prove certain $p$-resistance defines indeed a metric.
+is the oscillation of the function $u\in l(A)$ on the set AA.
+The idea of the following proof is from a recent result [42, Theorem 6.27], where a general result in the setting of $p$-energy forms was given to prove certain $p$-resistance defines indeed a metric.
+An alternative proof suggested by Kajino and Shimizu will be provided afterward.
 Proof.
-Let $Y=X\backslash B$, then $Y\subsetneqq X$ is a non-empty closed subset of $X$.
-For fixed $x,y\in B$, we have $u-u(x)\in\mathcal{F}_{e}$ is harmonic in $B\backslash\{x\}$, hence $u-u(x)=h_{Y\cup\{x\}}(u-u(x))$.
-By the maximum principle Corollary 8.8, we have $u-u(x)\leq(\mathrm{osc}_{Y}u)1_{Y}$ on $Y\cup\{x\}$, hence
+[…]
 $u-u(x)=h_{Y\cup\{x\}}(u-u(x))\leq h_{Y\cup\{x\}}\left((\mathrm{osc}_{Y}u)1_{Y}\right)=(\mathrm{osc}_{Y}u)h_{Y\cup\{x\}}(1_{Y}).$
-It remains to estimate $h_{Y\cup\{x\}}(1_{Y})(y)$. Indeed, by Corollary 8.9, we have
+Similarly, since $u-u(x)\geq-(\mathrm{osc}_{Y}u)1_{Y}$ on $Y\cup\{x\}$, we have
+$u-u(x)=h_{Y\cup\{x\}}(u-u(x))\geq-h_{Y\cup\{x\}}\left((\mathrm{osc}_{Y}u)1_{Y}\right)=-\left(\mathrm{osc}_{Y}u\right)h_{Y\cup\{x\}}(1_{Y}).$
+It suffices to estimate $h_{Y\cup\{x\}}(1_{Y})(y)$. Indeed, by Corollary 8.9, we have
 $R(x,X\backslash B)^{-1}=R(x,Y)^{-1}=\mathcal{E}(h_{Y\cup\{x\}}(1_{Y}),h_{Y\cup\{x\}}(1_{Y}))$
 $=\mathcal{E}(h_{Y\cup\{x\}}(1_{Y}),h_{Y\cup\{x\}}(1_{Y})-1)=-\mathcal{E}(h_{Y\cup\{x\}}(1_{Y}),h_{Y\cup\{x\}}(1_{\{x\}})).$
-Since $\mathcal{F}_{e}\ni h_{Y\cup\{x,y\}}(1_{\{x\}})-h_{Y\cup\{x\}}(1_{\{x\}})=0$ on $Y\cup\{x\}$ and $h_{Y\cup\{x\}}(1_{Y})\in\mathcal{F}_{e}$ is harmonic in $B\backslash\{x\}$, by Lemma 8.6 (4), we have
+[…]
 $\mathcal{E}(h_{Y\cup\{x\}}(1_{Y}),h_{Y\cup\{x,y\}}(1_{\{x\}})-h_{Y\cup\{x\}}(1_{\{x\}}))=0,$
@@
 $v_{1}$ $=h_{Y\cup\{x,y\}}(1_{\{x\}}),$
-then $u_{1},v_{1}\in\mathcal{F}_{e}$ are harmonic in $B\backslash\{x,y\}$. Let $u_{2}=u_{1}|_{Y\cup\{x,y\}}$, $v_{2}=v_{1}|_{Y\cup\{x,y\}}$, then $u_{2},v_{2}\in\mathcal{F}_{e}|_{Y\cup\{x,y\}}$ and
+[…]
 $\mathcal{E}(u_{1},v_{1})=\mathcal{E}|_{Y\cup\{x,y\}}(u_{2},v_{2}),$
@@
 Hence $u_{2}\wedge v_{2}=0$, $u_{2}\vee v_{2}=u_{2}+v_{2}$.
-By the following easy result Lemma 8.13, we have
+By the following easy result, Lemma 8.13, we have
 $\mathcal{E}(u_{1},v_{1})=\mathcal{E}|_{Y\cup\{x,y\}}(u_{2},v_{2})\leq 0,$
@@
 $=-h_{Y\cup\{x\}}(1_{Y})(y)\mathcal{E}(h_{\{x,y\}}(1_{\{y\}}),h_{Y\cup\{x,y\}}(1_{\{x\}})).$
-Since $\mathcal{F}_{e}\ni h_{Y\cup\{x,y\}}(1_{\{x\}})-h_{\{x,y\}}(1_{\{x\}})=0$ on $\{x,y\}$ and $h_{\{x,y\}}(1_{\{y\}})\in\mathcal{F}_{e}$ is harmonic in $X\backslash\{x,y\}$, by Lemma 8.6 (4), we have
+[…]
 $\mathcal{E}(h_{\{x,y\}}(1_{\{y\}}),h_{Y\cup\{x,y\}}(1_{\{x\}})-h_{\{x,y\}}(1_{\{x\}}))=0,$
@@
 which implies
-$h_{Y\cup\{x\}}(1_{Y})(y)\leq\frac{R(x,y)}{R(x,X\backslash B)}.$
-Thus we have
-$u(y)-u(x)\leq\left(\mathrm{osc}_{Y}u\right)\frac{R(x,y)}{R(x,X\backslash B)}.$
-Similarly, since $u-u(x)\geq-(\mathrm{osc}_{Y}u)1_{Y}$ on $Y\cup\{x\}$, we have
-$u-u(x)=h_{Y\cup\{x\}}(u-u(x))\geq-h_{Y\cup\{x\}}\left((\mathrm{osc}_{Y}u)1_{Y}\right)=-\left(\mathrm{osc}_{Y}u\right)h_{Y\cup\{x\}}(1_{Y}),$
-which implies
-$u(y)-u(x)\geq-\left(\mathrm{osc}_{Y}u\right)h_{Y\cup\{x\}}(1_{Y})(y)\geq-\left(\mathrm{osc}_{Y}u\right)\frac{R(x,y)}{R(x,X\backslash B)}.$
+[…]
 Therefore, we have
@@
 hence $\mathcal{E}(u,v)\leq 0$. ∎
+The alternative proof relies on the following Lipschitz continuity result for Green functions in the setting of Kigami’s resistance form theory.
+Lemma 8.14 ([46, THEOREM 4.1]).
+[…]
+$\lvert g_{D}(x,y)-g_{D}(x,z)\rvert\leq R(y,z).$
+An alternative proof of Proposition 8.12 using Lemma 8.14.
+[…]
+$\mathcal{E}(h,v)=v(x)\mathcal{E}(h,h)=\frac{v(x)}{R(x,X\backslash B)}\text{ for any }v\in\mathcal{F}_{e,B}.$
+[…]
+$h_{Y\cup\{x\}}(1_{Y})=1-h=\frac{g_{B}(x,x)-g_{B}(x,\cdot)}{R(x,X\backslash B)}\leq\frac{R(x,\cdot)}{R(x,X\backslash B)}.$
+The remainder of the argument proceeds as in the previous proof. ∎
 We need the following $L^{1}$-mean value inequality.
-Lemma 8.14.
-([31, THEOREM 6.3, LEMMA 9.2]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD, FK($\Psi$) and CS($\Psi$).
+Lemma 8.15 ([33, THEOREM 6.3, LEMMA 9.2]).
+Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD, FK($\Psi$) and CS($\Psi$).
 Then there exists $C\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $u\in\mathcal{F}$ which is harmonic in $2B$, we have
-$\lVert u\rVert_{L^{\infty}(B)}\leq C\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$
+$\lVert{u}\rVert_{L^{\infty}(B)}\leq C\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{2B}|u|\mathrm{d}m.$
 […]
@@
 By VD and Lemma 8.15, we have
-$M:=\lVert u\rVert_{L^{\infty}(\frac{3}{2}B)}\lesssim\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$
+$M:=\lVert{u}\rVert_{L^{\infty}(\frac{3}{2}B)}\lesssim\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m.$
 Let $v=(u\vee(-M))\wedge M$, then $v\in\mathcal{F}\cap L^{\infty}(X;m)$, $\lVert{v}\rVert_{L^{\infty}(X)}\leq M$ and $v=u$ in $\frac{3}{2}B$, hence $v\in\mathcal{F}\subseteq\mathcal{F}_{e}$ is bounded in $X$ and harmonic in $\frac{3}{2}B$.
@@
 hence
-$|u(x)-u(y)|\leq 2\lVert u\rVert_{L^{\infty}(\frac{3}{2}B)}\frac{R(x,y)}{R(x,X\backslash B(x,\frac{r}{2}))}\lesssim\frac{R(x,y)}{R(x,X\backslash B(x,\frac{r}{2}))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$
+$|u(x)-u(y)|\leq 2\lVert{u}\rVert_{L^{\infty}(\frac{3}{2}B)}\frac{R(x,y)}{R(x,X\backslash B(x,\frac{r}{2}))}\lesssim\frac{R(x,y)}{R(x,X\backslash B(x,\frac{r}{2}))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m.$
 By R($\Phi,\Psi$) and Lemma 8.11, we have
-$|u(x)-u(y)|\lesssim\frac{\left(\frac{\Psi}{\Phi}\right)(d(x,y))}{\left(\frac{\Psi}{\Phi}\right)(\frac{r}{2})}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m\asymp\frac{\left(\frac{\Psi}{\Phi}\right)(d(x,y))}{\left(\frac{\Psi}{\Phi}\right)(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m,$
+$|u(x)-u(y)|\lesssim\frac{\left(\frac{\Psi}{\Phi}\right)(d(x,y))}{\left(\frac{\Psi}{\Phi}\right)(\frac{r}{2})}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m\asymp\frac{\left(\frac{\Psi}{\Phi}\right)(d(x,y))}{\left(\frac{\Psi}{\Phi}\right)(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m,$
 hence HR($\Phi,\Psi$) holds. ∎
 Remark 8.16.
-A natural question is to consider the sharpness of the Hölder exponents $\beta_{1}-\alpha_{1}$ and $\beta_{2}-\alpha_{2}$ from HR($\Phi,\Psi$).
+A natural question concerns the sharpness of the Hölder exponents $\beta_{1}-\alpha_{1}$ and $\beta_{2}-\alpha_{2}$ arising from HR($\Phi,\Psi$).
 Indeed, on a class of unbounded p.c.f.
-self-similar sets (see [54]) and the corresponding cable systems, these exponents are sharp, see the proof of [25, Proposition 4.1, Proposition 4.3] for the Vicsek and the SG cable systems.
-However, on the unbounded SC and the SC cable system, it was conjectured in [1, Conjecture 5.4, Open Question 5.5] that these exponents are not sharp, and a new fractal dimension, that is, the so-called topological Hausdorff dimension (see [4]), should be involved.
+self-similar sets and the corresponding cable systems, these exponents are sharp, see the proof of [26, Proposition 4.1, Proposition 4.3] for the Vicsek and the Sierpiński gasket cable systems.
+However, on the unbounded Sierpiński carpet and the Sierpiński carpet cable system, it was conjectured in [2, Conjecture 5.4, Open Question 5.5] that these exponents are not sharp, and a new fractal dimension, the so-called topological Hausdorff dimension (see [5]), should be involved.
 9 Some examples related to the Sierpiński carpet and the Vicsek set
-In this section, we give some examples related to the Sierpiński carpet and the Vicsek set including the blowups of a fractal by another different fractal. In $\mathbb{R}^{2}$, let
+In this section, we give some examples related to the Sierpiński carpet and the Vicsek set, including blowups of one fractal by another. In $\mathbb{R}^{2}$, let
 $p_{1}=(0,0),p_{2}=(\frac{1}{2},0),p_{3}=(1,0),p_{4}=(1,\frac{1}{2}),$ (9.1)
@@
 Let $\alpha=\log 8/\log 3$.
-By [42, Theorem 1.5.7, Proposition 1.5.8], the Hausdorff dimension of $K$ is $\alpha$ and the $\alpha$-dimensional Hausdorff measure $\mathcal{H}^{\alpha}(K)\in(0,+\infty)$.
-The construction of Brownian motion on the Sierpiński carpet was given by [6] and [46] using approximation of stochastic processes.
+By [44, Theorem 1.5.7, Proposition 1.5.8], the Hausdorff dimension of $K$ is $\alpha$ and the $\alpha$-dimensional Hausdorff measure $\mathcal{H}^{\alpha}(K)\in(0,+\infty)$.
+The construction of Brownian motion on the Sierpiński carpet was given by [7] and [48] using approximations of stochastic processes.
 See also [35] for another construction using $\Gamma$-convergence of quadratic forms, [10] for the construction on higher dimensional Sierpiński carpets, [18] for the construction on unconstrained Sierpiński carpets.
@@
 It is obvious that V($\Phi$) holds.
-It was proved in [8, Theorem 1.1] that HK($\Psi$) and HHK($\Phi,\Psi$) hold.
-In a general setting of fractional diffusions introduced in [5, Section 3], it was proved in [1, Theorem 3.7, Lemma 3.4] that wBE($\Phi,\Psi$) and HHK($\Phi,\Psi$) hold.
-It was proved in [8, Theorem 8.3] that all functions in the domain of the generator of the Dirichlet form are $(\beta-\alpha)$-Hölder continuous.
-By the main result, Theorem 2.1, we now have the Hölder regularity for harmonic functions as HR($\Phi,\Psi$) and the Hölder estimate $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ with exponential terms for the heat kernel.
+It was proved in [9, Theorem 1.1] that HK($\Psi$) and HHK($\Phi,\Psi$) hold.
+In a general setting of fractional diffusions introduced in [6, Section 3], it was proved in [2, Theorem 3.7, Lemma 3.4] that wBE($\Phi,\Psi$) and HHK($\Phi,\Psi$) hold.
+It was proved in [9, Theorem 8.3] that all functions in the domain of the generator of the Dirichlet form are $(\beta-\alpha)$-Hölder continuous.
+By the main result, Theorem 2.1, we now have the Hölder regularity for harmonic functions HR($\Phi,\Psi$) and the Hölder estimate $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ with exponential terms for the heat kernel.
 Secondly, we consider the Sierpiński carpet cable system.
-Let $V_{0}=\{p_{1},\ldots,p_{8}\}$ and $V_{n+1}=\cup_{i=1}^{8}g_{i}(V_{n})$ for any $n\geq 0$, then $\{V_{n}\}_{n\geq 0}$ is an increasing sequence of finite subsets of $K$ and the closure of $\cup_{n\geq 0}V_{n}$ with respect to the Euclidean topology is $K$.
+Let $V_{0}=\{p_{1},\ldots,p_{8}\}$ and $V_{n+1}=\cup_{i=1}^{8}g_{i}(V_{n})$ for any $n\geq 0$, then $\{V_{n}\}_{n\geq 0}$ is an increasing sequence of finite subsets of $K$, and the closure of $\cup_{n\geq 0}V_{n}$ with respect to the Euclidean topology is $K$.
 For any $n\geq 0$, let $V^{(n)}=3^{n}V_{n}$, see Figure 4 for $V^{(0)}$, $V^{(1)}$ and $V^{(2)}$.
@@
 $E=\left\{\{p,q\}:p,q\in V,|p-q|=\frac{1}{2}\right\},$
-then $(V,E)$ is an infinite, locally bounded, connected (undirected) graph, whose corresponding cable system is the Sierpiński carpet cable system, see [25, Section 3] for more detailed construction.
+then $(V,E)$ is an infinite, locally bounded, connected (undirected) graph, whose corresponding cable system is the Sierpiński carpet cable system, see [26, Section 3] for a more detailed construction.
 Let $\alpha_{1}=1$, $\beta_{1}=2$, $\alpha_{2}=\alpha$, $\beta_{2}=\beta$.
@@
 Moreover, we also have GHK($\Phi,\Psi$), which can also be given directly by $\text{HHK}_{\text{exp}}(\Phi,\Psi)$.
-Thirdly, we consider the blowups of fractals related to the Sierpiński carpet and the Vicsek set.
-However, we blowup a fractal by another different fractal.
-We consider the setting in $\mathbb{R}^{2}$.
-Let $K,M$ be one of the following fractals in $\mathbb{R}^{2}$: the unit interval $[0,1]\times\{0\}$, the Vicsek set and the Sierpiński carpet.
-Let $\alpha_{1}$ and $\beta_{1}$ be the Hausdorff dimension and walk dimension of $K$, $\alpha_{2}$ and $\beta_{2}$ the Hausdorff dimension and walk dimension of $M$.
+Thirdly, we consider blowups of fractals related to the Sierpiński carpet and the Vicsek set. However, we blowup one fractal by a different fractal. In what follows, we work in $\mathbb{R}^{2}$.
+Let $K,M$ be one of the following fractals in $\mathbb{R}^{2}$: the unit interval $0,1]\times\{0\}$, the Vicsek set, and the Sierpiński carpet.
+Let $\alpha_{1}$ and $\beta_{1}$ be the Hausdorff dimension and the walk dimension of $K$, and let $\alpha_{2}$ and $\beta_{2}$ be the Hausdorff dimension and the walk dimension of $M$.
 Here, $K$ serves as the “cell” fractal and $M$ provides the “model” to blowup.
@@
 Obviously, $(X,d,m,\mathcal{E},\mathcal{F})$ is an unbounded MMD space satisfying V($\Phi$).
-Since two-sided heat kernel estimates hold on $K,M$, by [30, Theorem 3.14], we have the elliptic Harnack inequalities and the resistance estimates both hold on $K,M$.
+Since two-sided heat kernel estimates hold on $K,M$, by [31, Theorem 3.14], we have the elliptic Harnack inequalities and the resistance estimates both hold on $K,M$.
 Combining these two conditions on $K,M$, we obtain easily the corresponding elliptic Harnack inequalities and the resistance estimates on $X$.
-Then by [30, Theorem 3.14] again, we have HK($\Psi$) holds on $X$.
+Then by [31, Theorem 3.14] again, we have HK($\Psi$) holds on $X$.
 Indeed, in the strongly recurrent setting, only resistance estimates are enough for the above argument, see [13, 47, 39].
@@
     for any $x,y_{1},y_{2}\in X$ with $d(y_{1},y_{2})<1$.
-    Let $S_{0}=[(0,0),(1,1)]\cup[(0,1),(1,0)]$ be the union of the diagonals of the unit square $[0,1]^{2}\subseteq\mathbb{R}^{2}$.
+    Let $S_{0}=[(0,0),(1,1)]\cup[(0,1),(1,0)]$ be the union of the diagonals of the unit square $0,1]^{2}\subseteq\mathbb{R}^{2}$.
 For any $n\geq 0$, let $S_{n+1}=\cup_{i=1}^{N^{(1)}}g_{i}^{(1)}(S_{n})$.
-Then $\{S_{n}\}_{n\geq 0}$ is an increasing sequence of subsets of $K$ and the closure of $\cup_{n\geq 0}S_{n}$ is $K$.
+Then $\{S_{n}\}_{n\geq 0}$ is an increasing sequence of subsets of $K$, and the closure of $\cup_{n\geq 0}S_{n}$ is $K$.
 Let $\mathcal{S}_{0}=\cup_{n\geq 0}S_{n}$, $\nu^{(0)}$ the measure on $\mathcal{S}_{0}$ such that for any $n\geq 0$, $\nu^{(0)}|_{S_{n}}$ is the normalized 1-dimensional Hausdorff measure satisfying
@@
 For any cell $W\subseteq X$, we have $\mathcal{S}_{W}:=\mathcal{S}\cap W$ is a translation of $\mathcal{S}^{(0)}=\mathcal{S}_{0}$.
-For any two cells $W_{1},W_{2}\subseteq X$, we have either the cardinal #$\#(\mathcal{S}_{W_{1}}\cap\mathcal{S}_{W_{2}})\in\{0,1\}$ or $\mathcal{S}_{W_{1}}\cap\mathcal{S}_{W_{2}}$ is countably infinite, hence there exists a unique $\nu$ on $\mathcal{S}$ such that for any cell $W$, $\nu|_{\mathcal{S}_{W}}$ is the push forward of $\nu^{(0)}$ under the translation from $\mathcal{S}^{(0)}=\mathcal{S}_{0}$ to $\mathcal{S}_{W}$.
+For any two cells $W_{1},W_{2}\subseteq X$, we have either the cardinal $(\mathcal{S}_{W_{1}}\cap\mathcal{S}_{W_{2}})\in\{0,1\}$ or $\mathcal{S}_{W_{1}}\cap\mathcal{S}_{W_{2}}$ is countably infinite, hence there exists a unique $\nu$ on $\mathcal{S}$ such that for any cell $W$, $\nu|_{\mathcal{S}_{W}}$ is the push forward of $\nu^{(0)}$ under the translation from $\mathcal{S}^{(0)}=\mathcal{S}_{0}$ to $\mathcal{S}_{W}$.
 Note that $\nu(B(x_{0}),r)=+\infty$ for any ball $B(x_{0},r)$ and $\nu\perp m$, that is, $\nu$ and $m$ are mutually singular.
@@
     $|\partial_{y}p_{t}(x,y)|\leq\frac{C_{1}}{t}\exp\left(-\Upsilon\left(C_{2}d(x,y),t\right)\right)$
-    $\leq\begin{cases}\frac{C_{3}}{t}\exp\left(-C_{4}\left(\frac{d(x,y)}{t^{1/\beta_{1}}}\right)^{\frac{\beta_{1}}{\beta_{1}-1}}\right),&\text{ if }t<d(x,y),\\ \frac{C_{3}}{t}\exp\left(-C_{4}\left(\frac{d(x,y)}{t^{1/\beta_{2}}}\right)^{\frac{\beta_{2}}{\beta_{2}-1}}\right),&\text{ if }t\geq d(x,y).\\ \end{cases}$ (9.3)
+[…]
     Note that $\partial p_{t}(x,\cdot)$ exists and has the above estimate $\nu$-a.e.
 instead of $m$-a.e..
-At small scale, $\beta_{1}=\log 15/\log 3$, the gradient estimate has the sub-Gaussian behavior corresponding to the Vicsek set, while at large scale, $\beta_{2}\approx 2.09697$, the gradient estimate has the sub-Gaussian behavior corresponding to the Sierpiński carpet.
-See [15, Section 3] for the results about Lipschitz and gradient estimates for heat kernel on unbounded Vicsek set, that is, the blowup of the Vicsek set by the Vicsek set itself.
+At small scales, $\beta_{1}=\log 15/\log 3$, the gradient estimate has the sub-Gaussian behavior corresponding to the Vicsek set, while at large scales, $\beta_{2}\approx 2.09697$, the gradient estimate has the sub-Gaussian behavior corresponding to the Sierpiński carpet.
+See [16, Section 3] for the results about Lipschitz and gradient estimates for heat kernel on unbounded Vicsek set, that is, the blowup of the Vicsek set by the Vicsek set itself.
   4. (4)
@@
     Similar to the blowup of the Vicsek set by the Sierpiński carpet, we also have the Lipschitz estimate, Equation (9.2) and the gradient estimate, Equation (9.3).
-However, at small scale, $\beta_{1}=\log 15/\log 3$, the gradient estimate has the sub-Gaussian behavior corresponding to the Vicsek set, while at large scale, $\beta_{2}=2$, the gradient estimate has the Gaussian behavior corresponding to the unit interval.
+However, at small scales, $\beta_{1}=\log 15/\log 3$, the gradient estimate has the sub-Gaussian behavior corresponding to the Vicsek set, while at large scales, $\beta_{2}=2$, the gradient estimate has the Gaussian behavior corresponding to the unit interval.